### v1

# Algebraic geometry of bubbling Kähler metrics

Abstract.
We give an algebro-geometric or non-archimedean framework to study bubbling phenomena of Kähler metrics with Euclidean volume growth, after [DS17, Sun23, dBS23]. In particular, for any degenerating family to log terminal singularity, we algebraically construct a finite sequence of birational modifications of the family with milder degenerations, and compare with analytic bubbling constructions in [Sun23, dBS23]. We also provide approaches in terms of coordinates and valuations. Our discussion partially depends on the general framework of stability theory in [Od24b] after [HL14, AHLH23].
1. Introduction
We make a purely algebro-geometric study of degenerations of normal singularities in this paper, motivated by the recent works on asymptotic behaviour of Kähler metrics [DS17, Sun23, dBS23]. Our main setup is as follows.
Setup 1.
Let $\pi\colon\mathcal{X}\twoheadrightarrow\Delta\ni 0$, be either one of them:
  1. (a)
    (analytic setup over $\mathbb{C}$) $\Delta:=\{t\in\mathbb{C}\mid|t|<1\}$, $\pi$ is a flat klt (kawamata-log-terminal) affine 111meaning that $\pi$ is achievable as a restrction of ${\rm pr}_{1}\colon(\Delta\times\mathbb{C}^{l})\twoheadrightarrow\Delta$ for $l\in\mathbb{Z}_{>0}$ to an analytic closed subset of $(\Delta\times\mathbb{C}^{l})$ map of relative dimension $n$ over the base field $\mathbbm{k}=\mathbb{C}$ i.e., each fiber is a $n$-dimensional log terminal complex variety.
  2. (b)
    (algebraic setup over a field of characteristic 00) $\Delta$ is an affine curve over a field $\mathbbm{k}$ of characteristic 00, $\pi$ is a flat klt affine morphism between $\mathbbm{k}$-varieties of relative dimension $n$.
We also assume $K_{\mathcal{X}}$ to be $\mathbb{Q}$-Cartier as our main interest comes from Kähler-Einstein geometry. We set $X:=\mathcal{X}_{0}:=\pi^{-1}(0)$. We further fix a holomorphic (resp., algebraic) section $\sigma$ with $\sigma(0)=x$.
In the next section, we algebraically construct (almost) canonical, finite ordered sequences of modifications of $\mathcal{X}$ with milder degenerations. For instance, if the general fibers are $2$-dimensional smooth, degenerating to $X$ with a $2$-dimensional $A_{k}$-singularity, then we obtain other fillings with finite number of $A_{i}$-singularities with smaller Milnor number $i$ (cf., [dBS23]) and terminate at smooth filling. The process is different from the classical simultaneous resolution of Brieskorn-Grothendieck type in an essential manner. Morally speaking, our algorithm gives canonical “limit singularities” of various rescales beyond the given $X$ itself or its (metric) tangent cone, reflecting the whole formation process of the singularities of $X$, in the context of K-stability. While we explain its detailed meaning later, here is the brief overview of the differential geometric background.
Suppose we have a “good” family of (weak) Kähler metrics $g_{t}$ on $\mathcal{X}_{t}=\pi^{-1}(t)$ in the pluri-potential theoritic sense (cf., [BBEGZ19]), for $t\neq 0$, especially of Kähler-Einstein metrics, we would like to consider and answer the following question:
Main Question (see [Sun23, dBS23]).
How can we determine the set of possible non-collapsing pointed Gromov-Hausdorff limits of $(x_{t}\in\mathcal{X}_{t},c_{t}^{2}g_{t})$ for $t\to 0$ where $c_{t}>0$ are some sequences of diverging rescaling parameters?
In this paper, following [Sun23, dBS23], we call such a non-collapsing pointed Gromov-Hausdorff limit simply as bubbles of the family $(\mathcal{X}_{t},g_{t})$ for $t\to 0$. By the groundbreaking work [DS17, 1.4], extending some classical cases, such limits often naturally admits structure of an affine log terminal variety with (weak) Ricci-flat Kähler metric. Indeed, our aforementioned main construction of affine varieties is expected to underlie these bubbles. They have Euclidean volume growth simply by the Bishop’s inequality. This is a (Kähler) metric version of so-called bubbling phenomenon since [SU81], as later also studied by [Nak88, And89, Ban90, BKN89] in mainly $4$-dimensional Riemannian geometric setup, which in particular found out the finite energy (the $L^{2}$-norm of the curvature) rescaling provides smooth asymptotically locally Euclidean (ALE) manifolds limits as the deepest bubbles. Then, recently we witnessed a breakthrough [DS17, Sun23, dBS23] in the higher dimensional Kähler setup. In particular, [Sun23] proved the existence of “minimal” bubble in a differential geometric sense, finiteness of bubbles, and raised Conjecture 5.4 in loc.cit to expect a $2$-step algebraic construction of the underlying varieties of the minimal bubbles, and [dBS23] showed many examples and a refinement of the conjecture.
Motivated by loc.cit among others, this paper aims to provide a systematic algebro-geometric or non-archimedean approach to the problem.
The main contents of this paper give answer the above main question by solving large parts of the conjectures of [Sun23, dBS23] with some refinements. The following is a brief outlook of our statements but see the detailed statements in later sections.
Theorem 1.1.
Consider an arbitrary flat degeneration of pointed $n$-dimensional log terminal (hence normal) singularities $\pi\colon\mathcal{X}(\supset\sigma(\Delta))\to\Delta\ni 0$ (see Setup 1), which is strictly degenerating in a certain sense (see Theorem 2.4). Then, the following hold.
  1. ($\mathrm{i}$)
    (Theorem 2.4) There is a purely algebro-geometric $2$-step construction 222the following symbol $\rightsquigarrow$ means each step of the construction, and is not a map. of algebro-geometric version of minimum bubbles
    $(X\subset\mathcal{X})\rightsquigarrow\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow\mathcal{X}^{\prime\prime}_{\rm min,0}.$
    Here, the minimality morally means the mildness of singularity measured by the normalized volumes (at the base points) as we clarify later, as it is parallel to minimality of the diverging order of the rescaling constants $\{c_{t}\}_{t}$ (cf., also [Sun23] and Definition 2.1).
  2. ($\mathrm{ii}$)
    (Theorem 2.4, Theorem 2.8) The first step construction
    $X\subset\mathcal{X}\rightsquigarrow\mathcal{X}^{\prime}_{\rm min,0}$
    is uniquely (canonically) characterized, and coincides with more analytic construction following [Sun23, §4], which is called algebraic minimal bubble in [Sun23, §5.1, Conjecture 5.4].
  3. ($\mathrm{iii}$)
    (Theorem 2.6) Even at the total space level, the $1$-st step process $\mathcal{X}_{t}(t\to 0)\rightsquigarrow\mathcal{X}^{\prime}_{\rm min,0}$ can be achieved as a canonical birational modification of the original family, after a possibly non-finite type 333as far as we wants canonicity of the process base change i.e., a certain Novikov type ring spectrum (or an affine toric variety or even $\Delta$, if we do not mind non-canonicity of the base). We refer to these obtained total spaces as $\mathcal{X}^{\prime}_{\rm min}$ and $\mathcal{X}^{\prime\prime}_{\rm min}$.
[…]
To make sense of and prove the above ($\mathrm{iii}$), we use [Od24b] in which some algebraic (partly stack-theoretic) generalities to treat degenerations in “irrational direction” with higher dimensional algebraic base is prepared. We also compare our construction with more analytic and differential geometric constructions in [Sun23].
Since our construction is fairly general and look abstract first, although it is done after various concrete calculations as experiments, let us review a simple example after [dBS23] here. In this case, we only need a single process rather than the $2$-step processes i.e., $\mathcal{X}^{\prime}_{\rm min}=\mathcal{X}^{\prime\prime}_{\rm min}$, and they are achieved only after a finite base change of $\Delta$.
Example 1.2 (cf., [BKN89, Ban90] and [dBS23, §3.3]).
Consider a degeneration of pointed affine smooth surfaces to a $A_{3}$-singular surface as $\mathcal{X}\subset\Delta_{t}\times\mathbb{C}^{3}_{x,y,z}$ cut out by the equation
(1) $xy=z(z-t)(z-t^{2})(z-t^{3})$
whose base points are given by the 00-section $\sigma$. Here, the subindices $x,y,z,t$ are coordinates.
We can consider a family of hyperKähler metrics on each $t$ by the Gibbons-Hawking ansatz with the base $\mathbb{R}^{3}$ and monopole points $x_{1},x_{2},x_{3},x_{4}$ and harmonic functions of the form $\frac{1}{2}\sum_{i}\frac{1}{|x-x_{i}|}$. In this case, [dBS23, §3.3] describes how it bubbles at $t\to 0$ around the origin, and coincides with the algebraic procedure Theorem 1.1, and is explicit as follows. (Indeed, our general construction is originally obtained as a generalization of their examples).
Firstly, we consider the weighted blow up of $(0,0,0,0)\in\mathcal{X}$ with weights $(2,2,1,1)$ and removal of the strict transform of the central fiber $\mathcal{X}_{0}$, which gives $\mathcal{X}^{\prime}_{\rm min}=\mathcal{X}^{\prime\prime}_{\rm min}$ with the equation
(2) $xy=z(z-1)(z-t)(z-t^{2}),$
after replacing $\frac{x}{t^{2}}$ as $x$, $\frac{y}{t^{2}}$ as $y$, $\frac{z}{t}$ as $z$ while keeping $t$. This is a degeneration to $xy=z^{3}(z-1)$ when $t\to 0$, with $A_{2}$-singularity at $(0,0,0,0)\in\Delta_{t}\times\mathbb{C}_{x,y,z}^{3}$. This admits a ALE hyperKähler orbifold metric, again obtained via the Gibbons-Hawking ansatz over $\mathbb{R}^{3}$ (cf., [dBS23, §3.3]).
Then after the base change with respect to the double covering $\Delta_{\sqrt{t}}\to\Delta_{t}$, we replace $x,y,z,t$ by $\frac{x}{t^{1.5}},\frac{y}{t^{1.5}},\frac{z}{t},t^{0.5}$ and still denote it by $x,y,z,t$ for simplicity of notation. Then, the weighted blow up of the base changed family with weights $(3,3,2,2)$ (again with the strict transform of the central fiber removed) is defined by
(3) $xy$ $=z(t^{2}z-1)(z-1)(z-t^{2})$
(4) $=(t^{2}z-1)(z-1)\cdot z(z-t^{2}),$
which is a degeneration to a ALE hyperKähler orbifold bubbling with a $A_{1}$-singularity where the metric is infinitesimally $\mathbb{C}^{2}/\pm$.
As the final step, we again do a weighted blow up with weights $(2,2,2,1)$ for the coordinates $(x,y,z,t)$ and remove the strict transform of the central fiber. Then, by replacing the coorindates by new $(\frac{x}{t^{2}},\frac{y}{t^{2}},\frac{z}{t^{2}},t)$ which we denote again by $(x,y,z,t)$, we obtain the new equation
(5) $xy=(t^{4}z-1)(t^{2}z-1)\cdot z(z-1),$
inside $\Delta_{t}\times\mathbb{C}^{3}_{x,y,z}$. Note that this is finally a smooth family and we have an asymptotically conical hyperKähler metric of Eguchi-Hanson type at the central fiber.
As we clarify, in general, we need more complicated processes due to possible irregularity of the corresponding Sasaki-Einstein link i.e., irrationality of its Reeb vector field.
We make this kind of explicit study as Example 1.2 via coordinates more generally. That is, after the proof of Theorem 1.1 ($=$Theorems 2.4, 2.8), we propose more explicit ways of constructions of the candidates of bubbles, via coordinates/equations and valuations respectively in §2.4, §2.5. We conjecture that they coincide with the construction of Theorem 2.4. §2.2.3 and §2.3 may be of own independent interest, which are respectively the systematic treatment of negative weight deformation originated in [CH22, SJZ23] and an analogue of bubbling Calabi-Yau metric in the context of Kähler-Ricci flow. (The latter gives a yet another approach to the properness of K-moduli of Fano varieties cf., Theorem 2.21). Further, in §3, we discuss more on the structure on the “bubble tree” formed by all possible bubbles.
From some personal perspectives, this paper is also motivated by hope to deepen our understanding of the projective (resp., quasi-projective) K-moduli of Fano varieties (resp., Calabi-Yau varieties) in the sense of [Od10, 5.2], after recent completion of the construction (cf., e.g., [BL21, BHLLX21, LXZ22] resp., [Bir22, Od21]). Recall that those K-moduli were constructed and understood before loc.cit equivalently as Gromov-Hausdorff (partial) compactifications from more differential geometric viewpoints (cf., e.g., [OSS16, SSY16, Od15a, LWX19, Bir22, Od21]), following the breakthrough work of [DS14] and the aforementioned bubblings give a deepening of our understanding on it.
2. Minimal bubble
2.1. Definitions and Review
We work in Setup 1. The main objective of this section is to give an algebraic approach to the “minimum bubble” in the sense of [Sun23, dBS23], whereas the minimality roughly refers to that of the order of rescaling parameters $c_{t}$ for $t\to 0$. We review the definition for convenience.
Definition 2.1 ([Sun23, Definition 1.9]).
In the context of Setup 1, its minimal bubble refers to the pointed Gromov-Hausdorff limits of $(\sigma(t)\in\mathcal{X}_{t},c_{t}^{2}g_{t})$ for certain diverging sequence $c_{t}>0$, which is not a Riemannian metric cone, but whose tangent cone at infinity ([DS17, §3.4]) is the metric local tangent cone at $\sigma(0)\in X$, which we denote as $C_{\sigma(0)}(X)$.
Firstly, we discuss several a priori different algebraic approaches to it constructions, in different subsections, and will discuss the conjectural equivalence later. Since we crucially use the theory of local normalized volume of log terminal singularities due to [Li19, Liu18] (cf., also [Od15b]), with its origin in [DS17], we briefly review the definition and properties we use here. Its smooth case can be traced back to much earlier [Mus02, dFEM04] (and also related [ELS03]) and more detailed review are available at [LLX20], [Od24a, §2] for instance.
Definition 2.2 ([ELS03]).
[…]
Then, the normalized volume $\widehat{{\rm vol}}_{x}(v)$ is defined as $A_{X}(v)^{n}{\rm vol}_{x}(v)$ by [Li19], where $A_{X}(v)$ is the log discrepancy in the generalized sense by [JM12]. Then the following is known by now. We do not make any substantial contribution in this paper but rather use in this paper. See more references in [LLX20], [Od24a, §2].
Theorem 2.3 (cf., [DS17, Li19, Liu18, Blu18, HS17, LX18, LWX21, LLX20, XZ21, Od24a]).
For above $x\in X$, there is a unique quasi-monomial valuation $v=v_{X}$ up to the $\mathbb{R}_{>0}$-action, which minimizes the normalized volume $\widehat{\operatorname{vol}}(-)$.
  1. ($\mathrm{i}$)
    Then, the valuation $v_{X}$ coincides with the vanishing order function in [DS17, §3.2] for the local Kähler metric in the setup of loc.cit. We denote the rational rank of $v_{X}$ as $r$. Further, $\widehat{\rm vol}(v_{X})$ has a differential geometric meaning as the volume density of the link of the local metric tangent cone. We denote the minimized value $\widehat{\rm vol}(v_{X})$ as $\widehat{\rm vol}(x\in X)$ or $\widehat{\rm vol}_{X}(x)$, in this paper.
  2. ($\mathrm{ii}$)
    From more algebraic viewpoint, the graded ring ${\rm gr}_{v}(\mathcal{O}_{X,x}):=\oplus_{s\in{\rm Im}(v_{X})}(\{f\in\mathcal{O}_{X,x}\mid v(f)\geq s\}/\{f\in\mathcal{O}_{X,x}\mid v(f)>s\})$ is a finite type $\mathbbm{k}$-algebra on which there is a natural action of the algebraic $\mathbbm{k}$-torus $T\simeq\mathbb{G}_{m}^{r}$ on a K-semistable Fano cone $W:=\mathop{\mathrm{Spec}}\nolimits({\rm gr}_{v}(\mathcal{O}_{X,x})$.
  3. ($\mathrm{iii}$)
    For a certain rational polyhedral cone $\tau$ of the dual of the groupification of ${\rm Im}(v_{X})$, there is a positive weight degeneration of $X$ to $W$, over $U_{\tau}$ which is a $\mathbb{Q}$-Gorenstein family.
  4. ($\mathrm{iv}$)
    There is a (non-canonical) affine test configuration of $W$ to a unique K-polystable Fano cone $C$ with a (weak) Ricci-flat Kähler cone metric. In the setup of [DS17, §2-§3.3], it recovers the local metric tangent cone of local Kähler metric (as conjectured in loc.cit).
We note that the last $\mathbb{Q}$-Gorensteinness assertion of the positive weight degeneration in ($\mathrm{iii}$) seems missing in literature but can be confirmed as follows. As in [Od24b, §2], take homogeneous generating system $\overline{z}_{1},\cdots,\overline{z}_{l}$ of ${\rm gr}_{v}(\mathcal{O}_{X,x})$ with degrees $w_{1},\cdots,w_{l}$, and lift to $\Gamma(\mathcal{O}_{X})$ as $\overline{z}_{1},\cdots,\overline{z}_{l}$ which we suppose to give an embedding $W\hookrightarrow\mathbb{A}_{\mathbbm{k}}^{l}$ (otherwise add more $\overline{z}_{i}$s). Then we take rational approximation $\xi^{\prime}$ of $\xi:=(w_{1},\cdots,w_{l})$ and corresponding positive weight test configuration as $W_{\xi^{\prime}}$ over $\mathbb{A}^{1}$. For any $\xi^{\prime}$ close enough to $\xi$, we see that the central fiber is again $W$ just with a perturbed vector field induced on it. Note that there is a natural relative compactification $W_{\xi^{\prime}}\subset\overline{W_{\xi^{\prime}}}\to\mathbb{A}^{1}$ by adding the quotient by $\mathbb{G}_{m}$ for $\xi^{\prime}$-direction, which is automatically a $\mathbb{Q}$-Gorenstein family by the arguments of [OSS16, (proof of) Lemma 2.4]. Since this holds for any $\xi^{\prime}$ close enough to $\xi$, the map from the toric base $U_{\tau}$ of the positive weight degeneration is in the $\mathbb{Q}$-Gorenstein deformation component so that we obtain the assertion.
2.2. Re-construction via moduli-theoretic arguments
In this subsection, we use theorems in [Od24b] after [Lan75, HL14, AHLH23] to algebraic construction of the minimal bubble and discuss it. We also discuss reformulation of negative deformations in the sense of [CH22], after [Od24b].
2.2.1. Setup and algebraic construction
Setup 2.
In the context of Setup 1 (a) (resp., Setup 1 (b)), as Example 2.8 in [Od24b], denote by $v_{X}$ the valuation of $\mathcal{O}_{X,x}$ whose center is $x$ and minimizes the normalized volume $\widehat{\operatorname{vol}}(-)$ ([DS17, Li19, Blu18, XZ21]).
We set the groupification of ${\rm Im}(v_{X})$ as $N\simeq\mathbb{Z}^{r}$, $M:=N^{*}:={\rm Hom}_{\mathbb{Z}}(N,\mathbb{Z})$ denotes the dual lattice, and set $T:=N\otimes\mathbb{G}_{m}\simeq\mathbb{G}_{m}^{r}$, where $r$ is the $\mathbb{Q}$-rank of $v_{X}$.
For some technical reasons of our logics, i.e., to use Kollár’s hull and husk construction ([Kol08]) and also for precision of “good” family or singular limit of (weak) Kähler metrics later, we will henceforth sometimes assume the family $\pi$ can be compactified to or comes from a family of projective varieties. Note that it is also assumed in [DS14] (and [DS17]). We believe such assumption above will be unnecessary in the end (see Conjecture 3.1 ($\mathrm{iii}$)). That is, we temporary assume:
Setup 3.
In the contexts of Setup 1 $\land$ 2, we further assume $\pi\colon\mathcal{X}\to\Delta$ is compactifiable to a polarized projective family $\overline{\pi}\colon\overline{\mathcal{X}}\to\Delta$ such that $\overline{\mathcal{X}}_{t}=\overline{\pi}^{-1}(t)$ are all log terminal for any $t\in\Delta$ and $K_{\overline{\mathcal{X}}}$ is again $\mathbb{Q}$-Cartier.
The following main theorems 2.4, 2.6, 2.8 aim at algebraic reconstruction and generalization of the bubbles via metrics in [DS17, Sun23]. The first version below (Theorem 2.4) avoids the new terminology of [Od24b] for the readers convenience but the next one (Theorem 2.6), which uses loc.cit, will be more general and stronger.
Theorem 2.4 (Algebro-geometric construction of (minimal) bubbles).
We work under the Setup 1 (a) (resp., 1 (b)) $\land$ 2 $\land$ 3, and we further assume that the singularity of $\sigma(t)\in X_{t}$ strictly degenerates in the sense that
$\widehat{\operatorname{vol}}_{\mathcal{X}_{0}}(\sigma(0))<\widehat{\operatorname{vol}}_{\mathcal{X}_{t}}(\sigma(t))$
for any $t\neq 0$. Here $\widehat{\operatorname{vol}}(-)$ denotes the normalized volume 444also called under different names in the literature such as volume density ([SS17, HS17]) up to a constant, or local volume in [XZ24] (although different from the local volume of [Ful11]) in the sense of [Li19] (as its developments are also surveyed at [LLX20], [Od24a, §2]).
Then, possibly after replacing the base $\Delta\ni 0$ of $\pi$ by its finite base change, there exist a finite ordered set of pairs of birational transforms $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}\to\Delta$ and $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}\to\Delta$ which are both isomorphism over $\Delta\setminus\{0\}$ such that all the following conditions hold. We denote the projection $\mathcal{X}^{\prime}\to\Delta$ as $\pi^{\prime}$, and $\mathcal{X}^{\prime\prime}\to\Delta$ as $\pi^{\prime\prime}$ respectively.
Here are the conditions:
  1. ($\mathrm{i}$)
    $\pi^{\prime}\colon\mathcal{X}^{\prime}\to\Delta$ and $\pi^{\prime\prime}\colon\mathcal{X}^{\prime\prime}\to\Delta$ are affine faithfully flat morphisms and again (relatively/fiberwise) log terminal and they are both $\mathbb{Q}$-Gorenstein families in the sense that $K_{\mathcal{X}^{\prime}}$ and $K_{\mathcal{X}^{\prime\prime}}$ are both $\mathbb{Q}$-Cartier.
  2. ($\mathrm{ii}$)
    $\sigma|_{(\mathcal{X}\setminus\{0\})}$ extends to a holomorphic section $\Delta\to\mathcal{X}^{\prime}$ which we denote by $\sigma^{\prime}$ (resp., a holomorphic section $\Delta\to\mathcal{X}^{\prime\prime}$ which we denote by $\sigma^{\prime\prime}$).
  3. ($\mathrm{iii}$)
    Fix one of the pairs $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$. For the central fibers $\mathcal{X}^{\prime}_{0}$ of $\mathcal{X}$ (resp., $\mathcal{X}^{\prime\prime}_{0}$ of $\mathcal{X}^{\prime\prime}$), we have
    (*) $\widehat{\operatorname{vol}}_{\mathcal{X}^{\prime}_{0}}(\sigma^{\prime}(0))=\widehat{\operatorname{vol}}_{\mathcal{X}^{\prime\prime}_{0}}(\sigma^{\prime\prime}(0))>\widehat{\operatorname{vol}}_{\mathcal{X}_{0}}(\sigma(0)),$
[…]
  4. ($\mathrm{iv}$)
    The central fiber $\mathcal{X}^{\prime}_{0}:=\pi^{\prime-1}(0)$ of each $\mathcal{X}^{\prime}\ni\sigma^{\prime}_{0}(x)$, is an algebraic negative weight deformation ([Od24b, §2], after [CH22, Definition 1]) of a K-semistable $\mathbb{Q}$-Fano cone $(W,\xi)$ (in the sense of [CS18]) over $\mathbb{A}_{\mathbbm{k}}^{1}$ in a canonical manner, up to a finite base change, which we denote simply as $\mathcal{X}^{\prime}_{0}\rightsquigarrow W$. In particular, $\mathcal{X}^{\prime}_{0}:=\pi^{\prime-1}(0)$ is K-semistable in the sense of [Sun23, §5]. The total space $\mathcal{X}^{\prime}_{\rm min}$ is achieved as an elementary modification of $\mathcal{X}$ in the sense of [HL14, AHLH23].
  5. ($\mathrm{v}$)
[…]
  6. ($\mathrm{vi}$)
    The central fiber $\mathcal{X}^{\prime}_{{\rm min},0}$ of $\mathcal{X}^{\prime}_{\rm min}$ which satisfies all the above conditions is unique (cf., also Theorem 2.8 for the comparison with [Sun23]).
  7. ($\mathrm{vii}$)
    The central fiber $\mathcal{X}^{\prime\prime}_{0}:=\pi^{\prime-1}(0)$ of $\mathcal{X}^{\prime\prime}$ is an algebraic negative weight deformation 777 (See also discussions in [SRZ23, §6], which is related to [CH22, 1.7], though the stated existence of negative valuation is a weaker statement). ([Od24b, §2], after [CH22]) of a K-polystable $\mathbb{Q}$-Fano cone $(C,\xi_{C})$ in the sense of [CS18, CS19]. In particular, $\mathcal{X}^{\prime\prime}_{0}:=\pi^{\prime\prime-1}(0)$ (together with the corresponding negative valuation) is K-polystable in the sense of [Sun23, §5].
  8. ($\mathrm{viii}$)
[…]
    Further, the isotrivial degeneration is again a negative weight deformation999 (See the footnote 6 above). in the sense of [CH22, Definition 1.7 and Appendix I]. In particular, $\mathcal{X}^{\prime\prime}_{0}:=\pi^{\prime\prime-1}(0)$ (together with the corresponding negative valuation) is K-polystable in the sense of [Sun23, §5].
  9. ($\mathrm{ix}$)
    For each of the pairs $(\mathcal{X}^{\prime},\mathcal{X}^{\prime\prime})$, there is an affine $\mathbb{Q}$-Gorenstein test configuration of $\mathcal{X}_{0}^{\prime}$ degenerating to $\mathcal{X}_{0}^{\prime\prime}$. If $n=2$, this is the trivial degeneration i.e., $\mathcal{X}^{\prime}=\mathcal{X}^{\prime\prime}$.
  10. ($\mathrm{x}$)
    If the rational rank $r$ of $v_{X}$ is $1$, $\mathcal{X}^{\prime}_{\rm min,0}$ gives a unique divisorial valuation of $\mathcal{X}$ centered at $x$, which we call minimal critical (divisorial) valuation.
  11. ($\mathrm{xi}$)
    If we fix a test configuration of $W=\mathop{\mathrm{Spec}}\nolimits(gr_{v_{X}}(\mathcal{O}_{X,x}))\rightsquigarrow C_{\sigma(0)}(X)$, which we will denote as $\mathcal{T}^{\prime\prime}$ in the proof, the construction of $\mathcal{X}^{\prime\prime}_{\rm min}$ from $\mathcal{X}^{\prime}_{\rm min}$ is also canonical. If $n=2$, then this process is trivial i.e., $W=C_{\sigma(0)}(X)$ (as Fano cones).
Below, we use the framework of [Od24b] more to strengthen above theorem 2.4, following the same circle of ideas. (For the readers who prefer the naive more classical Setup 1, one can skip or only skim through the following Definition 2.5, Theorem 2.6.) In particular, the total spaces $\mathcal{X}^{\prime}_{\rm min}$ can be also taken as a canonical object, and we can treat much more general degeneration families. We begin with the following additional natural terminology, to compare with Setup 1 and partially review [Od24b].
Definition 2.5 (Restriction to one parameter families).
For a field $\mathbbm{k}$ and a submonoid $\Gamma_{\geq 0}\subset\mathbb{R}_{\geq 0}$, we consider N-ring $\mathbbm{k}[[\Gamma_{\geq 0}]]$ (see [Od24b, §2]). For a faithfully flat finite type morphism $\tilde{\pi}\colon\widetilde{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$, we call another faithfully flat finite type morphism $\mathcal{X}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$ is its restriction if there is a finitely generated submonoid $\overline{\Gamma}_{\geq 0}\subset\Gamma_{\geq 0}$ and a monoid surjective morphism $\varphi\colon\Gamma_{\geq 0}\twoheadrightarrow\mathbb{Z}_{\geq 0}$ such that
  • $\tilde{\pi}\colon\widetilde{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$ descends to $\overline{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\overline{\Gamma}_{\geq 0}]]$,
  • the base change of $\overline{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$ to $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$, through the $\mathbbm{k}$-algebra homomorphism $\mathbbm{k}[[\mathbb{Z}_{\geq 0}]]\to\mathbbm{k}[[\overline{\Gamma}_{\geq 0}]]$ induced by $\varphi$, is isomorphic to $\mathcal{X}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$.
We call the maps $\varphi$ or its associated $\mathbbm{k}$-algebra homomorphism $\mathbbm{k}[[\mathbb{Z}_{\geq 0}]]\to\mathbbm{k}[[\overline{\Gamma}_{\geq 0}]]$ a restriction map.
Theorem 2.6 (Algebro-geometric construction of (minimal) bubbles - ver2: with canonical base changes).
Following the setup of the above Definition 2.5, we also fix a section $\sigma\colon\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]\to\widetilde{\mathcal{X}}$ as setup 2 and further assume
  • $\mathbbm{k}$ has characteristic 00,
  • $\tilde{\pi}\colon\widetilde{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$ is also affine,
  • $K_{\widetilde{\mathcal{X}}/\mathbbm{k}[[\Gamma_{\geq 0}]]}$ is $\mathbb{Q}$-Cartier and the special fiber $\widetilde{\mathcal{X}}_{0}$ is log terminal (we call this condition (being) log terminal morphism following the literature)
  • strictly degenerating in the sense that $\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}_{0}}(\sigma(0))<\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}_{\eta}}(\sigma(\eta)),$ where $\eta(\text{resp., }0)\in\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$ is the generic point (resp., the closed point) and $\widetilde{\mathcal{X}}_{0}$ (resp., $\widetilde{\mathcal{X}}_{\eta}$) means its fiber.
  • $\tilde{\pi}$ has a restriction $\pi\colon\mathcal{X}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$ (which always holds for finite type $\tilde{\pi}$, and we fix the restriction map $\varphi$) in the sense of Definition 2.5.
Then, there is an iterative procedure of modifying $\tilde{\pi}$, as follows:
  1. (a)
    a finite (see Theorem 2.12) sequence of submonoids of $\mathbb{R}_{\geq 0}$ as
    $\Gamma_{\geq 0}$ $\subset\Gamma_{1,\geq 0}^{\prime}\subset\Gamma_{1,\geq 0}^{\prime\prime}$
    $\subset\Gamma_{2,\geq 0}^{\prime}\subset\Gamma_{2,\geq 0}^{\prime\prime}$
    $\subset\cdots\cdots\subset\mathbb{R}_{\geq 0}.$
  2. (b)
    vertical birational transforms of the base change of $\tilde{\pi}$ as faithfully flat affine log-terminal morphisms $\widetilde{\mathcal{X}}_{i}^{\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{i,\geq 0}^{\prime}]]$ and $\widetilde{\mathcal{X}}_{i}^{\prime\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{i,\geq 0}^{\prime\prime}]]$ for each $i=1,2,\cdots$,
  3. (c)
    their sections $\sigma_{i}^{\prime}\colon\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{i,\geq 0}^{\prime}]]\to\widetilde{\mathcal{X}}_{i}^{\prime}$,
  4. (d)
    their sections $\sigma_{i}^{\prime\prime}\colon\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{i,\geq 0}^{\prime\prime}]]\to\widetilde{\mathcal{X}}_{i}^{\prime\prime}$,
(the generic fibers with their sections are compatible with respect to the base change for (a)) such that
  • the set of pairs of restrictions of $\widetilde{\mathcal{X}}_{i}^{\prime}$, $\widetilde{\mathcal{X}}_{i}^{\prime\prime}$ (for same $i$), with respect to restriction maps which extend $\varphi$ compatibly, give the set of $(\mathcal{X}^{\prime},\mathcal{X}^{\prime\prime})$ in the previous Theorem 2.4 (in particular, $i=1$ case gives $(\mathcal{X}^{\prime}_{\rm min},\mathcal{X}^{\prime\prime}_{\rm min})$.
  • For each $i$,
    $\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}_{0}}(\sigma(0))$ $<\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}^{\prime}_{i,0}}(\sigma^{\prime}_{i}(0))$
    $=\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}^{\prime\prime}_{i,0}}(\sigma^{\prime\prime}_{i}(0))(=:c_{i})$
    $\leq\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}_{\eta}}(\sigma(\eta)),$
    and $c_{i}$ are monotonely increasing for $i=1,2,\cdots$, where $\widetilde{\mathcal{X}}^{\prime}_{i,0}$ and $\widetilde{\mathcal{X}}^{\prime\prime}_{i,0}$ are the central fibers of $\widetilde{\mathcal{X}}^{\prime}_{i,0}$ and $\widetilde{\mathcal{X}}^{\prime}_{i,0}$ respectively.
  • (Canonicity) The first step construction of $\Gamma_{1,\geq 0}^{\prime}$, $\widetilde{\mathcal{X}}_{1}^{\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{1,\geq 0}^{\prime}]]$ and its section $\sigma_{1}^{\prime}$ is canonical i.e., only need the data $\tilde{\pi}$.
Remark 2.7 (Locality and logarithmic generalization).
From the following proof of Theorem 2.4, 2.6 combined with the uniqueness statements in [Od24b, 3.8 (2) (also (1)), 3.15 (2)], it follows that the germs of $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ around the images of $\sigma^{\prime}$ (resp., $\sigma^{\prime\prime}$) is determined only by the germ of the image of $\sigma\colon\Delta\to\mathcal{X}$. Generalization of the above results to the so-called log pair setup i.e., family of kawamata-log-terminal log pairs $\{(X_{t},D_{t})\}_{t\in\Delta}$ (cf., e.g., [BL21, §2.4] for review of the definition) and corresponding weak conical singular Kähler metrics, is also almost verbatim which we omit here to avoid technical complication.
As a link to differential geometric works in [Sun23, dBS23], we first prove the following, which confirms [Sun23, Conjecture 5.4 (i)] and also makes a progress on Conjecture 5.4 (ii) of op.cit (cf., also [dBS23, §4.1]).
Theorem 2.8 (Partial identification with analytic construction of [Sun23]).
We work under the Setup 1 (a), 2, 3, and we further assume that
  • $\pi$-ample $\mathcal{L}$ is put on $\overline{\mathcal{X}}$ so that $g_{t}$ is restriction of the unique Kähler-Einstein metric on $\mathcal{X}_{t}$ whose Kähler class is $2\pi c_{1}(\mathcal{L}_{t}:=\mathcal{L}|_{\mathcal{X}_{t}})\in H^{1,1}(\overline{\mathcal{X}}_{t},\mathbb{R})$.
  • $(\mathcal{X}_{0},\mathcal{L}_{0},g_{0})$ is the polarized limit space of $(\overline{\mathcal{X}}_{t},\mathcal{L}_{t},g_{t})$ for $t\to 0$ in the sense of [DS14] (and [DS17]),
to make [DS17, Sun23] works. Then,
  1. ($\mathrm{i}$)
    our algebraic minimal bubble $\mathcal{X}^{\prime}_{\rm min,0}$ can be described by the notations in the proof of [Sun23, Therorem 4.3] after that discussions, and is unique for each $\pi$.
  2. ($\mathrm{ii}$)
    our $\mathcal{X}^{\prime\prime}_{\rm min,0}$ satisfies the conditions (a), (b) of [Sun23, Conjecture 5.4].
Question 2.9 (Yau-Tian-Donaldson type problem for bubbling).
Does our $\mathcal{X}^{\prime\prime}_{\rm min,0}$ for some $\mathcal{T}^{\prime\prime}$ in Theorem 2.4 coincide with the minimum bubble in the differential geometric sense of [Sun23, Definition 1.9] (Definition 2.1), obtained in [Sun23, Theorem 1.10]?
We confirmed the above question affirmatively in the following $1$ or $2$-dimensional classical cases considered in [dBS23]:
  • conical singular curves (“log curves”), considered in loc.cit §2.3,
  • $2$-dimensional $A_{k}$-singularities with the standard metric constructed by the Gibbons-Hawking ansatz, considered in loc.cit §3.3,
  • and higher dimensional $A_{1}$-singularity case, considered in loc.cit §4.2.
We omit the details here and write the proofs elsewhere (hopefully for more generality).
proof of Theorem 2.4.
First we discuss the construction of $\mathcal{X}^{\prime}_{\rm min}$. For simplicity, note that we can and do assume $\sigma(t)$ is the origin 0=($\vec{0}=(0,\cdots,0)$ for any $t$.
By the extended Rees construction (cf., e.g., [Mat89, Chapter 5, section 4], [Tei03, §2]), we can and do assume $v_{X}$ is associated to weights $w_{1},\cdots,w_{l}\in\mathbb{R}_{>0}$ on each coordinate of $\mathbb{A}_{\mathbbm{k}}^{l}$. See some complementary arguments in [Od24b, Example 2.9]. Below, we consider degree of polynomials on $\mathbb{A}_{k}^{l}$ with respect to this set of weights $\xi:=(w_{1},\cdots,w_{l})$. From the construction, we suppose the defining ideal of $W\subset\mathbb{A}_{\mathbbm{k}}^{l}$ is generated by the non-constant ($\xi$-)quasi-homogeneous elements $f_{1},\cdots f_{a}$ with the degrees (again with respect to $\xi$) bounded above by a positive constant $C>0$, and also that $\mathcal{X}_{t}$ for small enough $t$ is also defined by equations of degrees bounded above by $C>0$.
Now we consider the set of all possible polynomials $g_{1},\cdots,g_{a}$ on $\mathbb{A}_{\mathbbm{k}}^{l}$ of the degrees at most by $C$, without constant terms, and consider the corresponding universal family $\mathcal{U}\twoheadrightarrow\mathbb{A}_{\mathbbm{k}}^{m}$ for $m\gg 0$. Then we again apply the flattening stratification and universal hull construction [Kol08, 1, 2]. We then take its $T$-invariant strata $M$ which contains the base disk of $\mathcal{X}\to\Delta$ and the base of the positive weight degeneration $X$ degenerating to $W$ of Example 2.9 of [Od24b] after [Li19], and an isotrivial degeneration of $W$ to the metric tangent cone $C_{0}(X)$ (cf., [DS17, LWX21]). One can also take $\mathcal{T}$ a (non-canonical) affine test configuration induced by an algebraic subtorus $T^{\prime}$ of $T$ with the rank $1$ (the symbol $\mathcal{T}^{\prime}$ is reserved for different family to appear later as it is more natural notation for it), as a subfamily of the whole family $\mathcal{X}_{\tau}\to U_{\tau}$ of [Od24b, Example 2.9]. For $r>1$ case, this $T^{\prime}$ is clearly non-unique as a rational approximation of the algebraic positive degeneration $X\rightsquigarrow W$ ([LX18, §3], [Od24b, §2]). We denote the base change of $\mathcal{U}\to\mathbb{A}_{k}^{m}$ to $M$ as $\mathcal{V}\to M$. Further, we can and do replace the chart $M$ of $\mathcal{M}:=[M/T]$ by its equivariant resolution of singularities, and birationally modify $\mathcal{M}$ accordingly, which we still denote by the same letters $M$ and $\mathcal{M}$ for the simplicity of notation.
From our construction, the resulting $\varphi\colon[\mathcal{V}/T]\to\mathcal{M}$ is a $T$-equivariant $\mathbb{Q}$-Gorenstein family over an Artin stack which is smooth (as stack). By the openness of log terminality, we can further assume it is a log terminal morphism and all the fibers contain the origin 00.
Now we set $c:=\widehat{\operatorname{vol}}(X,0)$, the (minimized) normalized volume ([DS17, Li19], cf., also [Od24b, §2] for review) 101010recently it is also called local volume [XZ24] of the klt singularity $0\in X$ and take an open subscheme of $M$ whose $\mathbbm{k}$-valued points are $\{s\in M(\mathbbm{k})\mid\widehat{\operatorname{vol}}(0,\varphi^{-1}(s))\geq c\},$ which is indeed Zariski open by [BL21, Theorem 1], [Xu20, 1.3]. Then, we consider
$\{s\in M(\mathbbm{k})\mid\widehat{\operatorname{vol}}(0,\varphi^{-1}(s))=c\},$
which gives a Zariski closed subset $M_{c}$ of $M$ again by [BL21, Theorem 1]. We also consider reduced subscheme structure on $M_{c}$. We denote $\mathcal{M}_{c}:=[M_{c}/T]$.
As $\mathcal{M}_{c}\neq\emptyset$, we consider $\varphi^{-1}(s)$ for $s\in\mathcal{M}_{c}(\mathbbm{k})$. The construction of $W$ for each such $\varphi^{-1}(s)$ gives a structure of canonical higher $\Theta$-strata in the sense of [Od24b] (following [HL14, AHLH23]), of the quotient stack $[M/T]$ supported on $[M_{c}/T]$. The first step of its proof is that, in general, if you have a log terminal affine morphism with base points $\tilde{\mathcal{X}}(\supset\sigma(S)=\{0\}\times S)\to S$ and the positive weight degenerations $\tilde{X}_{s}\ni\vec{0}\rightsquigarrow W_{s}\ni\vec{0}$ of the pointed klt affine variety $X_{s}\ni\vec{0}$ to its K-semistable cone $W_{s}$ can be taken uniformly, i.e., if there is a flat family over $S\times_{\mathbbm{k}}U_{\tau}$ with a connected algebraic $\mathbbm{k}$-scheme $S$ and an affine toric variety $U_{\tau}$ which corresponds to a rational polyhedral cone $\tau$, the corresponding minimal normalized volume is constant i.e., not depending on $s\in S$. This follows from the comparison with the leading coefficient of the index character (cf., [MSY08, Li19, CS18] and also [BB73, 4.1], [Od24a, §2]). Therefore, in particular, $\Delta\times U_{\tau}\dashrightarrow M$ which is induced by the positive weight deformation of $W$ can not be a morphism. Another technical result we use is a natural higher analogue of the $\Theta$-reductivity type statement as proved in [Che24, 1.1]. See the statements [Od24a, §3.3, Theorem 3.6] for a complementary discussion and the actual statement we use, from which it follows that $\{\mathcal{M}_{c}\}_{c}$ give a higher $\Theta$-stratification of $\mathcal{M}$, in the sense of [Od24b].
Hence, we can apply semistable reduction type theorem [Od24b, Theorem 3.8 (also cf., 1.1)] (which generalizes [Lan75, AHLH23]) and we obtain an Langton or [AHLH23, 3.35] type modification of the original morphism $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}\to[M/T]$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$) which corresponds to $\mathcal{X}$. We denote the obtained family as $\mathcal{X}^{\prime}_{\rm min,0}$ and going to confirm the required conditions in our Theorem 2.4. As noted in [Od24b], note that the family is non-canonical which corresponds to a choice of $T^{\prime}$ while the central fiber $\mathcal{X}^{\prime}_{\rm min,0}$ does not depend on it, from our construction (and [Od24b, §3]). Here, $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ denotes the discrete valuation ring which consists of the convergent power series of $t$ (for some positive unspecified radius). The obtained map from $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$), or its finite extension $\mathbb{C}[[t^{\frac{1}{d}}]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t^{\frac{1}{d}}]]$) for some $d\in\mathbb{Z}_{>0}$, gives the desired family $\mathcal{X}^{\prime}_{\rm min}\to\Delta$ from the construction. The K-semistability of $\mathcal{X}^{\prime}_{\rm min,0}$ in the sense of [Sun23, §5] follows from the construction as $W$ is a K-semistable log terminal cone ($\mathbb{Q}$-Fano cone) in the sense of [CS18, CS19]. Furthermore, from the positivity of the weights on the coordinates of $T^{\prime}$-action and the definition of elementary modification [AHLH23, §3], we conclude that the obtained test configuration $\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow W$ is a negative weight deformation in the sense of [CH22]111111just of rank $1$ i.e., with constant $\xi_{i}$ in their definition, [Od24b, §2]. Note that we do not know yet if any good choice of $T^{\prime}$ gives rise to a $\Theta$-strata in the original sense, and independence of $T^{\prime}$ is a priori unclear (without using [Od24b]) even if that is the case. These points are why we need the higher rank setup and arguments in the generalized form [Od24b, Theorem 3.8] rather than the original in [AHLH23, §6].
Now we are ready to discuss algebraic construction of $\mathcal{X}^{\prime\prime}_{\rm min}$ from $\mathcal{X}^{\prime}_{\rm min}$, which follows a similar line of ideas as those for $\mathcal{X}^{\prime}_{\rm min}$, but also with some differences. We take a (non-canonical) algebraic subtorus $T^{\prime\prime}$ of $T$, with the rank $1$, which degenerates $W$ to the K-polystable degeneration as in [LWX21, Theorem 1.2]. Denote the corresponding base i.e., the partial compactification of $T^{\prime\prime}$ as $\overline{T}^{\prime\prime}$ which is isomorphic to $\mathbb{A}^{1}$. Recall that the construction in loc.cit is far from canonical. We still take one of them and denote the corresponding test configuration as $\mathcal{T}^{\prime\prime}\to\overline{T}^{\prime\prime}$. Henceforth we can and do assume $\mathcal{T}^{\prime\prime}$ is also achieved inside $M$, by enlarging $M$ inside the original base $\mathbb{A}_{k}^{m}$ if necessary, and $\mathbb{C}^{l}$ (resp., $\mathbb{A}_{k}^{l}$) by retaking larger $l$, $\mathbb{C}^{l}$ (resp., $\mathbb{A}_{k}^{l}$) and the embedding $\mathcal{X}\hookrightarrow\mathbb{C}^{l}\times\Delta$ (resp., $\mathcal{X}\hookrightarrow\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$) is also realized in $M$ if necessary. Now we consider a couple of morphisms $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}\to M$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]\to M$) corresponding to $\mathcal{T}^{\prime}$ of degeneration of $X^{\prime}\rightsquigarrow W$, and corresponding to $\mathcal{T}^{\prime\prime}$ whose degeneration is of the form $W\rightsquigarrow C$.
We denote the base, with its associated moduli map, of the negative weight degeneration $\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow W$ by $\Delta^{\prime}\to M$. Then, consider the natural rational map from $m\colon\Delta^{\prime}\times\overline{T}^{\prime\prime}\to M$. If it extends to a morphism, one can take the image of $c\times\{0\}(\in\Delta^{\prime}\times\overline{T}^{\prime\prime})$ for $c\neq 0$ as $\mathcal{X}^{\prime\prime}_{\rm min,0}$. Otherwise, we apply [AHLH23, Theorem 6.3] (or [Od24b, Theorem 3.8]) to them, with respect to the simple $\Theta$-strata $[\overline{m(\Delta^{\prime}\times T^{\prime\prime})}/T^{\prime\prime}]$ supported at $m(0\in\Delta)$. Here, $\overline{m(\Delta^{\prime}\times T^{\prime\prime})}$ means the closure of the image $\overline{m(\Delta^{\prime}\times T^{\prime\prime})}$ of $m$. Then, it follows that possibly after passing to a finite ramified covering of the base $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$), we obtain a isotrivial family which degenerates $\mathcal{X}^{\prime}_{\rm min,0}$ to $\mathcal{X}^{\prime\prime}_{\rm min,0}\subset\mathbb{A}_{\mathbbm{k}}^{l}$. In this case, we can avoid the full use of [AHLH23, Od15b] by a more standard direct and canonical arguments using a toric surface as follows.
Note that $T^{\prime}$ and $T^{\prime\prime}$ are commutative so that they are both algebraic subgroups of an algebraic tori $T_{2}$ of rank $2$, which acts equivariantly on the family $\mathcal{U}\to\mathbb{A}_{\mathbbm{k}}^{m}$ and its base change $\mathcal{V}\to M$. Since the obtained elementary modification maps $(0,0)\in[\mathbb{A}^{2}/\mathbb{G}_{m}]={\rm ST}_{k[[t]]}$ (cf., [HL14, AHLH23]) to $[C\in\mathcal{M}]$ with the inertia group $T$, the elementary modification is represented by a $T_{2}$-equivariant morphism $f\colon\mathbb{A}_{\mathbbm{k}}^{2}\to M$. The normalization of its image in $M$ is a toric surface $T_{N\simeq\mathbb{Z}^{2}}\Sigma$ where the fan contains a ray $\mathbb{R}_{>0}(1,0)$ (resp., $\mathbb{R}_{>0}(0,1)$) which corresponds to $\mathcal{T}^{\prime}$ (resp., $\mathcal{T}^{\prime\prime}$) and its support ${\rm Supp}(\Sigma)=\mathbb{R}_{\geq 0}^{2}$ from the construction. Take the ray $\mathbb{R}_{>0}(s_{1},s_{2})(s_{i}\in\mathbb{Z}_{>0})$ which is adjacent to $\mathbb{R}_{>0}(1,0)$ and consider the one parameter subgroup $\lambda\colon\mathbb{G}_{m}\to T_{2}$. Then it gives a test configuration $\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow\mathcal{X}^{\prime\prime}_{\rm min,0}$ for some $X^{\prime\prime}\in S$ such that the inverse action of $\mathcal{T}^{\prime\prime}$ degenerates $\mathcal{X}^{\prime\prime}_{\rm min,0}$ to $C$ as desired.
Now we glue a family $\mathcal{X}^{\prime}_{\rm min,0}$ and $\pi_{X^{\prime}}$, by the same technique as [LWX21, §3], [Od22a, §3], to obtain the desired $\mathcal{X}^{\prime\prime}_{\rm min,0}\to\Delta$, as a vertical birational modification of $\mathcal{X}^{\prime}_{\rm min,0}\to\Delta$, which satisfies the conditions listed in the statements of our theorem. The only part which does not follow immediately, is the equality assertion of ($\mathrm{iii}$). It holds because of the condition satisfied by the “minimally” constructed $\mathcal{X}^{\prime}_{\rm min,0}\ni\sigma^{\prime}(0)$ i.e., that it achieves the possible minimum normalized volume, combined with the semicontinuity [BL21, Theorem 1], [Xu20, Theorem 1.3] of $\widehat{\rm vol}$ applied to $\pi_{X^{\prime}}$. Here we complete the construction of $\mathcal{X}^{\prime\prime}_{\rm min,0}$ from $\mathcal{X}^{\prime}_{\rm min,0}$. If $n=2$, since any log terminal singularities are quotient singularities so that K-semistable Fano cones are automatically K-polystable, we have $\mathcal{X}^{\prime\prime}_{\rm min,0}=\mathcal{X}^{\prime}_{\rm min,0}$ from our construction.
Finally, to obtain more bubbles $(\mathcal{X}^{\prime},\mathcal{X}^{\prime\prime})$, we replace the given $\mathcal{X}_{\rm min}$ by $\mathcal{X}^{\prime\prime}_{\rm min}$ and repeat the whole process above. We show this repetition terminates after finite times in Theorem 2.12. We complete the proof. ∎
Remark 2.10.
From the above proof, it easily follows that the analytic/formal germs of $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ around the images of $\sigma^{\prime}$ (resp., $\sigma^{\prime\prime}$) is determined only by the analytic germ of the image of $\sigma\colon\Delta\to\mathcal{X}$. Generalization of the above results to the so-called log pair setup i.e., family of kawamata-log-terminal log pairs $\{(X_{t},D_{t})\}_{t\in\Delta}$ (cf., e.g., [BL21, §2.4] for review of the definition) is also almost verbatim which we omit here to avoid technical complication.
Remark 2.11.
The above replacement $\mathcal{X}$ by $\mathcal{X}^{\prime}$ or $\mathcal{X}^{\prime\prime}$ or the existence of bubbles can first look in a sharp contrast to what happens in analogous global setup i.e., the separatedness of the moduli of log terminal projective Calabi-Yau varieties ([Od12, 4.2 (i), 4.3],  [^12] However, although 4.3 is fine, there is an obvious typo in its 4.2(i), mixing up notation with canonically polarized case: in the left hand side $-K_{\mathcal{X}/\mathcal{C}}$ should be read $\mathcal{L}$ and we assume K-stability (log terminality) of the fibers of $\pi$. [Bou14, 2.1]).
proof of Theorem2.6.
Our proof follows that of Theorem 2.4 closely. For $\pi^{-1}(0)=\tilde{\mathcal{X}}_{0}$, we apply the same arguments as the beginning of the proof of Theorem 2.12. In particular, we take the same log terminal family $\mathcal{U}\subset\mathbb{A}_{\mathbbm{k}}^{l}\times\mathbb{A}_{\mathbbm{k}}^{m}$, which restricts to $\mathcal{V}\to M$.
We set $\mathcal{M}:=[M/T]$ again and consider the map $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]\to M\to\mathcal{M}$ which corresponds to $\tilde{\pi}$. As in the previous proof of Theorem 2.4, take the algebraic torus $T$ for $W$, the semistable tangent cone of $X$ and a rational polyhedral cone $\tau$ in the cocharacter lattice $N:={\rm Hom}(\mathbb{G}_{m},T)$ with $\xi\in\tau$ such that the base change of $\mathcal{V}$ over $U_{\tau}$ gives a flat isotrivial degeneration (over possibly higher dimensional base) with the central fiber $W$. If there is a uniform family of such isotrivial degenerations $X\rightsquigarrow W$ over a connected base, then from the same reason as in the proof of Theorem 2.4, i.e., by the use of index character ([MSY08, CS18, CS19]), the normalized volume are constant. Similarly as before, combined with the uniqueness of the $\widehat{\rm vol}$-minimizing valuation, this gives the canonical structure of higher $\Theta_{\tau}$-strata on $M_{c}$ for $c:=\widehat{\rm vol}(X,0)$. Then, we again apply the higher $\Theta$-stable reduction theorem [Od24b, §3.2, Theorem 3.8] to construct $\Gamma_{1,\geq 0}^{\prime}$, $\widetilde{\mathcal{X}}_{1}^{\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{1,\geq 0}^{\prime}]]$ and its section $\sigma_{1}^{\prime}$. From the above discussions, this construction is completely canonical. The rest of the constructions are almost the same, so we omit the essential repetition. ∎
2.2.2. Comparison with the differential geometric construction
Now, we discuss comparison with the differential geometric construction in [DS14, Sun23] and prove Theorem 2.8.
proof of Theorem2.8.
We prove the item ($\mathrm{i}$) of Theorem 2.8. In this proof, we closely follow the arguments and notations of [Sun23] (cf., also [DS17]) by fitting it to our notation in Theorem 2.8. In the notation of loc.cit, suppose that his $X_{j}(j=1,2,\cdots)$ are taken as fibers of our $\pi$ i.e., $X_{j}=X_{s_{j}}$ for some sequence $s_{j}\in\Delta$ with $s_{j}\to 0$ for $j=1,2,\cdots$ and $X_{\infty}=\pi^{-1}(0)=X$. Our $\xi$ corresponds to the real $N\times N$ diagonal matrix $\Lambda={\rm diag}(2^{\zeta_{1}},\cdots,2^{\zeta_{N}})$ with $\zeta_{i}\in\mathbb{R}_{>0}$s acting on the ambient space $\mathbb{C}^{N}$ in his notation. We crucially use the local holomorphic functions $G^{i},S^{i}_{j},L^{i}_{j}$ (and $H^{i}$, $H^{i}_{j}$ with higher degrees with respect to $\xi$) on the ambient spaces $\mathbb{C}^{N}$ of $X_{j}$s ($i$ was denoted as $b$ in loc.cit), $X_{\infty}$ and the possible bubbles $Z$, the bounds of rescaling orders $\underline{k}$ and $\overline{k_{j}}$ as prepared in loc.cit. Then, we replace the sequence $\{j\}$ by its subsequence so that there is a sequence of real numbers $b_{j}(j=1,2,\cdots)$ with $b_{j}\to\infty$ obtain that
$W_{1}:$ $=\lim_{j\to\infty}\Lambda^{b_{j}}\cdot X_{j}$
Let us write $\operatorname{deg}_{\xi}$ for the degree with respect to $\xi$ as in [Sun23, §4]. Here, each $L^{i}_{\infty}$ is obtained as a rescaled limit of $L^{i}_{j}$ for $j\to\infty$. Indeed, it can be achieved as follows. For each $i$, consider the set $L(i)$ of monomials $\chi^{\vec{m}}$ on the ambient space $\mathbb{C}^{N}$ whose $\xi$-degree is less than that of $G^{i}$. Its order is finite as $\zeta_{1},\cdots,\zeta_{N}$ are all positive. We denote the coefficient of $\chi^{\vec{m}}$ in $L^{i}_{j}$ as $c^{i}_{j}(\vec{m})$, then of course we have $c^{i}_{j}(\vec{m})\to 0\hskip 11.38109pt(j\to\infty)$ as $L^{i}_{j}\to 0\hskip 11.38109pt(j\to\infty)$. Now, for each fixed $j$, we consider
$b_{j}:=\max_{i,\vec{m}\in L(i)}\dfrac{\log(|c^{i}_{j}(\vec{m})|)}{\log(2)\cdot(\operatorname{deg}_{\xi}(G^{i})-\operatorname{deg}_{\xi}(\chi^{\vec{m}}))}$
which exists by the finiteness of $\{i\}$ (the Noetherian property) and each $L(i)$. Hence, by passing to a subsequence of $j$, we can and do assume that the above maximum is attained by the same $i$ and $m$. Further, by the compactness of subset $\prod_{i}\{t\in\mathbb{C}\mid|t|\leq 1\}^{L(i)}$, as the set of coefficients of $2^{-\operatorname{deg}_{\xi}(G^{i})}(\Lambda^{b_{j}}\cdot L^{i}_{j})$, if we pass to a further subsequence, we can and do assume that for any $i$, $2^{-\operatorname{deg}_{\xi}(G^{i})}(\Lambda^{b_{j}}\cdot(G^{i}+L^{i}_{j}))$ converges when $j\to\infty$. We denote its limit as $G^{i}+c(i)L^{i}_{\infty}$ with some $c(i)\in\mathbb{C}$ for each $i$ and $\xi$-homogeneous polynomials $L^{i}_{\infty}$ with $\operatorname{deg}_{\xi}(L^{i}_{\infty})<\operatorname{deg}_{\xi}(G^{i}).$ Since $\Lambda^{\overline{k_{j}}}\cdot X_{s_{j}}\to W(j\to\infty)$ with the notation of [Sun23], for each i and $\vec{m}$, we have $2^{\overline{k_{j}}}|c^{i}_{j}(\vec{m})|\to 0$ for $j\to\infty$. On the other hand, for some $i$ and $\vec{m}$, we have $2^{b_{j}}|c^{i}_{j}(\vec{m})|\to 1$ for $j\to\infty$ by the construction. Therefore, we have $\lim_{j\to\infty}(b_{j}/\overline{k_{j}})=0$. Further, from the construction of $W_{1}$, it is clear that $(\Lambda^{-1})^{\mathbb{R}_{>0}}\cdot W_{1}$ is enhanced to be a test configuration degenerating to $W$. All these confirm the conditions for the uniqueness theorem [Od24b, Theorem 3.12 (ii)] so that we can apply it to conclude that $W_{1}$ is $\mathcal{X}^{\prime}_{{\rm min},0}$ in Theorem 2.4. Note that $\mathcal{X}^{\prime}_{{\rm min}}$ is the family level enhancement, which can be taken even canonical over a convergent N-type spectrum by [Od24b, Theorem 3.12]. Item (ii) is shown in Theorem 2.4 (but repeated for the comparison with [Sun23, Conjecture 5.4]). ∎
Now we discuss finiteness of bubbles, for fixed $\pi$, as it naturally arises in the previous theorem 2.4. A differential geometric counterpart is proven for the case when $\mathcal{X}_{t}(t\neq 0)$ are all smooth, compactifiable to metrized Kähler manifolds, in [Sun23, Corollary 1.12].
Theorem 2.12.
We fix a positive integer $n$. In the setup 1, 2, 3, we modify $\mathcal{X}$ by $\mathcal{X}^{\prime\prime}_{\rm min}$. Then this modification process does not continue for infinitely many times for $\operatorname{dim}(X)=n$. Hence, we algebraically obtain a sequence of (weighted) algebraic bubbles $Z_{0},\cdots,Z_{2k}$ similarly to [Sun23, Cor 1.12].
In particular, as the terminal object of this repetition, possibly after a finite base change of $\Delta$, there is an affine faithfully flat klt family $\mathcal{Y}\to\Delta$ which coincides with $\mathcal{X}\to\Delta$ over $\Delta\setminus\{0\}$, $Y=Z_{2k}$, and the section $\sigma$ to a section $\tau$ from $\Delta$ so that
(6) $\widehat{\operatorname{vol}}_{\mathcal{Y}_{0}}(\tau(0))$ $=\widehat{\operatorname{vol}}_{\mathcal{Y}_{t}}(\tau(t))$
(7) $=\widehat{\operatorname{vol}}_{\mathcal{X}_{t}}(\sigma(t)).$
for an arbitrary $0<|t|\ll 1$.
Proof.
The former half of Theorem 2.12 is a simple corollary of Theorem 2.4, that enhances its statements. The modification process $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}_{\rm min}$ terminates by [XZ24]. We take $\mathcal{Y}$ as the end object of the repetition of the process $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}$, towards deeper bubbles. Then the last equality (6) holds as otherwise we can repeat the process by Theorem 2.4, and the equality (7) is obvious from identification of $\mathcal{X}$ and $\mathcal{Y}$ outside fibers over $0\in\Delta$. ∎
Note that the last statement of the above Theorem 2.12 is a semistable reduction type theorem corresponding to virtual compactness of “moduli” of singularities with fixed normalized volumes. In particular, the following corollary to above is pointed out by M.Mauri. Our main reason of writing this is on our different method than usual resolution of singularities, rather than the result itself.
Corollary 2.13 (Semistable resolution).
Consider any affine pointed flat morphism $\pi\colon\mathcal{X}\to\Delta\ni 0$ of relative dimension $n$, with a section $\sigma\colon\Delta\to\mathcal{X}$ such that fibers $\pi^{-1}(s)=\mathcal{X}_{s}\ni\sigma(s)$ are all smooth points (resp., log terminal) for $s\neq 0$ (resp., $s=0$), over a field $\mathbbm{k}$ of characteristic 00, as in Setup 1.
Then there is a birational modification $\tilde{\pi}\colon\mathcal{Y}\twoheadrightarrow\Delta$ along $\mathcal{X}_{0}$, possibly after a finite base change of $\Delta$, such that $\sigma|_{\Delta\setminus 0}$ extends to whole $\Delta$ as $\sigma_{\mathcal{Y}}\colon\Delta\to\mathcal{Y}$, $\tilde{\pi}$ is smooth surjective around $\sigma_{\mathcal{Y}}(s)$ for any $s$ near 00.
If $n=2$, more generally, if $\pi$ is a log terminal morphism, one can construct $\mathcal{Y}$ with $\sigma_{\mathcal{Y}}$ which are étale locally trivial around $\sigma_{\mathcal{Y}}(s)$ around the neighborhood of 00.
Proof.
Suppose the last assertion is not satisfied for the original $\mathcal{X}$. Otherwise, the statements is obvious. Now we apply Theorem 2.4 and Theorem 2.12 to obtain the deepest bubbling family $\tilde{\pi}$. From (6) and (7) of Theorem 2.12, $\widehat{\rm vol}_{\mathcal{Y}_{0}}(\sigma_{\mathcal{Y}}(0))=n^{n}$. Then, [LX19, Theorem A.4] (see also implicit arguments in [HS17, SS17, LX18]) shows that $\sigma_{\mathcal{Y}}(0)\in\mathcal{Y}_{0}$ is a smooth point. Hence we complete the proof of the former claim for general $n$. The claim for $n=2$ follows from the same arguments, once we combine with the simpler proof of Theorem 3.6 of [Od24a] using the comparison of the orders of the local fundamental groups. ∎
Further, note that the above analogue to resolution of singularities $\tilde{X}$ is often taken canonically if $\mathcal{X}_{0}$ has quasi-regular minimizer of $\widehat{\rm vol}(-)$.
In our companion paper [Od24a], we restrict our focus to cones where the similar method applies to prove the “proper”-ness part of the following theorem.
Theorem 2.14 (K-moduli of Fano cone [Od24a]).
Consider the set (resp., all $\mathbb{Q}$-Gorenstein families) of K-polystable (resp., K-semistable) $\mathbb{Q}$-Fano cone of $n$-dimension, with the fixed normalized volume $V$ of the vertices. Then, it forms a good proper (and separated) moduli algebraic space and its stacky enhancement, generalizing the K-moduli of $\mathbb{Q}$-Fano varieties.
2.2.3. Negative deformation
We discuss the notion of negative weight deformation in the sense of [CH22] and equivalently rephrase more algebraically in terms of the generalized test configurations of [Od24b]. This finishes the proof of the nontrivial direction of [Od24b], and answers the Open Question 1.4 (2) of [CH22]. Note that, in the context of the previous section, the degenerations $\mathcal{X}^{\prime}_{0}\rightsquigarrow W$ and $\mathcal{X}^{\prime\prime}_{0}\rightsquigarrow C$ are both negative weight deformations by the construction.
Theorem 2.15 (Negative weight deformation).
  1. ($\mathrm{i}$)
    ([Od24b, Lemma 2.19]) Take an arbitrary algebraic negative weight deformation (cf., Definition 2.6 of [Od24b]) of a Fano cone $T\curvearrowright(Y,\xi)$ as $p\colon\mathcal{Y}\twoheadrightarrow U_{\tau}$, with $\xi\in-\tau$. Then, the general fiber of $p$ is a negative $\xi$-weight deformation in the original sense of Definition 1.7 of [CH22].
  2. ($\mathrm{ii}$)
    Suppose that for a Fano cone $(T\curvearrowright Y,\xi)$, $Y$ has only an isolated singularity. Then, conversely to the above ($\mathrm{i}$), any negative $\xi$-weight deformation $X$ of normal affine variety of $Y$ with $\xi$ in the sense of [CH22] arises in this manner.
  3. ($\mathrm{iii}$)
[…]
In general, the direction of ($\mathrm{ii}$) and ($\mathrm{iii}$) has a difficulty to treat a priori infinite dimensional deformation space (cf., [SRZ23, §6]). The latter ($\mathrm{iii}$) is modelled on and include the examples of [Li19, CR17, HN, Sze19, Sze20, Chi22].
Proof.
The above item ($\mathrm{i}$) is proved in [Od24b], so it remains to prove ($\mathrm{ii}$) and ($\mathrm{iii}$).
We start with the case ($\mathrm{ii}$). We first confirm the existence of $T$-equivariant semi-universal deformation of $T\curvearrowright Y$ after Pinkham. The tangent space $\mathbb{T}^{1}$ of the deformation of $Y$ naturally admits $T$-action by [Pin74, Proposition 2.2] and the arguments (2.5) (2.6) of loc.cit. The existence of the equivariant family follows from the arguments in the proof of Schlessinger [Sch68, Theorem 2.11], as in [Pin74, (2.9), (2.10)]. Indeed, the inductive construction of $J_{q}$ in (2.9) of loc.cit (or [Sch68, 2.11]) extends without any change. We similarly take $F^{0}$ as a column vector of $T$-homogeneous equations $f_{i}(i=1,\cdots,N)$ of some $T$-equivariant embedding $Y\hookrightarrow\mathbb{A}^{N}$, and construct the lift $F^{q}$ whose entries are all $T$-equivariant by projection to $T$-eigenspace at the inductive step for each $q=0,1,2,\cdots$. Thus, the upshot of the necessary modification is simply to replace the degree (with respect to the $\mathbb{G}_{m}$-action) to the weight vector with respect to the $T$-action. Summarizing up, we obtained a $T$-equivariant semiuniversal deformation $\mathcal{U}\to{\rm Def}(Y)$ as a formal family. It can be retaken as an algebraic family by Artin’s theorem [Art69], which we denote by the same symbol.
From the construction, note that ${\rm Def}(Y)$ comes with an $T$-equivariant étale morphism $u\colon{\rm Def}(Y)\to\mathbb{T}^{1}$, where the latter means the affine space scheme for $\mathbb{T}^{1}$.
We decompose $\mathbb{T}^{1}$ to $T$-eigensubspaces as $\mathbb{T}^{1}_{\alpha}$ for each character $\alpha\in M:={\rm Hom}(T,\mathbb{G}_{m})$. Then we extract $\xi$-negative part as
$\mathbb{T}^{1}_{\xi<0}$ $:=\oplus_{\alpha\text{ s.t. }\langle\xi,\alpha\rangle<0}\mathbb{T}^{1}_{\alpha},$
${\rm Def}(Y)_{\xi<0}$ $:=\mathbb{T}^{1}_{\xi<0}\cap{\rm Def}(Y),$
$\mathcal{U}_{\xi_{<}0}$ $:=u^{*}({\rm Def}(Y)_{\xi<0}),$
and consider the flat affine deformation $u|_{\xi<0}\colon\mathcal{U}_{\xi_{<}0}\to{\rm Def}(Y)_{\xi<0}$ as the restriction of $u$. Because of the $T$-equivariance of $u$ and the conditions $\langle\xi,\alpha\rangle<0$, ${\rm Def}(Y)_{\xi<0}\to\mathbb{T}^{1}$ is injective, hence an immersion.
Now, if we consider the negative $\xi$-deformation of $Y$ in the sense of [CH22, Definition 1.7] and take the realizing sequence as $(X_{i},\xi_{i})_{i=1,2,\cdots}$. From the conditions (1) and (2) of [CH22, Definition 1.7] (also cf., its (4)), it follows that $[X_{i}]$ are in the same $T(\mathbbm{k})$-orbit of ${\rm Def}_{\xi<0}(\mathbbm{k})$. Hence if we take its normalization of the closure of the $T(\mathbbm{k})$-orbit, we obtain an affine toric variety which we denote as $U_{\tau}$ with $\tau\subset N_{\mathbb{R}}$. Thus we obtain the desired assertion of ($\mathrm{ii}$).
The proof of ($\mathrm{iii}$) is similar, after proving the finite dimensionality of subspace of versal deformation space in concern. Suppose that $X$ is a $T$-equivariant weighted hypersurface cone i.e., there is a polynomial $f(z_{0},\cdots,z_{n})\in\mathbbm{k}[z_{0},\cdots,z_{n}]$ which is in particular $\xi=(w_{0},\cdots,w_{n})$homogeneous i.e., every monomial terms of $f$ has the same degree $d$ with respect to $\xi$.
If $y$ is an isolated singularity, as in ($\mathrm{ii}$), there are only finitely many monomials $g_{1},\cdots,g_{s}$ of $z_{i}$s whose degree with respect to $\xi$ is less than $d$. In this case, the semi-universal negative $\xi$-deformation is written as
$({\rm Def}_{x}(Y))_{\xi<0}$ $=\mathbb{A}^{s}_{t_{1},\cdots,t_{s}},$
$\mathcal{U}_{\xi<0}$ $=V(f+\sum_{i=1}^{s}t_{i}g_{i}).$
If $Y$ have a $1$-dimensional singularity, the singular locus is a line i.e., $1$-dimensional linear subspace of $\mathbb{A}^{n+1}_{z_{0},z_{1},\cdots,z_{n}}$ which we can assume to be of the form $(z_{1}=\cdots=z_{n}=0)$ after a $T$-equivariant linear transformation of the coordinates. Then, we can assume $f$ is a polynomial of $z_{1},\cdots,z_{n}$. Similarly as above, we take a homogeneous basis of the Jacobian ring ${\rm Jac}(f)=\mathbbm{k}[z_{1},\cdots,z_{n}]/\langle\{\frac{\partial f}{\partial z_{i}}\}\rangle$ and denote the lift as $g_{i}$s. Similarly as above, in this case, the semi-universal negative $\xi$-deformation achieved on $\mathbb{A}^{n}$ is written as
$({\rm Def}_{x}(Y))_{\xi<0,\mathbb{A}^{n}}$ $=\mathbb{A}^{s}_{t_{1},\cdots,t_{s}},$
$\mathcal{U}_{\xi<0,\mathbb{A}^{n}}$ $=V(z_{0}+f+\sum_{i=1}^{s}t_{i}g_{i})$
$\simeq({\rm Def}_{x}(Y))_{\xi<0,\mathbb{A}^{n}}\times\mathbb{A}^{n}.$
This is the semi-universal object of those negative deformations which are affine spaces $\mathbb{A}^{n}$ as varieties. The rest of the proof is the same as ($\mathrm{ii}$). ∎
Note that in the proof of above Theorem 2.15 ($\mathrm{ii}$) and ($\mathrm{iii}$), the following is proved, which we extract for later convenience.
Corollary 2.16 (Semi-universal negative weight deformation).
  1. (a)
    For each Fano cone $(T\curvearrowright Y,\xi\in N_{\mathbb{R}})$ with the only isolated singularity $y$, there is a semi-universal negative weight deformation $u_{\xi<0}\colon\mathcal{U}_{\xi<0}\twoheadrightarrow({\rm Def}_{x}(Y))_{\xi<0}\subset{\rm Def}_{x}(Y)$ over a finite type base with a good $T$-action $({\rm Def}_{x}(Y))_{\xi<0}$ (in the sense of [LS13]) such that any (algebraic) negative $\xi$-weight deformation can be obtained by pullbacks of $\pi_{\xi<0}$.
  2. (b)
    For each Fano cone $(T\curvearrowright Y,\xi\in N_{\mathbb{R}})$ with the vertex $y$, we suppose $Y$ is an affine (algebraic) hypersurface with $1$-dimensional singular locus, which is smooth (line).
    Then if we consider all negative deformations $X$ of $(Y,\xi)$ which are isomorphic to $\mathbb{A}^{n}$, there is a semi-universal deformation of them $\mathcal{U}_{\xi<0,\mathbb{A}^{n}}\to({\rm Def}_{x}(Y))_{\xi<0,\mathbb{A}^{n}}$.
After the construction of the semi-universal negative deformation as above, it is natural to consider the corresponding moduli theory (cf., also [SRZ23, §6]). Here are some questions, which we hope to explore more in near future.
Definition 2.17.
For the above Kuranishi space ${\rm Def}_{\xi<0}(Y)$ of Corollary 2.16 (a) (resp., ${\rm Def}_{\xi<0,\mathbb{A}^{n}}(Y)$ of Corollary 2.16 (b)) for the negative $\xi$-weight deformation in Corollary 2.16, note that there is a canonical $G:={\rm Aut}_{T}(Y)$-action (automorphism group of $Y$ as Fano cone) with canonical linearization which is trivial on the origin $0=[Y]$. If $({\rm Def}_{\xi<0}(Y)\setminus 0)\curvearrowleft G$ is prestable or semistable in the sense of [Mum65, Chapter I, §4], we define the corresponding GIT quotient $M^{\xi<0}_{Y}:=({\rm Def}_{\xi<0}(Y)\setminus 0)//G$ (resp., $({\rm Def}_{\xi<0,\mathbb{A}^{n}}(Y)\setminus 0)//G$).
Recall $G$ is reductive as proved in [DS17, Appendix] for $\mathbbm{k}=\mathbb{C}$, and for general $\mathbbm{k}$, it is algebraically proved [Od24a, Corollary 3.13].
Question 2.18.
In the above setup of Definition 2.17, does the corresponding GIT quotients
$M^{\xi<0}_{Y}:=({\rm Def}_{\xi<0}(Y)\setminus 0)//G$
admit moduli interpretation of affine $\partial\bar{\partial}$-exact complete Calabi-Yau metrics with Euclidean volume growth? Does the prestability or semistability condition always hold?
Remark 2.19.
For instance, if $G=T=\mathbb{G}_{m}$, the latter is true and $M_{Y}^{\xi<0}$ always comes with a canonical realization as a closed (projective) subscheme inside a weighted projective space. If $Y$ is toric with only isolated singularity, then it is also true by [Alt94, Alt97] (indeed, for $n\geq 4$, the tangent space is trivial). Note that the recent interesting non-existence result of asymptotically conical Kähler metric on $G_{2}$ type complex symmetric spaces $X$ of rank $2$ with a specified asymptotic (horospherical) cone $(Y,\xi)$ [Ngh24, Corollary D] is due to the discrepancy of the (metric) Reeb vector field and the negative valuation i.e., $X\notin{\rm Def}_{\xi}^{-}(Y)$, hence is of different nature from what is discussed here. We thank T.T.Nghiem for the clarification on this point.
2.3. Analogue for Kähler-Ricci solitons - Properness of moduli and bubbling solitons
This subsection takes a brief and temporary detour from the main discussion, for a few pages, to consider another application of our methods to more global problem. That is, we apply to the (expected) properness problem of moduli of compact shrinking Kähler-Ricci soliton (cf., [PSS23, GPSS18, CSW18, Ino19a, HL20, BLXZ23]), and provides a global analogue of bubbling in that context.
In particular, we prove the expected properness modulo some natural but technical input from birational geometry of K-stability. The reason this material is put here is that technically its proof is very similar to the above proofs of Theorem 2.4, 2.6 (and [Od24a]) and is even somewhat easier, since the objects are proper rather than affine. In this subsection, we assume the base field $\mathbbm{k}$ is algebraically closed field of characteristic 00.
Recall that the work of Chen-Sun-Wang on Fano manifolds [CSW18] provides another $2$-step degeneration
$X\rightsquigarrow(T\curvearrowright X^{\prime},\xi^{\prime})\rightsquigarrow(T\curvearrowright X^{\prime\prime},\xi^{\prime\prime}),$
where $(T\curvearrowright X^{\prime},\xi^{\prime})$ and $(T\curvearrowright X^{\prime\prime},\xi^{\prime\prime})$ are (modified) K-semistable and K-polystable in the sense of [BN14] respectively, hence the latter admitting (weak) Kähler-Ricci solitons as the polarized limit space (Gromov-Hausdorff limit) of the normalized Kähler-Ricci flow. Later it is more algebraically reformulated by the work of [DeSz20, Ino19a, HL20, BLXZ23]. In particular, Dervan-Szekelyhidi [DeSz20] introduces  [^13] Here, we use the opposite sign convention of [DeSz20] following [HL20, BLXZ23]. the $h$-invariant $h(X)\in\mathbb{R}_{\leq 0}$ for any $\mathbb{Q}$-Fano variety $X$ ([HL20] for the kawamata-log-terminal singular pairs case).
So, given our construction [Od24a] of the moduli of possibly irregular Calabi-Yau cones, especially the proof of the properness by using the higher $\Theta$-stable reduction [Od24b], it now becomes natural to expect that an analogous method may work for the moduli of Fano varieties with (weak) shrinking Kähler-Ricci solitons.
As the setup, let $N$ be a free abelian group of finite rank and we consider an algebraic $\mathbbm{k}$-torus $T=N\otimes\mathbb{G}_{m}$ which acts on log terminal $\mathbb{Q}$-Fano varieties $X$ of dimension $n$, together with $\xi\in N\otimes\mathbb{R}$ so that the triple $(T\curvearrowright X,\xi)$ is K-semistable and its anticanonical volume $(-K_{X})^{n}$ is $V$ (or consider its natural logarithmic generalization, for fixed coefficients set). We consider their moduli Artin stack $\mathcal{M}^{s}_{n,V,T,\xi}$, as an easy proof of its existence is included in our Theorem 2.21 for readers convenience, following e.g., [PSS23, Ino19a, BLXZ23]. We also fix $n,V,h$ and consider the finite type moduli Artin stack $\mathcal{M}_{n,V,h}$ (resp., $\mathcal{M}_{n,V,h^{+}}$), after [BLXZ23, 6.1, 6.4], of $n$-dimensional $\mathbb{Q}$-Fano varieties $X$ with anticanonical volume $(-K_{X})^{n}=V$ and $h(X)=h$ (resp., $h(X)\geq h$).
The latter is (essentially) constructed in [BLXZ23, §6, Theorem 6.1 (and 6.4)], as loc.cit Theorem 6.1 proves the existence of $\mathcal{M}^{\rm Fano}_{n,V,(1,)h^{+}}$ which is our $\mathcal{M}_{n,V,h^{+}}$, and it follows from their Theorem 6.4 that $\mathcal{M}_{n,V,h}$ is a locally closed substack of $\mathcal{M}_{n,V,h^{+}}$. Indeed, by the same theorem, it even follows that $\mathcal{M}_{n,V,h^{+}}$ has a finite stratification by of $\mathcal{M}_{n,V,h^{\prime}}$ for finite possibilities of $h^{\prime}(\geq h)$. Now we hope that the following analogue of [Che24, 1.1], [Od24a, Theorem 3.5] hold.
Conjecture 2.20 (Higher $\Theta$-stratification by $h$-invariant of Fano varieties).
The $h$-invariants of $\mathbb{Q}$-Fano families and the corresponding first step degeneration of [CSW18, BLXZ23] (as reformulated in [HL20, §2], [Od24b, Example 2.19, Corollary 2.17]) satisfies the following general properties.
  1. ($\mathrm{i}$)
    Consider any finite type $\mathbbm{k}$-scheme $S$ and a $\mathbb{Q}$-Gorenstein 141414locally stable in the sense of [Kol23] family $\pi\colon\mathcal{X}\to S$, whose $h$-invariants and volumes are both constant i.e., $h(X_{s})$, $(-K_{X_{s}})^{n}$ does not depend on closed points $s\in S$, where $X_{s}:=\pi^{-1}(s)$. Then, there is a finite rank free abelian group $N$ and a rational polyhedral cone $\tau\subset N\otimes\mathbb{R}$ together with $\xi\in S$ such that there is a $\mathbb{Q}$-Gorenstein family $\tilde{\pi}\colon\tilde{\mathcal{X}}\to U_{\tau}\times S$ which restricts to $\pi$ (resp., the corresponding first step degeneration of $X_{s}$ [CSW18, BLXZ23]) on $(1,\cdots,1)\times S$ (resp., $\Delta_{\xi}\times s\to U_{\tau}\times s$, on each $s$). Here, $U_{\tau}$ means the affine toric variety corresponding to $N_{\mathbb{R}}\supset\tau$ and $\Delta_{\xi}$ is the N-spectrum in [Od24b, §2] (cf., also [HL20, §2]).
  2. ($\mathrm{ii}$)
    For a positive integer $n$ (the dimension), real numbers $V$ (the anticanonical volume, hence positive) and $h$ (lower bound for the $h$-invariant), the finite type moduli Artin stack of $\mathbb{Q}$-Fano varieties $X$ with $h(X)\geq h$ $\mathcal{M}_{n,V,h^{+}}$ (cf., [BLXZ23, 6.1]) and its natural finite stratification $\{\mathcal{M}_{n,V,h^{\prime}}\}_{h^{\prime}(\geq h)}$ by the values of $h(X)$ ([BLXZ23, Theorem 6.4]) admits the structure of a higher $\Theta$-stratification in the sense of [Od24b, §3.3], by encoding the aforementioned first step degeneration of $\mathbb{Q}$-Fano varieties [CSW18, BLXZ23] (reformulated in [Od24b, Ex 2.19]).
Note that (the special case when $S$ is a smooth curve of) the above ($\mathrm{i}$) easily implies ($\mathrm{ii}$). Indeed, because the $H^{\rm NA}$-invariant for a special $\mathbb{R}$-test configuration only depends on the central fiber with the vector field (cf., [HL20, Example 2.32 (87), §2.5 (73)]), together with the fact that the Duistermaat-Heckman measure of $\mathbb{R}$-filtration is determined by the dimensions of the each filters (cf., [BC11], [Od24b, §2.6]), such $\tilde{\pi}$ can exists only if $h(X_{s})$ is constant. The rest follows from the similar arguments as [Od24a, 3.13 (§3.3, §3.4)] after [HL14, AHR20]. See also a related criterion of the $\Theta$-stratification in [HL14], given by two conditions (S) and (B) therein.
Then if the above ($\mathrm{ii}$) holds, now that [Od24b] prepared, the higher $\Theta$-stable reduction theorem §3 loc.cit applies and we readily obtain its application as below. This is similar to [Od24a, §3] for the Calabi-Yau cone case and our Theorems 2.4, 2.6.
Theorem 2.21 (Compactness).
Under the above setup, we assume the conjecture 2.20 holds. Then, both the moduli algebraic $\mathbbm{k}$-stacks $\mathcal{M}_{n,V,h}$ and $\mathcal{M}^{s}_{n,V,T,\xi}$ are universally closed. 151515i.e., satisties the existence part of the valuative criterion of properness of an algebraic stack. The actual properness requires separatedness in some of the literature.
Also, the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ of (modified) K-semistable pair $(X\curvearrowleft T,\xi)$ exists. If $\mathbbm{k}=\mathbb{C}$ and the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ admits a coarse good moduli space $M^{s}_{n,V,T,\xi}$, its complex analytification $M^{s}_{n,V,T,\xi}(\mathbb{C})$ is compact.
Remark 2.22.
In particular, by setting $h:=0$, we give a yet another proof of the properness of K-moduli stack of $\mathbb{Q}$-Fano varieties (modulo Conjecture 2.20), different from [BHLLX21, LXZ22] which uses the $\delta$-invariant [FO18, BJ18].
Proof.
Consider any stalk of the structure sheaf on a complete $\mathbbm{k}$-curve at a closed point and denote the DVR as $R$, its fraction field (the function field) as $K$. Take an arbitrary $\mathbbm{k}$-morphism $\varphi^{o}\colon{\rm Spec}K\to\mathcal{M}_{n,V,h_{1}=h}$, corresponding to a $\mathbb{Q}$-Fano family $\pi^{o}\colon\mathcal{X}^{o}\to{\rm Spec}(K)$. By [LX18], after a finite base change (we keep the notation of $R,K$ though), $\pi^{o}$ extends to a $\mathbb{Q}$-Fano family $\pi\colon\mathcal{X}\to{\rm Spec}(R)$. We suppose the $h$-invariant of the central fiber (resp., generic fiber) of $\pi$ is $h_{0}$ (resp., $h_{1}$). By [BLXZ23], we have $h_{0}\leq h_{1}$. Correspondingly, so that $\varphi^{o}$ extends to a $\mathbbm{k}$-morphism from ${\rm Spec}(R)$ to $\mathcal{M}_{n,V,h_{0}^{+}}$ for some $h_{0}\leq h_{1}$.
What we need to prove is that, if $h_{0}<h_{1}$, we have another extension of $\varphi^{o}$ to $\mathcal{M}_{n,V,h_{1}=h}$ as $\varphi\colon{\rm Spec}(R)\to\mathcal{M}_{n,V,h_{1}=h}$ possibly after finite extension of $R$. If conjecture (2.20) ($\mathrm{ii}$) is true, then the assertion is a direct corollary to the the higher $\Theta$-stable reduction along the higher $\Theta$-stratification [Od24b, §3.3, Theorem 3.18]. Note that by boundedness results obtained in [BLXZ23, Jia20], possible values of $h(X)$ for $X\in\mathcal{M}_{n,V,h^{+}}$ are finite.
The existence of $\mathcal{M}^{s}_{n,V,T,\xi}$ as an algebraic $\mathbbm{k}$-stack follows similarly as [Od24a]: it is equivalent to the stack of $T$-equivariant $\mathbb{Q}$-Gorenstein families of $n$-dimensional $\mathbb{Q}$-Fano varieties of volume $V$, whose $H^{\rm NA}$-minimizer is $\xi$, by its uniqueness ([HL20, 1.2], [BLXZ23, 1.2]), and the constructible lower semicontinuity of the $h$-invariant ([BLXZ23, 6.4]). Indeed, the whole stack $\mathcal{M}^{T}$ of $T$-equivariant $\mathbb{Q}$-Gorenstein families of $n$-dimensional $\mathbb{Q}$-Fano varieties $X$ of volume $V$ is an algebraic $\mathbbm{k}$-stack. This comes from the fact that the classical affine cone $C(-NK_{X}):={\rm Spec}\oplus_{m\geq 0}H^{0}(-mNK_{X})$ for a uniform sufficiently divisible positive integer $N$, together with the action of $T\times\mathbb{G}_{m}$ on it, recovers $X$. Indeed, the data $C(-NK_{X})\curvearrowleft(T\times\mathbb{G}_{m})$ forms an algebraic stack due to the existence of the (quasi-projective) multi-Hilbert scheme [HM04] and the hulls and husks construction by Kollár [Kol08]. Then our $\mathcal{M}^{s}_{n,V,T,\xi}$ is straightforwardly seen to be an open substack of $\mathcal{M}^{T}$ by the aforementioned results in [HL20, BLXZ23].
Take a connected component $\mathcal{M}$ of $\mathcal{M}^{s}_{n,V,T,\xi}$, put the corresponding $h$-invariant as $h$ and consider the connected component of $\mathcal{M}_{n,V,h}$ which includes the natural image of $\mathcal{M}$, which we denote as $\mathcal{M}_{h}$. Therefore, the forgetful morphism $\mathcal{M}\to\mathcal{M}_{h}$ between the two algebraic $\mathbbm{k}$-stacks is radiciel and finite. Since the latter is universally closed, so is the former. We complete the proof. ∎
In conclusion, we can expect very similar methods of constructing the coarse (good) moduli space $\mathcal{M}^{s}_{n,V,T,\xi}$ of (weak) Kähler-Ricci solitons, to the construction of the coarse good moduli space of the moduli stack of the Calabi-Yau cones in [Od24a]. Also note that the above algebraic compactness result is compatible with the precompactness type result by [PSS23, GPSS18].
2.4. Re-construction of minimal bubble via coordinates
Now, finishing the digression in the previous subsection, we come back to the original context of affine klt varieties and the local bubbling problem. From what follows, for simplicity, we work over $\mathbbm{k}=\mathbb{C}$ under the Setups 1 (a) $\land$ 2 $\land$ 3. Further we assume that $\mathcal{X}\to\Delta$ is a restriction of algebraic morphism to a unit disk $\Delta$. As it is often obvious, many of the following discussions also work verbatim for algebraic families over other base fields $\mathbbm{k}$ (as in Setup 1 (b)).
In this subsection, we discuss a priori different (but conjecturally equivalent) and more practical reconstruction of the minimum bubble, in terms of coordinates and defining equations. For that, we work again under the setup 1, 2, and further suppose that $X$ and $W$ is obtained inside the affine space $\mathbb{C}^{l}$ and $v_{X}$ is obtained by weights $(w_{1},\cdots,w_{l})\in\mathbb{R}_{>0}^{l}$ corresponding to the coordinates $z_{i}$ of $\mathbb{C}^{l}$. For certain sufficiently divisible $N$ and $w_{i}$s, this is always achievable ([DS17, §3.3]). We would like to put the weight on $t$, the base coordinate, as $w_{0}\in\mathbb{R}_{>0}$ and $w:=(w_{0},\cdots,w_{l})$. In this subsection, we further assume $\mathcal{X}$ is the restriction of affine morphism over $\mathbb{A}_{t}^{1}$ for technical simplicity.
We now temporarily fix the extended embedding of $\mathcal{X}$ into $\Delta\times\mathbb{C}^{l}$, and consider $\mathcal{W}:=\mathcal{W}(w_{0};w_{1},\cdots,w_{l}):=\mathop{\mathrm{Spec}}\nolimits(gr_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x}))$ as a family over $\mathbb{A}^{1}_{t}(\mathbb{C})=\mathbb{C}$ where $t$ is the coordinate of $\Delta$, as in [dBS23]. Here, $gr_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ is the natural graded ring
$\oplus_{m\geq 0}\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{C}^{l},0},w(f)\geq m\}/\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{C}^{l},0},w(f)>m\},$
where for $f=\sum a_{i_{0},i_{1},\cdots,i_{N}}t^{i_{0}}\prod_{1\leq j\leq N}z_{j}^{i_{j}}$, the weight $w(f)$ of $f$ means $\min\{\sum_{j}w_{j}i_{j}\mid a_{i_{0},i_{1},\cdots,i_{N}}\neq 0\}$ and $\overline{f}$ means the restriction.
Proposition 2.23.
In the above context,
  1. ($\mathrm{i}$)
    there is a critical value $w_{\rm cri}\in\mathbb{R}_{>0}$, such that
    • $w_{0}>w_{\rm cri}$, then $\mathcal{W}$ is canonically isomorphic to $\mathbb{A}^{1}_{t}\times W$ over $\mathbb{A}^{1}_{t}$.
    • $w_{0}=w_{\rm cri}$, then $\mathcal{W}$ is a normal affine faithfully flat and $\mathbb{G}_{m}$-equivariant family with the central fiber $W$, where $\mathbb{G}_{m}$ acts on the base $\mathbb{A}_{t}^{1}$ with the weight $w_{0}$.
  2. ($\mathrm{ii}$)
    In particular, for $w_{0}\geq w_{\rm cri}$, the weights $w_{0},w_{1},\cdots,w_{l}$ gives a unique valuation on $\mathcal{O}_{\mathcal{X},x}$.
This proposition is inspired by [Sun23, dBS23] and the discussions with the authors.
Proof.
As we assumed in the beginning of this subsection, $\mathcal{X}\subset\Delta\times\mathbb{C}^{l}$ is cut out by some defining polynomials $F_{1},\cdots,F_{m}$ of $z_{1},\cdots,z_{N},t$. We (again) take enough $F_{i}$s to form a universal Gröbner basis of the generating ideal ([BCRV22, §5.1]). As $x=0\in\mathcal{X}$, these do not contain constant terms. Then $\mathcal{W}\subset\mathbb{A}_{t}^{1}\times\mathbb{C}^{l}$ is cut out by the initial terms ${\rm in}_{w_{0},w_{1},\cdots,w_{l}}(F_{i})$ of $F_{i}$ with respect to the weights $w_{0},w_{1},\cdots,w_{l}$. Thus in particular, it admits the natural $\mathbb{C}^{*}$-action which equivariantly projects down to $\mathbb{A}_{t}^{1}$ with the weight $w_{0}$.
We set $f_{i}:=F_{i}|_{t=0}$. Since $F_{i}=f_{i}+tg_{i}$ for some polynomials $g_{i}$s, for large enough $w_{0}$, there is a critical value $w_{\rm cri}$ such that for the initial term of $F_{i}$ is $f_{i}$ for any $i$ if and only if $w_{0}>w_{\rm cri}$. From the construction, if $w_{0}=w_{\rm cri}$, ${\rm in}_{w_{0},w_{1},\cdots,w_{l}}(F_{i})|_{t=0}=f_{i}$ for each $i$ so that the central fiber of $\mathcal{W}$ is still $W$. For general weights $w_{0},\cdots,w_{l}$, $gr_{w_{0},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ may not be flat over $\mathbb{A}^{1}$, but in our case, since the central fiber is normal, $n$-dimensional, and $\mathcal{W}$ is $\mathbb{G}_{m}$-acted, $\mathcal{W}(w_{\rm cri},w_{1},\cdots,w_{l})$ is normal and faithfully flat over $\mathbb{A}^{1}$. This completes the proof of the first claim ($\mathrm{i}$) of the proposition. In particular, $gr_{w_{\rm cri},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ is integral so that $w_{\rm cri},w_{1},\cdots,w_{l}$ naturally gives a valuation on $\mathcal{O}_{\mathcal{X},x}$ (cf., [Tei14, proof of Proposition 3.3. a) and comments]). This completes the proof of the item ($\mathrm{ii}$). ∎
Conjecture 2.24 (cf., [Sun23, dBS23]).
For a good choice of $N$, embedding of $\mathcal{X}$ to $\Delta\times\mathbb{C}^{l}$ and $w_{1},\cdots,w_{l}$, $\mathcal{W}(w_{\rm cri};w_{1},\cdots,w_{l})\to\mathbb{A}^{1}$ is a test configuration $\mathcal{T}$, which degenerates the central fiber $\mathcal{X}^{\prime}_{\rm min,0}$ of $\mathcal{X}$ in Theorem 2.4 isotrivially to $W$ (compare the proof of Theorem 2.4).
Note that [dBS23] implicitly confirmed the above conjecture in classical examples in low dimension which inspire above (we also did some similar calculations to loc.cit, confirming the above in higher dimensional examples e.g. generic hypersurfaces). Indeed, to obtain the minimum bubble in this general, we need a good choice of extension of the embedding $X$ into $\mathbb{C}^{l}$ to $\mathcal{X}$ to $\Delta\times\mathbb{C}^{l}$ as the construction of [Sun23, Lemma 4.6] and the following bad example suggests.
Example 2.25.
Consider the case $N=3$, when $\mathcal{X}$ is defined by $((z_{1}+tz_{3})^{2}+(z_{2}+tz_{3})^{2}=z_{3}^{m}+t^{h})$ in $\Delta\times\mathbb{C}^{l}$ for $h\gg m\geq 2$. Then if we still consider the coordinates $z_{i}$s and $t$,
  • $\mathcal{W}(w_{0};w_{1},\cdots,w_{l})\to\mathbb{A}^{1}$ for $w_{0}>w_{\rm cri}$ is $(z_{1}^{2}+z_{2}^{2}=z_{3}^{m})\simeq W\times\mathbb{A}^{1}$,
  • $\mathcal{W}(w_{\rm cri};w_{1},\cdots,w_{l})\to\mathbb{A}^{1}$ is $((z_{1}+tz_{3})^{2}+(z_{2}+tz_{3})^{2}=z_{3}^{m})\simeq W\times\mathbb{A}^{1}$.
2.5. Re-construction of minimal bubble via valuation
Taking Theorem 2.4 ($\mathrm{x}$) and Proposition 2.23 ($\mathrm{ii}$) into account, for instance, it is reasonable to pass our context to general valuations and discuss characterization of the critical valuations in the sense of Theorem 2.4 ($\mathrm{x}$). Our discussions in this subsection are totally coordinates-free.
We set the space of real valuations $v$ (resp., $v_{X}$) of $\mathcal{O}_{X,x}$ (resp., $\mathcal{O}_{X,x}$) whose center $c_{\mathcal{X}}(v)$ (resp., $c_{X}(v)$) is $x\in\mathcal{X}$ (resp., $x\in X$) as $\mathop{\mathrm{Val}}\nolimits(x\in\mathcal{X})$ (resp., $\mathop{\mathrm{Val}}\nolimits(x\in X)$). Their subsets which consist of quasi-monomial valuations will be denoted by $\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in\mathcal{X})$ (resp., $\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in X)$).
For each $v\in\mathop{\mathrm{Val}}\nolimits(x\in\mathcal{X})$ (resp., $v_{X}\in\mathop{\mathrm{Val}}\nolimits(x\in X)$), we have valuation ideals
$\mathcal{J}_{m}(v)$ :=$:=\{F\in\mathcal{O}_{\mathcal{X},x}\mid v(F)\geq m\}$
[…]
as a sequence of coherent ideals. For later convenience, we also set $\mathcal{J}_{>m}(v):=\{f\in\mathcal{O}_{\mathcal{X},x}\mid v(f)>m\}$. The following may evoke the discussion in [Li19, §6] after [MSY08]. S.Boucksom pointed out it is somewhat analogous to the Gauss extension in the isotrivial case (cf., e.g., [BJ18, §1]) in spirit.
Proposition 2.26.
  1. ($\mathrm{i}$)
    For each $v_{X}\in\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in X)$, $a\in\mathbb{R}_{>0}$, with $a\ll 1$, say $a\in(0,a_{0})$ for some $a_{0}\in\mathbb{R}_{>0}$, we have a unique valuation $v\in\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in\mathcal{X})$ with $\mathcal{J}_{m}(v)|_{X}=J_{m}(v_{X})$ and $v(t)=\frac{1}{a}$.
  2. ($\mathrm{ii}$)
    Once one fixes a holomorphic section $\sigma\colon\Delta\to\mathcal{X}$, and a SNC model of $X$ (resp., [BFJ08, BFJ16]) which realizes $v$ as a quasi-monomial valuation, then in a longer interval (than ($\mathrm{i}$)) of $a$ as $a\in(0,a_{1})$ with $a_{1}>a_{0}$, $v_{a}$ of ($\mathrm{i}$) is still canonically determined. Henceforth, we sometimes denote this $a_{1}$ as $a_{1}(\varphi)$.
Proof.
There is a log resolution $\varphi\colon\tilde{X}\to X$ where the center of $v_{X}$ is a strata of simple normal crossing divisor $S=\cap_{i\in I}E_{i}$. We write $E_{i}=(z_{i}=0)$ with $z_{i}\in\mathcal{O}_{\tilde{X},\eta_{S}}$ for each $i\in I$, where $\eta_{S}$ means the generic point of $S$. By replacing $\varphi$ by a further blow up if necessary, we can and do assume $\varphi$ extends to a log resolution $\tilde{\varphi}\colon\tilde{\mathcal{X}}\to\mathcal{X}$. Then from the quasi-monomiality assumption of $v_{X}$, there is a set of real positive numbers $c_{i}$s that the $v_{X}$ is determined by $E_{i}$s with the condition $v_{X}(z_{i})=c_{i}$. For convenience, we also take other coordinates $z_{i}(i\notin I)$ which vanishes at the center of $v_{X}$ in $S\subset\tilde{X}$ to form an analytic local coordinate system $\{z_{i}\}$. Take an extension $z_{i}$ to $\mathcal{O}_{\tilde{\mathcal{X}},\eta_{S}}$ as $\tilde{z_{i}}$ and consider the quasi-monomial valuation $v_{a}$ determined by $\tilde{\varphi}$, $c_{i}$s and $v_{a}(t)=\frac{1}{a}$. Hence the existence claim is proven.
Next we consider the uniqueness and canonicity of $v_{a}$. For a different choice of $\tilde{z_{i}}$ as $\tilde{z_{i}}^{\prime}$, there is a holomorphic function $f_{i}$ of $z_{j}$s such that
(8) $\tilde{z_{i}}^{\prime}=z_{i}+tf_{i}(z_{1},\cdots,z_{n}).$
If $v_{a}(t)=\frac{1}{a}>v_{a}(z_{i})$, then by (8), the independencity of $v_{a}$ with respect to the lifted coordinates $\tilde{z_{i}}$ easily follows. This confirms ($\mathrm{i}$). On the other hand, if we fix a section $\sigma$, then we can impose the condition $\tilde{z_{i}}(\sigma(t))$ for each $i$. Under such constraints, $f_{i}$ of (8) does not have constant terms i.e., positive multiplicities with respect to $\tilde{z_{i}}$s. In that case, if $v_{a}(t)=\frac{1}{a}>\max\{c_{i}\}-\min\{c_{i}\}$, then the obtained quasi-monomial $v_{a}$ does not depend on the lift $\tilde{z_{i}}$s. Hence the proof of ($\mathrm{ii}$). ∎
Remark 2.27.
S.Boucksom kindly suggested to compare the construction with the composite (cf., e.g., [FK18, §6.4]) of the discrete valuation ${\rm ord}_{\mathcal{X}_{0}}$ on $\mathcal{O}_{\mathcal{X},x}$ and $v_{X}$ on $\mathcal{O}_{\mathcal{X}_{0},x}$. Although it gives a height $2$ (not real-valued) valuation, the author sees this as a kind of infinitesimal analogue of the above construction $\{v_{a}\}_{a}$, which may be its integral curve in some unknown sense.
Example 2.28.
For many ADE singularities with $n=2$, $a_{0}$ can be taken as $\infty$, while [dBS23, Example 7] shows that it is not necessarily the case for $n(>2)$-dimensional $A$-type singularities.
In this paper, we introduce yet another couple of “Yau-Tian-Donaldson type” conjectures for mild degenerations, after [Sun23, dBS23]. It hopefully connects the complex differential geometry of bubble (tree) to our algebro-geometric or non-archimedean framework.
Now, again take a section $\sigma\colon\Delta\to\mathcal{X}$ as Setup 2 and fix it. Then we have a metric analogue of (2.26):
Definition 2.29.
Under the assumption (3), for $a\in\mathbb{R}_{>0}$, we define $v^{\rm DG}_{a}\colon\mathcal{O}_{\mathcal{X},x}\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ as follows.  [^16] We do not expect this value would take $\infty$. For $G\in\mathcal{O}_{\mathcal{X},x}$, we set
$v^{\rm DG}_{a}(G):=\limsup_{t\to 0}\dfrac{\log\sup_{y\in B_{|t|^{a}}(\sigma(t))}(|G(y)|)}{-\log|t|}.$
Here, $B_{|t|^{a}}(\sigma(t))$ denotes the geodesic ball in $\pi^{-1}(t)=X_{t}$ centered at $\sigma(t)$ with radius $|t|^{a}$.
The following conjecture for the minimum bubble refines the inspiring predictions and discussions in [Sun23, dBS23].
Conjecture 2.30 (cf., [Sun23, 5.4], [dBS23, §4.1]).
Under the Setup 1 with a fixed closed base point $x\in X$, we now again take a quasi-monomial valuation $v_{X}$ of $\mathcal{O}_{X,x}$ which minimizes $\widehat{\operatorname{vol}}_{X}(-)$ ([DS17, §3.2], [Li19, Blu18]) and apply Proposition 2.26 and Definition 2.29 to consider $v_{a}$ and $v_{a}^{\rm DG}$. Then the following holds:
  1. ($\mathrm{i}$)
[…]
    • $a\in(0,a_{\rm min}]$, $A_{(\mathcal{X},X)}(v_{a})$ stays constant while
    • $A_{(\mathcal{X},X)}(v_{a})>A_{(\mathcal{X},X)}(v_{a_{\rm min}})$ for $a>a_{\rm min}$.
    Below, we set $v_{a_{\rm min}}=:v_{\rm min}$ as a notation.  [^17] Note that this is a valuation of the stalk of the total space, unlike $v_{X}$. If $r=1$, then this should be the minimal critical valuation (Theorem 2.4 ($\mathrm{x}$)).
  2. ($\mathrm{ii}$)
    After suitably multiplying $v_{X}$ by a positive constant, $(\sigma(t)\in X_{t},|t|^{-a_{\rm min}}g_{t})$ converges to a minimum bubble in the sense of [Sun23].
  3. ($\mathrm{iii}$)
[…]
  4. ($\mathrm{iv}$)
    If $v_{\rm min}$ is divisorial, the birational map $\mathcal{X}^{\prime}_{\rm min}\dashrightarrow\mathcal{X}$ of Theorem 2.4 corresponds to $v_{\rm min}$. Further, it is obtained as an open subset of a plt blow up ([Sho96, 3.1], [Pro00, 2.1]) of $\mathcal{X}$ (e.g., often they are open subsets of weighted blow ups of $\mathcal{X}$ as [dBS23, Remark 11]).
  5. ($\mathrm{v}$)
    If we define ${\rm gr}_{v_{\rm min}}(\mathcal{O}_{\mathcal{X},x}):=\oplus_{m\in{\rm Im}(v_{\rm min})}\mathcal{J}_{m}(v_{\rm min})/\mathcal{J}_{>m}(v_{\rm min})$, then it is a finitely generated $\mathbb{C}[t]$-algebra and $\mathcal{W}(v_{\rm min}):={\rm Spec}_{\mathbb{C}[t]}({\rm gr}_{v_{\rm min}}(\mathcal{O}_{\mathcal{X},x}))$ is a $\mathbb{C}^{*}$-equivariant flat family over $\mathbb{A}^{1}_{t}$ i.e., an affine test configuration. Further, for any $\mathbb{G}_{m}(\mathbb{C})$-invariant section $\sigma\colon\mathbb{A}^{1}_{t}\to\mathcal{W}(v)$, $\sigma(0)$ is always the same which we denote as $p(v)\in\mathcal{W}(v_{\rm min})|_{t=0}$.
  6. ($\mathrm{vi}$)
    There is an isotrivial degeneration181818with a different coordinate/direction from $t$, hence giving test configuration for each fiber of $\mathcal{W}(v_{\rm min})\to\mathbb{A}^{1}_{t}$ $\mathcal{W}(v_{\rm min})\rightsquigarrow\mathcal{B}(v_{\rm min})$ i.e., to be more precise, of $\mathcal{W}(v_{\rm min})\to\mathbb{A}^{1}_{t}$ to another test configuration $\mathcal{B}(v_{\rm min})\to\mathbb{A}^{1}_{t}$ (whose general fiber $\mathcal{B}(v_{\rm min})|_{t=c(\neq 0)}$ is the bubble $B(v)$ in Conjecture 3.1). The limit of $p(v_{\rm min})\in\mathcal{W}(v)|_{t=0}$ is denoted as $q_{\rm min}(v)\in\mathcal{B}(v)|_{t=0}$.
  7. ($\mathrm{vii}$)
    $p(v_{\rm min})\in\mathcal{W}(v_{\rm min})|_{0}$ (resp., $q(v_{\rm min})\in\mathcal{W}(v_{\rm min})|_{0}$) with the natural algebraic torus action, which is associated to the graded structure of ${\rm gr}_{v_{\rm min}}(\mathcal{O}_{\mathcal{X},x})$ and isotrivial degeneration structure of $\mathcal{W}$ and $\mathcal{B}$ give the $2$-step degeneration of [DS14, Li19, LLX20] to the metric tangent cone of $x\in X$.
2.6. A family version of local normalized volume
We analyze the following family version of the normalized local volumes (see the review of the original version in §2.1) motivated by a discussion with C.Spotti (cf., [dBS23, §4.1, esp. after Remark 10]) to whom the author is grateful. On the other hand, we do not mean this subsection to be the main focus of this paper; rather, it is a supplementary for our purpose and arose alongside Conjecture 2.30. ($\mathrm{i}$).
Definition 2.31.
In the above setup,
  1. ($\mathrm{i}$)
    for each $v\in\mathop{\mathrm{Val}}\nolimits(x\in\mathcal{X})$, we define
    $\operatorname{vol}_{\mathcal{X}}(v):=\limsup_{m\to\infty}\dfrac{{\rm length}(\mathcal{O}_{\mathcal{X},x}/\mathcal{J}_{m}(v)|_{X})}{m^{n}/n!},$
    after [ELS03].
  2. ($\mathrm{ii}$)
    For the same $v$, we also define a normalized version:
    $\widetilde{\operatorname{vol}}_{\mathcal{X}}(v):=(A_{(\mathcal{X},X)}(v))^{n}\cdot\operatorname{vol}_{\mathcal{X}}(v).$
  3. ($\mathrm{iii}$)
    Here is its sibling:
    $\widetilde{\operatorname{vol}}(\mathcal{J}):=(\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J}))^{n}\cdot\operatorname{mult}(\mathcal{J}|_{X}),$
    where $\mathcal{J}$ is an arbitrary coherent ideal of $\mathcal{O}_{\mathcal{X}}$ with cosupport $x\in\mathcal{X}$, and $\mathop{\mathrm{lct}}\nolimits$ stands for the log canonical threshold and $\operatorname{mult}$ stands for the multiplicity (here on the level of $X$).
  4. ($\mathrm{iv}$)
    More generally, if we have a graded sequence of coherent ideals $\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}\}_{m=1,2,\cdots}$ with cosupport $x\in\mathcal{X}$ i.e., satisfying $\mathcal{J}_{m_{1}}\mathcal{J}_{m_{2}}\subset\mathcal{J}_{m_{1}+m_{2}}$ for any $m_{1},m_{2}\in\mathbb{Z}_{>0}$ and $\mathcal{J}_{m}\supset\mathfrak{m}_{\mathcal{X},x}^{N_{m}}$ if $N_{m}\gg 0$ for each $m>0$, we similarly define
    $\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J}_{\bullet}):=(\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J}))^{n}\cdot\operatorname{mult}(\mathcal{J}_{\bullet}|_{X}).$
A subtle but crucial difference here with [Li19, Liu18] is that both terms of $A_{(\mathcal{X},X)}(v)$ and $\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J})$ in the above definition reflect the geometry of total space or the process of degeneration, not determined only by the central fiber $X\ni x$. We also note that for any real positive number $c>0$, $\widetilde{\operatorname{vol}}(cv)=\widetilde{\operatorname{vol}}(v)$ by the normalization effect.
Proposition 2.32.
In the above setup, we have
(9) $\inf_{c_{\mathcal{X}}(v)=x}\widetilde{\operatorname{vol}}_{\mathcal{X}}(v)$ $=\inf_{\mathcal{J}\subset\mathcal{O}_{\mathcal{X},x}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J})$
(10) $=\inf_{\begin{subarray}{c}\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}\}_{m},\\ \mathcal{J}_{m}\subset\mathcal{O}_{\mathcal{X},x}\end{subarray}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J}_{\bullet})$
(11) $\geq\widehat{\operatorname{vol}}(x\in X).$
We call the quantity (9)$=$(10) as $\widetilde{\operatorname{vol}}_{\mathcal{X}}(x)$. We conjecture (11) to be the equality, where $\widehat{\operatorname{vol}}$ is in the sense of [Li19].
Proof.
It naturally follows from similar arguments to [Liu18, §4.1] but we sketch the proof. The notion of valuation ideal gives a central tool for the proof.
First, for a given coherent ideal $\mathcal{J}$, from the definition of $\mathop{\mathrm{lct}}\nolimits$, there is a divisorial valuation $v_{0}$ with $\mathop{\mathrm{lct}}\nolimits_{(\mathcal{X},X)}(\mathcal{J})=\frac{A_{(\mathcal{X},X)}(v_{0})}{v_{0}(\mathcal{J})}$. For simplicity, we rescale the valuation to make $v_{0}(\mathcal{J})=1$. Then we have $\operatorname{mult}(\mathcal{J}|_{X})\geq\operatorname{vol}(\mathcal{J}_{v}|_{X})$ since $\mathcal{J}^{m}\subset\{f\in\mathcal{O}_{\mathcal{X},x}\mid v_{0}(f)\geq m\}$. Hence, we have $\inf_{c_{\mathcal{X}}(v)=x}\widetilde{\operatorname{vol}}_{\mathcal{X}}(v)\leq\inf_{\mathcal{J}\subset\mathcal{O}_{\mathcal{X},x}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J})$. For (10), from the arguments in [JM12] and [ELS03], we have $\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J}_{\bullet})=\lim_{m\to\infty}m\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J}_{m})$ so that
$\inf_{\mathcal{J}\subset\mathcal{O}_{\mathcal{X},x}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J})\leq\inf_{\begin{subarray}{c}\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}\}_{m},\\ \mathcal{J}_{m}\subset\mathcal{O}_{\mathcal{X},x}\end{subarray}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J}_{\bullet}).$
Further, for each $v$, the construction of the valuation ideals $\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}(v)\}_{m}$ easily implies
$\inf_{\begin{subarray}{c}\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}\}_{m},\\ \mathcal{J}_{m}\subset\mathcal{O}_{\mathcal{X},x}\end{subarray}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J}_{\bullet})\leq\inf_{c_{\mathcal{X}}(v)=x}\widetilde{\operatorname{vol}}_{\mathcal{X}}(v).$
Finally, note
$\mathop{\mathrm{lct}}\nolimits(\mathcal{X},\mathcal{J}|_{\mathcal{X}_{0}})\leq\mathop{\mathrm{lct}}\nolimits((\mathcal{X},\mathcal{X}_{0});\mathcal{J}),$
which follow from the inversion of adjunction or comparison of the adjoint ideal to the multiplier ideal (cf., e.g., [Eis11]), that matches to simple consequences of the normal volume comparison in Riemannian geometry under the control of Ricci curvatures from both sides. ∎
We consider the space of real valuations of $\mathcal{O}_{\mathcal{X},x}$ with the center $c_{\mathcal{X}}(v)=x$ and $v(t)=1$ and denote it by $\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$. We denote its subset of quasi-monomial valuations as $\mathop{\mathrm{Val}}\nolimits^{\rm qm,t\mapsto 1}(x\in\mathcal{X})$. If $\mathcal{X}$ is associated with a variety $\mathcal{X}_{\eta}$ over convergent power series ring $\mathbb{C}[[t]]^{\rm conv}(\subset\mathbb{C}[[t]])$, $\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$ is an open subset of its Berkovich analytification $\mathcal{X}_{\eta}^{\rm an}$ over the fraction field ${\rm Frac}(\mathbb{C}[[t]]^{\rm conv})=\mathbb{C}((t))^{\rm conv}$, so we have the weak topology on it. A subtle caution here is such valuation in $\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$ does not necessarily restrict to (unique) valuation at $x\in X$ in general.
Example 2.33 (Equisingular case).
If $\mathcal{X}=X\times\Delta$, it is natural to consider $\mathbb{C}^{*}$-invariant valuation i.e., valuation $v$ of the form $v(\sum_{i}f_{i}t^{i})=\min_{i}\{v_{X}(f_{i})+iw\}$ for some valuation $v_{X}$ of $\mathcal{O}_{X,x}$ with the center $\mathfrak{m}_{X,x}$ and a positive number $w$. Then it follows from a standard calculation to see $A_{(\mathcal{X},x)}(v)=A_{X}(v_{X})$, $\operatorname{vol}_{\mathcal{X}}(\mathcal{J}_{m}(v))=\operatorname{vol}(v_{X})$ so that $\widetilde{\operatorname{vol}}_{\mathcal{X}}(v)=\widehat{\operatorname{vol}}_{X}(v_{X})$. Hence the problem of minimizing $\widetilde{\operatorname{vol}}$ should be reduced to $\widehat{\operatorname{vol}}$ in the sense of [Li19].
3. Discussions on more general bubbles
3.1. Deeper bubbles for klt degenerations
For the structure formed by more general bubbles, we have rougher predictions. For convenience of discussions, we implicitly take a rescaling normalization of valuations as $v(t)=1$ in this section i.e., multiply $a$ to those considered in section 2.5.
Conjecture 3.1.
We fix an affine finite type log-terminal faithfully flat morphism $\pi\colon\mathcal{X}\to\Delta^{\rm conv}(\mathbb{C})$, where $\Delta^{\rm conv}(\mathbb{C})$ denotes the convergent spectrum $\mathop{\mathrm{Spec}}\nolimits(\mathbb{C}[[t]]^{\rm conv}:=\mathcal{O}_{\mathbb{A}^{1},0}^{\rm an})$, with $v(t)=1$. (Note that in particular, we have its restrictions of the form $\mathcal{Y}\to\Delta_{\epsilon}:=\{t\in\mathbb{C}\mid|t|<\epsilon\}$ for some $\epsilon>0$ with the same central fibers.) Then the following holds.
  1. ($\mathrm{i}$)
    Consider the rescaling equivalence classes of non-flat bubbles at $x\in\mathcal{X}$, which we write $\mathfrak{B}(x\in\mathcal{X})=\{(p\in B,g)\}/\mathbb{R}_{>0}$, and the class $(p\in B,g)$ represents as $[p\in B,g]$. There is a bijective correspondence between $\mathfrak{B}(x\in\mathcal{X})=\{(p\in B,g)\}/\mathbb{R}_{>0}$ and a subset $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})\subset\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$. We write the latter as $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})$ and the bijective maps as
    $\nu\colon$ $\mathfrak{B}(x\in\mathcal{X})\xrightarrow{\,\smash{\raisebox{-1.50696pt}{$\scriptstyle\simeq$}}\,}\mathop{\mathrm{Crit}}\nolimits(x\in X)\subset\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X}),$
    $B\colon$ $\mathop{\mathrm{Crit}}\nolimits(x\in X)\xrightarrow{\,\smash{\raisebox{-1.50696pt}{$\scriptstyle\simeq$}}\,}\mathfrak{B}(x\in\mathcal{X}),$
    which are inverse to each other. In particular, the set of bubbles is embedded into a non-archimedean Berkovich analytic space.
  2. ($\mathrm{ii}$)
    $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})\subset\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$ and compatibly the moduli of bubble classes $\mathfrak{B}(x\in\mathcal{X})$ has a locally PL structure.
  3. ($\mathrm{iii}$)
    $\mathfrak{B}(x\in\mathcal{X})$ and $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})$ only depend on the germ of $\pi$ around $x$.
To make sense of the item ($\mathrm{ii}$), we prepare the following after the general ideas in [Berk99, KS06, BFJ08, BFJ16].
Lemma 3.2.
In the above setup, we further assume $\mathcal{X}$ descends to the family over $\Delta$ i.e., $\mathcal{X}=\mathcal{Y}$ for simplicity. Consider the projective system $(S,>)$ of isomorphism classes of $\varphi\colon\tilde{\mathcal{X}}\to\mathcal{X}$ which run over birational projective morphism such that $E_{x}$ is a divisorial part of $\varphi^{-1}(x)$ and $(\tilde{\mathcal{X}},\tilde{\mathcal{X}}_{0}+E_{x})$ is dlt and $\mathbb{Q}$-factorial, so that the dual intersection complex $\Delta(E_{x})$ is well-defined. (In case $x\in X$ is an isolated singularity, we can instead put stronger assumptions $\tilde{\mathcal{X}}$ is smooth and $\tilde{\mathcal{X}}_{0}+E_{x}$ is simple normal crossing for simplicity.) $(\varphi_{1}\colon\tilde{X}_{1}\to\mathcal{X})>(\varphi_{2}\colon\tilde{X}_{2}\to\mathcal{X})$ is simply defined for domination by a commuting birational morphism $\mathcal{X}_{1}\to\mathcal{X}_{2}$.
  1. ($\mathrm{i}$)
    For each $\tilde{\mathcal{X}}\to\mathcal{X}$ in $S$, consider $\Delta(E_{x})$, the dual intersection complex (of the largest normal crossing open subset of $E_{x}$). Then there is a natural injective continous map $\Delta(E_{x})\hookrightarrow{\rm Val}^{\rm qm,t\mapsto 1}(x\in\mathcal{X})$. Through this map, we identify $\Delta(E_{x})$ as a subset of ${\rm Val}^{\rm qm,t\mapsto 1}(x\in\mathcal{X})$.
  2. ($\mathrm{ii}$)
    Further, we have the following structure:
[…]
    (13) $\subset\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$
[…]
    (15) $\subset\mathcal{X}_{\eta}^{\rm an},$
    where $\varphi\in S$ and the domination “$>$” induces PL maps as in loc.cit. In particular, (14) is a dense inclusion while (15) is not.
Proof.
[…]
3.2. Towards bubbling for log canonical degeneration
We end our paper by making some further discussions towards understanding of bubbling along worse degenerations than log terminal singularities, following our previous works ([Od22a, Od22b]). This subsection is independent of other (sub)sections. As a common setup of diverging or collapsing Ricci-flat Kähler metrics in the literature, one often consider the cases when the section hits a non-log-terminal, but strictly (semi-)log canonical point, like a non-smooth point at simple normal crossing degeneration. In such a situation, natural metric limits are expected to collapse. Therefore, to extend our approach, one can instead naturally ask the existence of minimal non-collapsing bubble (with infinite diameter) as defined in [Od22a, Definition 4.2] if it exists. Nevertheless the situation is more intricate (cf., e.g., [Od22a, 4.4], [HSVZ22, SRZ23]) and far beyond the main setup of this paper.
Here, as we expect it to play some roles in such general bubbling problem, we first review a simpler local version ([Od22b, Definition 2.5, Proposition 2.6]) of the CM degree minimization ([Od12, §4]) as follows. The following is a variant of weight function by [KS06, MN15, Tem14].
Definition 3.3 (Log canonical height).
For a projective flat generically K-trivial191919only means the canonical divisor of the generic fiber is numerically trivial family over a smooth proper curve $\mathcal{X}\to C\ni 0$ and a holomorphic section $\sigma\colon C\to\mathcal{X}$, where $\mathcal{X}$ is normal and $\mathbb{Q}$-factorial, we define
$h_{\rm lc}(\sigma):=\operatorname{deg}_{C}\hskip 2.84526pt\sigma^{*}(\mathcal{O}_{\mathcal{X}}(K_{\mathcal{X}/C}+\mathcal{X}_{0,{\rm red}})),$
and call it the log canonical height of $\sigma$. Here, $\mathcal{X}_{0,{\rm red}}$ refers to the reduced scheme structure on $\mathcal{X}_{0}$.
Though we work over a proper base $C$, note that it is only a matter of formal convenience to make sense of the degree. Indeed, along any vertical birational transforms of $\mathcal{X}$, one can always make sense of the difference even over non-proper $C$. Simply for this convenience, we keep the proper base setup. The following observation tells us the importance of log canonical minimal models.
Proposition 3.4.
For a flat proper family of log terminal projective K-trivial varieties $\mathcal{X}^{*}\to(C\setminus\{0\})$, we fix the meromorphic section $\sigma|_{C\setminus 0}$. Here, $C$ is a smooth proper curve. We assume $\mathcal{X}^{*}$ is $\mathbb{Q}$-factorial. Then among all fillings i.e., proper flat $\mathcal{X}\to C$ possibly after a finite base change of $C$ of degree $d$, we consider the quantity
$h_{lc}^{\nu}(\sigma):=\frac{h_{lc}(\sigma)}{d}.$
Then, $h_{lc}^{\nu}(\sigma)$ is minimized if and only if $\mathcal{X}$ around $\sigma(0)$ is a germ of some relative log canonical minimal model $(\mathcal{X},\mathcal{X}_{0,{\rm red}})$ over $C$.
Proof.
Firstly, if $(\mathcal{X},\mathcal{X}_{0,{\rm red}})$ is not log canonical around $\sigma(0)$, by the log canonical modification ([OX12]) $h_{\rm lc}(\sigma)$ would decrease by the negativity lemma. Therefore, we can and do assume $(\mathcal{X},\mathcal{X}_{0,{\rm red}})$ is at least log canonical locally around $\sigma(0)$. We replace $(\mathcal{X},\mathcal{X}_{0,{\rm red}})$ by a log canonical pair, keeping isomorphism around $\sigma(0)$ by [HX13], and run the relative minimal model program with ample $\mathbb{Q}$-Cartier scaling $\mathcal{H}$ to obtain the relative log canonical models of $(\mathcal{X},\mathcal{X}_{0,{\rm red}}+a\mathcal{H})$, by [HH20, Theorem 1.5]. We can assume $\sigma(0)$ is not contained in $\mathcal{H}$. If the corresponding divisorial contraction or log canonical flip is not locally isomorphism around $\sigma(0)$, then the log canonical height will decrease by the negativity lemma again. ∎
Recall that among the relative log canonical minimal models showed up above, there are dlt minimal models which are more special and simpler. Further, there is the technique of replacing the log canonical minimal model by its blow up which is dlt minimal model, in a log crepant manner, the so-called (log crepant) dlt modification or dlt blow up (cf., e.g., [Fuj11, KK10, OX12]). Thus, it sounds very natural to consider dlt minimal models to study these general bubbles, which led to the notion of “galaxy models” ([Od22b, §2]): connected locally ringed space which contains infinitely many bubbling log Calabi-Yau varieties densely.
Acknowledgement
The author heartily thanks S.Sun and C.Spotti for the helpful conversations over many years on this topic. He also thanks S.Boucksom, S.K.Donaldson, D.Halpern-Leistner, M.Jonsson, M.Mauri and T.Nghiem for helpful comments and discussions. During the research, the author was partially supported by Grant-in-Aid for Scientific Research (B) 23H01069, 21H00973, Grant-in-Aid for Scientific Research (A) 21H04429, 20H00112, and Fund for the Promotion of Joint International Research (Fostering Joint International Research) 23KK0249. Part of this work is done during the author’s stay at Paris, notably Institut de mathématiques de Jussieu - Paris Rive Gauche from March 2024 and Imperial College London in June 2024. The author received there above-mentioned benefitial comments, and nice pleasant working environments. We appreciate their kind hospitality.