### v2

# Algebraic geometry of bubbling Kähler metrics

Abstract.
We give an algebro-geometric or non-archimedean framework to study bubbling phenomena of Kähler metrics with Euclidean volume growth, after [DS17, Sun23, dBS24]. In particular, for any degenerating family to log terminal singularity, we give an algorithm to algebraically construct a finite sequence of birational modifications of the family with milder degenerations, and partially relate with analytic bubbling constructions in [Sun23, dBS24].
On the way, we also introduce and discuss a global analogue which in particular reduces compactness of the moduli of compact Kähler-Ricci solitons, again to a finite generation type problem in birational geometry. Our whole discussion depends on the general framework of stability theory for irrational type degenerations in [Od24b].
1. Introduction
We make a purely algebro-geometric study of degenerations of normal singularities in this paper, motivated by the recent works on asymptotic behaviour of Kähler metrics [DS17, Sun23, dBS24]. Our main setup is as follows.
Setup 1.
Let $\pi\colon\mathcal{X}\twoheadrightarrow\Delta\ni 0$, be either one of them:
  1. (a)
    (analytic setup over $\mathbb{C}$) $\Delta:=\{t\in\mathbb{C}\mid|t|<1\}$, $\pi$ is a flat klt (kawamata-log-terminal) affine 111meaning that $\pi$ is achievable as a restrction of ${\rm pr}_{1}\colon(\Delta\times\mathbb{C}^{l})\twoheadrightarrow\Delta$ for $l\in\mathbb{Z}_{>0}$ to an analytic closed subset of $(\Delta\times\mathbb{C}^{l})$ map of relative dimension $n$ over the base field $\mathbbm{k}=\mathbb{C}$ i.e., each fiber is a $n$-dimensional log terminal complex variety.
  2. (b)
    (algebraic setup over a field of characteristic 00) $\Delta$ is an affine curve over a field $\mathbbm{k}$ of characteristic 00, $\pi$ is a flat klt affine morphism between $\mathbbm{k}$-varieties of relative dimension $n$.
Occasionally, to make the base coordinate $t$ clear, we denote $\Delta$ as $\Delta_{t}$. We also assume $K_{\mathcal{X}}$ to be $\mathbb{Q}$-Cartier as our main interest comes from Kähler-Einstein geometry. We set $X:=\mathcal{X}_{0}:=\pi^{-1}(0)$. We further fix a holomorphic (resp., algebraic) section $\sigma$ with $\sigma(0)=x$.
In the next section, we algebraically construct (almost) canonical, finite ordered sequences of modifications of $\mathcal{X}$ with milder degenerations. For instance, if the general fibers are $2$-dimensional smooth, degenerating to $X$ with a $2$-dimensional $A_{k}$-singularity, then we obtain other fillings with finite number of $A_{i}$-singularities with smaller Milnor number $i$ (cf., [dBS24]) and terminate at smooth filling. The process is, weakly related to but substantially different from the classical simultaneous resolution of Brieskorn-Grothendieck type in an essential manner. Morally speaking, our algorithm gives canonical “limit singularities” of various rescales beyond the given $X$ itself or its (metric) tangent cone, reflecting the whole formation process of the singularities of $X$, in the context of Kähler-Einstein geometry and the K-stability. While we explain its detailed meaning later, here is the brief overview of the differential geometric background.
Suppose we have a “good” continuous family of (singular) Kähler metrics $g_{t}$ on $\mathcal{X}_{t}=\pi^{-1}(t)$ in the pluri-potential theoritic sense (cf., e.g., [BBEGZ19]), for $t\neq 0$, especially of Kähler-Einstein metrics, we would like to consider and answer the following question:
Main Question (see [Sun23, dBS24]).
How can we determine the set of possible non-collapsing pointed Gromov-Hausdorff limits of $(x_{t}\in\mathcal{X}_{t},c_{t}^{2}g_{t})$ for $t\to 0$ where $c_{t}>0$ are some sequences of diverging rescaling parameters?
In this paper, following [Sun23, dBS24], we call such a non-collapsing pointed Gromov-Hausdorff limit simply as bubbles of the family $(\mathcal{X}_{t},g_{t})$ for $t\to 0$. By the groundbreaking work [DS17, 1.4], extending some classical cases, such limits often naturally admits structure of an affine log terminal variety with (weak) Ricci-flat Kähler metric. Indeed, our aforementioned main construction of affine varieties is expected to underlie these bubbles. They have Euclidean volume growth simply by the Bishop’s inequality. This is a (Kähler) metric version of so-called bubbling phenomenon since [SU81], as later also studied by [Nak88, And89, Ban90, BKN89] in mainly $4$-dimensional Riemannian geometric setup, which in particular found out the finite energy (the $L^{2}$-norm of the curvature) rescaling provides smooth asymptotically locally Euclidean (ALE) manifolds limits as the deepest bubbles. Then, recently we witnessed a breakthrough [DS17, Sun23, dBS24] in the higher dimensional Kähler setup. In particular, [Sun23] proved the existence of “minimal” bubble in a differential geometric sense, finiteness of bubbles, and raised Conjecture 5.4 in loc.cit to expect a $2$-step algebraic construction of the underlying varieties of the minimal bubbles, and [dBS24] showed many examples and a refinement of the conjecture.
Motivated by loc.cit among others, this paper aims to provide a systematic algebro-geometric or non-archimedean approach to the problem.
The main contents of this paper give answer the above main question by solving large parts of the conjectures of [Sun23, dBS24] with some refinements. The following is a brief outlook of our statements for our algorithm, but see the detailed statements in later referred subsections.
Theorem 1.1.
Consider an arbitrary flat degeneration of pointed $n$-dimensional log terminal (hence normal) singularities $\pi\colon\mathcal{X}(\supset\sigma(\Delta))\to\Delta\ni 0$ (see Setup 1), which is strictly degenerating in a certain sense (see Theorem 2.4). Then, the following hold.
  1. ($\mathrm{i}$)
    (Theorem 2.4) There is a purely algebro-geometric $2$-step construction 222the following symbol $\rightsquigarrow$ means each step of the construction, and is not a map. of algebro-geometric version of minimum bubbles as affine varieties
    $(X\subset\mathcal{X})\rightsquigarrow\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow\mathcal{X}^{\prime\prime}_{\rm min,0}.$
    Here, $\mathcal{X}^{\prime}_{\rm min,0},\mathcal{X}^{\prime\prime}_{\rm min,0}$ are both still $n$-dimensional affine log terminal varieties. The minimality morally means the mildness of singularity measured by the local normalized volumes (at the base points) as we clarify later. It is also parallel to minimality of the diverging order of the rescaling constants $\{c_{t}\}_{t}$ (cf., also [Sun23] and Definition 2.1).
    We call $\mathcal{X}^{\prime}_{\rm min,0}$ (resp., $\mathcal{X}^{\prime\prime}_{\rm min,0}$) (algebraic) semistable minimal bubble (resp., (algebraic) polystable minimal bubble) for simplicity. 333[Sun23] (resp., [dBS24]) conjectured such construction and called the semistable minimal bubble $\mathcal{X}^{\prime}_{\rm min,0}$ under different names as algebraic minimal bubble (resp., (minimal) weighted bubble). Note that our construction of $\mathcal{X}^{\prime\prime}_{\rm min,0}$ is also algebraic.
  2. ($\mathrm{ii}$)
    (Theorem 2.4, Theorem 2.11, Lemma 2.25) The first step construction
    $X\subset\mathcal{X}\rightsquigarrow\mathcal{X}^{\prime}_{\rm min,0}$
[…]
We refer to these obtained (non-canonical) total spaces of the new degenerations to $\mathcal{X}^{\prime}_{\rm min,0}$ and $\mathcal{X}^{\prime\prime}_{\rm min,0}$ respectively as $\mathcal{X}^{\prime}_{\rm min}$ and $\mathcal{X}^{\prime\prime}_{\rm min}$.
By replacing $\mathcal{X}$ by $\mathcal{X}^{\prime\prime}_{\rm min}$ and do iteration of the above procedure ($\mathrm{i}$), we obtain deeper modifications of $\mathcal{X}$.
This whole process can be seen as a certain algorithm of improving the degenerations with some differential geometric motivation. Its double iterative nature may make look too complicated or even artificial, but it is unavoidable and indeed parallels the actual bubbling structure and hence we believe it is the nature of the arts in general. Nevertheless, it becomes simpler in many classical examples. For instance, in the $1$ or $2$-dimensional classical cases considered in [dBS24, §2.3, 3.3, 4.2], the second step of ($\mathrm{i}$) each time is trivial so that $\mathcal{X}^{\prime}_{\rm min,0}=\mathcal{X}^{\prime\prime}_{\rm min,0}$ recovers the differential geometric construction i.e., a certain pointed Gromov-Hausdorff limit (at the underlying affine varieties level) and the above construction of $\mathcal{X}^{\prime}_{\rm min,0}$ is canonical. Thus by iterations, we obtain the whole set of possible differential geometric bubbles along $\mathcal{X}\supset\sigma(\Delta)$ with the local metrics given in loc.cit. See Theorem 2.13.
Theorem 2.6 gives a subtle technical improvement of the above statements, including the following points. Firstly, we allow degenerations along more general bases. As a related advantage, the $1$-st step construction of the total space $\mathcal{X}^{\prime}_{\rm min}$ can be achieved in a more canonical manner i.e., as a more canonical birational modification of the original family, once we pass to a possibly non-finite type base change i.e., a certain Novikov type ring spectrum. To make sense of it, we use [Od24b] in which some algebraic (partly stack-theoretic) generalities to treat degenerations in “irrational direction” with higher dimensional algebraic base is prepared. We also compare our construction with more analytic and differential geometric constructions in [Sun23].
Since our construction is fairly general and look abstract first, it is designed after various explicit calculations as experiments. We do not write all the calculations but review one of the simplest examples after [dBS24] here. In this case, we only need a single process rather than the $2$-step processes i.e., $\mathcal{X}^{\prime}_{\rm min}=\mathcal{X}^{\prime\prime}_{\rm min}$, and they are achieved only after a finite base change of $\Delta$.
Example 1.2 (cf., [BKN89, Ban90], [dBS24, §3.3] and Theorem 2.13).
Consider a degeneration of pointed affine smooth surfaces to a $A_{3}$-singular surface as $\mathcal{X}\subset\Delta_{t}\times\mathbb{C}^{3}_{x,y,z}$ cut out by the equation
(1) $xy=z(z-t)(z-t^{2})(z-t^{3})$
whose base points are given by the 00-section $\sigma$. Here, the subindices $x,y,z,t$ are coordinates.
We can consider a family of hyperKähler metrics on each $t$ by the Gibbons-Hawking ansatz with the base $\mathbb{R}^{3}$ and monopole points $x_{1},x_{2},x_{3},x_{4}$ and harmonic functions of the form $\frac{1}{2}\sum_{i}\frac{1}{|x-x_{i}|}$. In this case, [dBS24, §3.3] describes how it bubbles at $t\to 0$ around the origin, and coincides with the algebraic procedure Theorem 1.1, and is explicit as follows. (Indeed, our general construction is originally obtained as a generalization of their examples).
Firstly, we consider the weighted blow up of $(0,0,0,0)\in\mathcal{X}$ with weights $(2,2,1,1)$ and removal of the strict transform of the central fiber $\mathcal{X}_{0}$, which gives $\mathcal{X}^{\prime}_{\rm min}=\mathcal{X}^{\prime\prime}_{\rm min}$ with the equation
(2) $xy=z(z-1)(z-t)(z-t^{2}),$
after replacing $\frac{x}{t^{2}}$ as $x$, $\frac{y}{t^{2}}$ as $y$, $\frac{z}{t}$ as $z$ while keeping $t$. This is a degeneration to $xy=z^{3}(z-1)$ when $t\to 0$, with $A_{2}$-singularity at $(0,0,0,0)\in\Delta_{t}\times\mathbb{C}_{x,y,z}^{3}$. This admits a ALE hyperKähler orbifold metric, again obtained via the Gibbons-Hawking ansatz over $\mathbb{R}^{3}$ (cf., [dBS24, §3.3]).
Then after the base change with respect to the double covering $\Delta_{\sqrt{t}}\to\Delta_{t}$, we replace $x,y,z,t$ by $\frac{x}{t^{1.5}},\frac{y}{t^{1.5}},\frac{z}{t},t^{0.5}$ and still denote it by $x,y,z,t$ for simplicity of notation. Then, the weighted blow up of the base changed family with weights $(3,3,2,2)$ (again with the strict transform of the central fiber removed) is defined by
(3) $xy$ $=z(t^{2}z-1)(z-1)(z-t^{2})$
(4) $=(t^{2}z-1)(z-1)\cdot z(z-t^{2}),$
which is a degeneration to a ALE hyperKähler orbifold bubbling with a $A_{1}$-singularity where the metric is infinitesimally $\mathbb{C}^{2}/\pm$.
As the final step, we again do a weighted blow up with weights $(2,2,2,1)$ for the coordinates $(x,y,z,t)$ and remove the strict transform of the central fiber. Then, by replacing the coorindates by new $(\frac{x}{t^{2}},\frac{y}{t^{2}},\frac{z}{t^{2}},t)$ which we denote again by $(x,y,z,t)$, we obtain the new equation
(5) $xy=(t^{4}z-1)(t^{2}z-1)\cdot z(z-1),$
inside $\Delta_{t}\times\mathbb{C}^{3}_{x,y,z}$. Note that this is finally a smooth family and we have an asymptotically conical hyperKähler metric of Eguchi-Hanson type at the central fiber.
As we clarify, in general, we need more complicated processes due to possible irregularity of the corresponding Sasaki-Einstein link i.e., irrationality of its Reeb vector field.
We make this kind of explicit study as Example 1.2 via coordinates more generally. That is, after the proof of Theorem 1.1 ($=$Theorems 2.4, 2.11), we propose more explicit ways of constructions of the candidates of bubbles, via coordinates/equations and valuations respectively in §2.4, §2.5. We conjecture that they coincide with the construction of Theorem 2.4. §2.2.5 and §2.3 may be of own independent interest, which are respectively the systematic treatment of negative weight deformation originated in [CH22, SJZ23] and an analogue of bubbling Calabi-Yau metric in the context of normalized Kähler-Ricci flow on Fano varieties. The latter would give a yet another approach to the properness of K-moduli of Fano varieties cf., Theorem 2.22. Further, in §3, we discuss more on the structure on the “bubble tree” formed by all possible bubbles.
From some personal perspectives, this paper is also motivated by hope to deepen our understanding of the projective (resp., quasi-projective) K-moduli of Fano varieties (resp., Calabi-Yau varieties) in the sense of [Od10, 5.2], after recent completion of the construction (cf., e.g., [BL21, BHLLX21, LXZ22] resp., [Bir22, Od21]). Recall that those K-moduli were constructed and understood before loc.cit equivalently as Gromov-Hausdorff (partial) compactifications from more differential geometric viewpoints (cf., e.g., [OSS16, SSY16, Od15a, LWX19, Bir22, Od21]), following the breakthrough work of [DS14] and the aforementioned bubblings give a deepening of our understanding on it.
2. Minimal bubble
2.1. Reviewing definitions
We work in Setup 1 in the introduction. In this section, we give algebraic approaches to the “minimum bubble” in the sense of [Sun23, dBS24], whereas the minimality roughly refers to that of the order of rescaling parameters $c_{t}$ for $t\to 0$. We review the definition for convenience.
Definition 2.1 ([Sun23, Definition 1.9]).
In the context of Setup 1 (a), its (differential geometric) minimal bubble refers to the pointed Gromov-Hausdorff limits of $(\sigma(t)\in\mathcal{X}_{t},c_{t}^{2}g_{t})$ for certain diverging sequence $c_{t}>0$, which is not a Riemannian metric cone, but whose tangent cone at infinity ([DS17, §3.4]) is the metric local tangent cone at $\sigma(0)\in X$, which we denote as $C_{\sigma(0)}(X)$.
Firstly, we discuss several a priori different algebraic approaches to it constructions, in different subsections, and will discuss the conjectural equivalence later. We crucially use the theory of local normalized volume of log terminal singularities due to [Li19, Liu18] (cf., also [Od15b]), with its origin in [DS17]. We briefly review the definition and properties we use here. Its smooth case can be traced back to much earlier [Mus02, dFEM04] (and also related [ELS03]) and more detailed review are available at [LLX20], [Od24a, §2] for instance.
Definition 2.2 ([ELS03]).
[…]
Then, the normalized volume $\widehat{{\rm vol}}_{x}(v)$ is defined as $A_{X}(v)^{n}{\rm vol}_{x}(v)$ by [Li19], where $A_{X}(v)$ is the log discrepancy in the generalized sense by [JM12]. Then the following is known by now. We do not make any substantial contribution in this paper but rather use in this paper. See more references in [LLX20], [Od24a, §2]. Morally speaking, this process is to modify a “locally K-stable” i.e., log terminal singularity (cf., [Od15b, GT23]) to cone, which is still (K-semi/poly)stable.
Theorem 2.3 (Scale up conification cf., [DS17, Li19, Liu18, Blu18, HS17, LX18, LWX21, LLX20, XZ21, Od24a]).
For any log terminal singularity $x\in X$, there is a unique quasi-monomial valuation $v=v_{X}$ with the center $x$ up to the $\mathbb{R}_{>0}$-action, which minimizes the normalized volume $\widehat{\operatorname{vol}}(-)$.
  1. ($\mathrm{i}$)
    Then, the valuation $v_{X}$ coincides with the vanishing order function in [DS17, §3.2] for the local Kähler metric in the setup of loc.cit. We denote the rational rank of $v_{X}$ as $r$. Further, $\widehat{\rm vol}(v_{X})/n^{n}$ has a differential geometric meaning as the volume density of the link of the local metric tangent cone. We denote the minimized value $\widehat{\rm vol}(v_{X})$ as $\widehat{\rm vol}(x\in X)$ or $\widehat{\rm vol}_{X}(x)$, in this paper.
  2. ($\mathrm{ii}$)
    From more algebraic viewpoint, the graded ring
    ${\rm gr}_{v}(\mathcal{O}_{X,x}):=\oplus_{s\in{\rm Im}(v_{X})}(\{f\in\mathcal{O}_{X,x}\mid v(f)\geq s\}/\{f\in\mathcal{O}_{X,x}\mid v(f)>s\})$
    is a finite type $\mathbbm{k}$-algebra on which there is a natural action of the algebraic $\mathbbm{k}$-torus $T\simeq\mathbb{G}_{m}^{r}$ on a K-semistable Fano cone $W:=\mathop{\mathrm{Spec}}\nolimits({\rm gr}_{v}(\mathcal{O}_{X,x})$.
  3. ($\mathrm{iii}$)
    For a certain rational polyhedral cone $\tau$ of the dual of the groupification of ${\rm Im}(v_{X})$, there is a positive weight degeneration of $X$ to $W$, over $U_{\tau}$ which is a $\mathbb{Q}$-Gorenstein family.
  4. ($\mathrm{iv}$)
    There is a (non-canonical) affine test configuration of $W$ to a unique K-polystable Fano cone $C$ with a (weak) Ricci-flat Kähler cone metric. In the setup of [DS17, §2-§3.3], it recovers the local metric tangent cone of local Kähler metric (as conjectured in loc.cit).
We note that the last $\mathbb{Q}$-Gorensteinness assertion of the positive weight degeneration in ($\mathrm{iii}$) seems missing in literature but can be confirmed as follows. As in [Od24b, §2], take large enough homogeneous generating system $\overline{z}_{1},\cdots,\overline{z}_{l}$ of ${\rm gr}_{v}(\mathcal{O}_{X,x})$ with degrees $w_{1},\cdots,w_{l}$, and lift to $\Gamma(\mathcal{O}_{X})$ as $\overline{z}_{1},\cdots,\overline{z}_{l}$ which we suppose to give an embedding $W\hookrightarrow\mathbb{A}_{\mathbbm{k}}^{l}$ (otherwise add more $\overline{z}_{i}$s). Then we take rational approximation $\xi^{\prime}$ of $\xi:=(w_{1},\cdots,w_{l})$ and corresponding positive weight test configuration as $W_{\xi^{\prime}}$ over $\mathbb{A}^{1}$. For any $\xi^{\prime}$ close enough to $\xi$, we see that the central fiber is again $W$ just with a perturbed vector field induced on it. Note that there is a natural relative compactification $W_{\xi^{\prime}}\subset\overline{W_{\xi^{\prime}}}\to\mathbb{A}^{1}$ by adding the quotient by $\mathbb{G}_{m}$ for $\xi^{\prime}$-direction, which is automatically a $\mathbb{Q}$-Gorenstein family by the arguments of [OSS16, (proof of) Lemma 2.4]. Since this holds for any $\xi^{\prime}$ close enough to $\xi$, the map from the toric base $U_{\tau}$ of the positive weight degeneration is in the $\mathbb{Q}$-Gorenstein deformation component so that we obtain the assertion.
2.2. Construction via deformation theory
We use theorems in [Od24b] after [Lan75, HL14, AHLH23] to algebraic construction of the conjectural (minimal) bubble and discuss it. We also discuss reformulation of negative weight deformations in the sense of [CH22], after [Od24b], in which we also called it scale down deformation.
2.2.1. Setup and algebraic construction
In the following two subsubsections, we set up stages and list some statements in more precise form we prove in the later subsubsections.
Setup 2.
In the context of Setup 1 (a) (resp., Setup 1 (b)), as Example 2.8 in [Od24b], denote by $v_{X}$ the valuation of $\mathcal{O}_{X,x}$ whose center is $x$ and minimizes the normalized volume $\widehat{\operatorname{vol}}(-)$ ([DS17, Li19, Blu18, XZ21]).
We set the groupification of ${\rm Im}(v_{X})$ as $N\simeq\mathbb{Z}^{r}$, $M:=N^{*}:={\rm Hom}_{\mathbb{Z}}(N,\mathbb{Z})$ denotes the dual lattice, and set $T:=N\otimes\mathbb{G}_{m}\simeq\mathbb{G}_{m}^{r}$, where $r$ is the $\mathbb{Q}$-rank of $v_{X}$.
For some technical reasons of our logics, i.e., to use Kollár’s hull and husk construction ([Kol08]) and also for precision of “good” family or singular limit of (weak) Kähler metrics later, we will henceforth sometimes assume the family $\pi$ can be compactified to or comes from a family of projective varieties. Note that it is also assumed in [DS14] (and [DS17]). We believe such assumption above will be unnecessary in the end (see Conjecture 3.1 ($\mathrm{iii}$)). That is, we temporary assume:
Setup 3.
In the contexts of Setup 1 $\land$ 2, we further assume $\pi\colon\mathcal{X}\to\Delta$ is compactifiable to a polarized projective family $\overline{\pi}\colon\overline{\mathcal{X}}\to\Delta$ such that $\overline{\mathcal{X}}_{t}=\overline{\pi}^{-1}(t)$ are all log terminal for any $t\in\Delta$ and $K_{\overline{\mathcal{X}}}$ is again $\mathbb{Q}$-Cartier.
The following main theorems 2.4, 2.6, 2.11 aim at algebraic reconstruction and generalization of the bubbles via metrics in [DS17, Sun23]. The first version below (Theorem 2.4) avoids the new terminology of [Od24b] for the readers convenience but the next one (Theorem 2.6), which uses loc.cit, will be more general.
Theorem 2.4 (Algebro-geometric construction of (minimal) bubbles).
We work under the Setup 1 (a) (resp., 1 (b)) $\land$ 2 $\land$ 3, and we further assume that the singularity of $\sigma(t)\in X_{t}$ strictly degenerates in the sense that
$\widehat{\operatorname{vol}}_{\mathcal{X}_{0}}(\sigma(0))<\widehat{\operatorname{vol}}_{\mathcal{X}_{t}}(\sigma(t))$
for any $t\neq 0$. Here $\widehat{\operatorname{vol}}(-)$ denotes the normalized volume 444also called under different names in the literature such as volume density ([SS17, HS17]) up to a constant, or local volume in [XZ24] (although different from the local volume of [Ful11]) in the sense of [Li19] (see §2.1 or [LLX20], [Od24a, §2] in more details).
Then, possibly after replacing the base $\Delta\ni 0$ of $\pi$ by its finite base change, there exist a finite ordered set of pairs of birational transforms $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}\to\Delta$ and $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}\to\Delta$ which are both isomorphism over $\Delta\setminus\{0\}$ such that all the following conditions hold. We denote the projection $\mathcal{X}^{\prime}\to\Delta$ as $\pi^{\prime}$, and $\mathcal{X}^{\prime\prime}\to\Delta$ as $\pi^{\prime\prime}$ respectively.
Here are the conditions:
  1. ($\mathrm{i}$)
    $\pi^{\prime}\colon\mathcal{X}^{\prime}\to\Delta$ and $\pi^{\prime\prime}\colon\mathcal{X}^{\prime\prime}\to\Delta$ are affine faithfully flat morphisms and again (relatively/fiberwise) log terminal and they are both $\mathbb{Q}$-Gorenstein families in the sense that $K_{\mathcal{X}^{\prime}}$ and $K_{\mathcal{X}^{\prime\prime}}$ are both $\mathbb{Q}$-Cartier.
  2. ($\mathrm{ii}$)
    $\sigma|_{(\mathcal{X}\setminus\{0\})}$ extends to a holomorphic section $\Delta\to\mathcal{X}^{\prime}$ which we denote by $\sigma^{\prime}$ (resp., a holomorphic section $\Delta\to\mathcal{X}^{\prime\prime}$ which we denote by $\sigma^{\prime\prime}$).
  3. ($\mathrm{iii}$)
    Fix one of the pairs $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$. For the central fibers $\mathcal{X}^{\prime}_{0}$ of $\mathcal{X}$ (resp., $\mathcal{X}^{\prime\prime}_{0}$ of $\mathcal{X}^{\prime\prime}$), we have
    (*) $\widehat{\operatorname{vol}}_{\mathcal{X}^{\prime}_{0}}(\sigma^{\prime}(0))=\widehat{\operatorname{vol}}_{\mathcal{X}^{\prime\prime}_{0}}(\sigma^{\prime\prime}(0))>\widehat{\operatorname{vol}}_{\mathcal{X}_{0}}(\sigma(0)),$
[…]
  4. ($\mathrm{iv}$)
    The central fiber $\mathcal{X}^{\prime}_{0}:=\pi^{\prime-1}(0)$ of each $\mathcal{X}^{\prime}\ni\sigma^{\prime}_{0}(x)$, is an algebraic negative weight deformation ([Od24b, §2], after [CH22, Definition 1]) of a K-semistable $\mathbb{Q}$-Fano cone $(W,\xi)$ (in the sense of [CS18]) over $\mathbb{A}_{\mathbbm{k}}^{1}$ in a canonical manner, up to a finite base change, which we denote simply as $\mathcal{W}\twoheadrightarrow\mathbb{A}_{\mathbbm{k}}^{1}$ or $\mathcal{X}^{\prime}_{0}\rightsquigarrow W$. In particular, $\mathcal{X}^{\prime}_{0}:=\pi^{\prime-1}(0)$ is K-semistable in the sense of [Sun23, §5]. The total space $\mathcal{X}^{\prime}_{\rm min}$ is achieved as an elementary modification of $\mathcal{X}$ in the sense of [HL14, AHLH23].
  5. ($\mathrm{v}$)
[…]
  6. ($\mathrm{vi}$)
    Suppose $W=X$ and they have only isolated singularity with unobstructed deformations, and $\mathcal{X}$ is weakly non-degenerating (compare [BR15]) in the sense we reveal during the proof. For instance, if it holds for $n=2$ ADE singular case.
    Then, the central fiber $\mathcal{X}^{\prime}_{{\rm min},0}$ of $\mathcal{X}^{\prime}_{\rm min}$ can be constructed canonically (cf., also Theorem 2.11, Theorem 2.13 for partial comparison results with differential geometric bubbles [Sun23, dBS24]).
  7. ($\mathrm{vii}$)
    The central fiber $\mathcal{X}^{\prime\prime}_{0}:=\pi^{\prime-1}(0)$ of $\mathcal{X}^{\prime\prime}$ is an algebraic negative weight deformation 777 (See also discussions in [SRZ23, §6], which is related to [CH22, 1.7], though the stated existence of negative valuation is a weaker statement). ([Od24b, §2], after [CH22]) of a K-polystable $\mathbb{Q}$-Fano cone $(C,\xi_{C})$ in the sense of [CS18, CS19]. In particular, $\mathcal{X}^{\prime\prime}_{0}:=\pi^{\prime\prime-1}(0)$ (together with the corresponding negative valuation) is K-polystable in the sense of [Sun23, §5].
  8. ($\mathrm{viii}$)
[…]
    Further, the isotrivial degeneration is again a negative weight deformation999 (See the footnote 6 above). in the sense of [CH22, Definition 1.7 and Appendix I]. In particular, $\mathcal{X}^{\prime\prime}_{0}:=\pi^{\prime\prime-1}(0)$ (together with the corresponding negative valuation) is K-polystable in the sense of [Sun23, §5].
  9. ($\mathrm{ix}$)
    For each of the pairs $(\mathcal{X}^{\prime},\mathcal{X}^{\prime\prime})$, there is an affine $\mathbb{Q}$-Gorenstein test configuration of $\mathcal{X}_{0}^{\prime}$ degenerating to $\mathcal{X}_{0}^{\prime\prime}$. If $n=2$, this is the trivial degeneration i.e., $\mathcal{X}^{\prime}=\mathcal{X}^{\prime\prime}$.
  10. ($\mathrm{x}$)
    If the rational rank $r$ of $v_{X}$ is $1$, $\mathcal{X}^{\prime}_{\rm min,0}$ gives a unique divisorial valuation of $\mathcal{X}$ centered at $x$, which we call minimal critical (divisorial) valuation.
  11. ($\mathrm{xi}$)
    If we fix a test configuration of $W=\mathop{\mathrm{Spec}}\nolimits(gr_{v_{X}}(\mathcal{O}_{X,x}))\rightsquigarrow C_{\sigma(0)}(X)$, which we will denote as $\mathcal{T}^{\prime\prime}$ in the proof, the construction of $\mathcal{X}^{\prime\prime}_{\rm min}$ from $\mathcal{X}^{\prime}_{\rm min}$ is also canonical. If $n=2$, then this process is trivial i.e., $W=C_{\sigma(0)}(X)$ (as Fano cones).
Below, we use the framework of [Od24b] more to refine and generalize above Theorem 2.4 to Theorem 2.6, following the same circle of ideas. (For the readers who prefer the naive more classical Setup 1, one can skip or only skim through it.) In particular, the total space $\mathcal{X}^{\prime}_{\rm min}$ can be also taken “more canonically”, and we can treat much more general degeneration families over Novikov type rings. We prepare the following additional natural terminology, to compare with Setup 1 and partially review [Od24b].
Definition 2.5 (Restriction to one parameter families).
For a field $\mathbbm{k}$ and a submonoid $\Gamma_{\geq 0}\subset\mathbb{R}_{\geq 0}$, we consider N-ring $\mathbbm{k}[[\Gamma_{\geq 0}]]$ (see [Od24b, §2]). For a faithfully flat finite type morphism $\tilde{\pi}\colon\widetilde{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$, we call another faithfully flat finite type morphism $\mathcal{X}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$ is its restriction if there is a finitely generated submonoid $\overline{\Gamma}_{\geq 0}\subset\Gamma_{\geq 0}$ and a monoid surjective morphism $\varphi\colon\Gamma_{\geq 0}\twoheadrightarrow\mathbb{Z}_{\geq 0}$ such that
  • $\tilde{\pi}\colon\widetilde{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$ descends to $\overline{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\overline{\Gamma}_{\geq 0}]]$,
  • the base change of $\overline{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$ to $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$, through the $\mathbbm{k}$-algebra homomorphism $\mathbbm{k}[[\mathbb{Z}_{\geq 0}]]\to\mathbbm{k}[[\overline{\Gamma}_{\geq 0}]]$ induced by $\varphi$, is isomorphic to $\mathcal{X}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$.
We call the maps $\varphi$ or its associated $\mathbbm{k}$-algebra homomorphism $\mathbbm{k}[[\mathbb{Z}_{\geq 0}]]\to\mathbbm{k}[[\overline{\Gamma}_{\geq 0}]]$ a restriction map.
Theorem 2.6 (Algebro-geometric construction of (minimal) bubbles - ver2: with canonical base changes).
Following the setup of the above Definition 2.5, we also fix a section $\sigma\colon\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]\to\widetilde{\mathcal{X}}$ as setup 2 and further assume
  • $\mathbbm{k}$ has characteristic 00,
  • $\tilde{\pi}\colon\widetilde{\mathcal{X}}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$ is also affine,
  • $K_{\widetilde{\mathcal{X}}/\mathbbm{k}[[\Gamma_{\geq 0}]]}$ is $\mathbb{Q}$-Cartier and the special fiber $\widetilde{\mathcal{X}}_{0}$ is log terminal (we call this condition (being) log terminal morphism following the literature)
  • strictly degenerating in the sense that $\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}_{0}}(\sigma(0))<\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}_{\eta}}(\sigma(\eta)),$ where $\eta(\text{resp., }0)\in\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]$ is the generic point (resp., the closed point) and $\widetilde{\mathcal{X}}_{0}$ (resp., $\widetilde{\mathcal{X}}_{\eta}$) means its fiber.
  • $\tilde{\pi}$ has a restriction $\pi\colon\mathcal{X}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$ (which always holds for finite type $\tilde{\pi}$, and we fix the restriction map $\varphi$) in the sense of Definition 2.5.
Then, there is an iterative procedure of modifying $\tilde{\pi}$, as follows:
  1. (a)
    a finite (see Theorem 2.8) sequence of submonoids of $\mathbb{R}_{\geq 0}$ as
    $\Gamma_{\geq 0}$ $\subset\Gamma_{1,\geq 0}^{\prime}\subset\Gamma_{1,\geq 0}^{\prime\prime}$
    $\subset\Gamma_{2,\geq 0}^{\prime}\subset\Gamma_{2,\geq 0}^{\prime\prime}$
    $\subset\cdots\cdots\subset\mathbb{R}_{\geq 0}.$
  2. (b)
    vertical birational transforms of the base change of $\tilde{\pi}$ as faithfully flat affine log-terminal morphisms $\widetilde{\mathcal{X}}_{i}^{\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{i,\geq 0}^{\prime}]]$ and $\widetilde{\mathcal{X}}_{i}^{\prime\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{i,\geq 0}^{\prime\prime}]]$ for each $i=1,2,\cdots$,
  3. (c)
    their sections $\sigma_{i}^{\prime}\colon\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{i,\geq 0}^{\prime}]]\to\widetilde{\mathcal{X}}_{i}^{\prime}$,
  4. (d)
    their sections $\sigma_{i}^{\prime\prime}\colon\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{i,\geq 0}^{\prime\prime}]]\to\widetilde{\mathcal{X}}_{i}^{\prime\prime}$,
(the generic fibers with their sections are compatible with respect to the base change for (a)) such that
  • the set of pairs of restrictions of $\widetilde{\mathcal{X}}_{i}^{\prime}$, $\widetilde{\mathcal{X}}_{i}^{\prime\prime}$ (for same $i$), with respect to restriction maps which extend $\varphi$ compatibly, give the set of $(\mathcal{X}^{\prime},\mathcal{X}^{\prime\prime})$ in the previous Theorem 2.4 (in particular, $i=1$ case gives $(\mathcal{X}^{\prime}_{\rm min},\mathcal{X}^{\prime\prime}_{\rm min})$.
  • For each $i$,
    $\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}_{0}}(\sigma(0))$ $<\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}^{\prime}_{i,0}}(\sigma^{\prime}_{i}(0))$
    $=\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}^{\prime\prime}_{i,0}}(\sigma^{\prime\prime}_{i}(0))(=:c_{i})$
    $\leq\widehat{\operatorname{vol}}_{\widetilde{\mathcal{X}}_{\eta}}(\sigma(\eta)),$
    and $c_{i}$ are monotonely increasing for $i=1,2,\cdots$, where $\widetilde{\mathcal{X}}^{\prime}_{i,0}$ and $\widetilde{\mathcal{X}}^{\prime\prime}_{i,0}$ are the central fibers of $\widetilde{\mathcal{X}}^{\prime}_{i,0}$ and $\widetilde{\mathcal{X}}^{\prime}_{i,0}$ respectively.
We prove both Theorems 2.4, 2.6 in the subsection §2.2.3.
Remark 2.7 (Locality and logarithmic generalization).
From the following proof of Theorem 2.4, 2.6 combined with the uniqueness statements in [Od24b, 3.8 (2) (also (1)), 3.15 (2)], it follows that the germs of $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ around the images of $\sigma^{\prime}$ (resp., $\sigma^{\prime\prime}$) is determined only by the germ of the image of $\sigma\colon\Delta\to\mathcal{X}$. Generalization of the above results to the so-called log pair setup i.e., family of kawamata-log-terminal log pairs $\{(X_{t},D_{t})\}_{t\in\Delta}$ (cf., e.g., [BL21, §2.4] for review of the definition) and corresponding weak conical singular Kähler metrics, is also almost verbatim which we omit here to avoid technical complication.
For the sake of organization of our explanations, here we focus and finish the issue of finiteness of bubbles, for fixed $\pi$, as a part of the previous theorem 2.4. We use [XZ24] but we note that a differential geometric counterpart is proven before for the case when $\mathcal{X}_{t}(t\neq 0)$ are all smooth, compactifiable to metrized Kähler manifolds, in [Sun23, Corollary 1.12].
Theorem 2.8.
We fix a positive integer $n$. In the setup 1, 2, 3, we modify $\mathcal{X}$ by $\mathcal{X}^{\prime\prime}_{\rm min}$ as in Theorem 2.4. Then this modification process does not continue for infinitely many times for $\operatorname{dim}(X)=n$. Hence, we algebraically obtain a sequence of (weighted) algebraic bubbles $Z_{0},\cdots,Z_{2k}$ similarly to [Sun23, Cor 1.12].
In particular, as the terminal object of this repetition, possibly after a finite base change of $\Delta$, there is an affine faithfully flat klt family $\mathcal{Y}\to\Delta$ which coincides with $\mathcal{X}\to\Delta$ over $\Delta\setminus\{0\}$, $Y=Z_{2k}$, and the section $\sigma$ to a section $\tau$ from $\Delta$ so that
(6) $\widehat{\operatorname{vol}}_{\mathcal{Y}_{0}}(\tau(0))$ $=\widehat{\operatorname{vol}}_{\mathcal{Y}_{t}}(\tau(t))$
(7) $=\widehat{\operatorname{vol}}_{\mathcal{X}_{t}}(\sigma(t)).$
for an arbitrary $0<|t|\ll 1$.
Proof.
The modification process $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}_{\rm min}$ terminates by [XZ24]. We take $\mathcal{Y}$ as the end object of the repetition of the process $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}$, towards deeper bubbles. Then the last equality (6) holds as otherwise we can repeat the process by Theorem 2.4, and the equality (7) is obvious from identification of $\mathcal{X}$ and $\mathcal{Y}$ outside fibers over $0\in\Delta$. ∎
Note that the last statement of the above Theorem 2.8 is a semistable reduction type theorem corresponding to virtual compactness of “moduli” of singularities with fixed normalized volumes. In particular, the following corollary to above, of a weak version of the resolution of singularities type, is pointed out by M.Mauri.
Corollary 2.9 (Alteration of log terminal singularities).
Consider any affine pointed flat morphism $\pi\colon\mathcal{X}\to\Delta\ni 0$ of relative dimension $n$, with a section $\sigma\colon\Delta\to\mathcal{X}$ such that fibers $\pi^{-1}(s)=\mathcal{X}_{s}\ni\sigma(s)$ are all smooth points (resp., log terminal) for $s\neq 0$ (resp., $s=0$), over a field $\mathbbm{k}$ of characteristic 00, as in Setup 1.
Then there is a birational modification $\tilde{\pi}\colon\mathcal{Y}\twoheadrightarrow\Delta$ along $\mathcal{X}_{0}$, possibly after a finite base change of $\Delta$, such that $\sigma|_{\Delta\setminus 0}$ extends to whole $\Delta$ as $\sigma_{\mathcal{Y}}\colon\Delta\to\mathcal{Y}$, $\tilde{\pi}$ is smooth surjective around $\sigma_{\mathcal{Y}}(s)$ for any $s$ near 00.
If $n=2$, more generally, if $\pi$ is a log terminal morphism, one can construct $\mathcal{Y}$ with $\sigma_{\mathcal{Y}}$ which are étale locally trivial around $\sigma_{\mathcal{Y}}(s)$ around the neighborhood of 00.
Our main reason of writing this is on our different method than usual resolution of singularities, rather than the result itself.
Proof.
Suppose the last assertion is not satisfied for the original $\mathcal{X}$. Otherwise, the statements is obvious. Now we apply Theorem 2.4 and Theorem 2.8 to obtain the deepest bubbling family $\tilde{\pi}$. From (6) and (7) of Theorem 2.8, $\widehat{\rm vol}_{\mathcal{Y}_{0}}(\sigma_{\mathcal{Y}}(0))=n^{n}$. Then, [LX19, Theorem A.4] (see also implicit arguments in [HS17, SS17, LX18]) shows that $\sigma_{\mathcal{Y}}(0)\in\mathcal{Y}_{0}$ is a smooth point. Hence we complete the proof of the former claim for general $n$. The claim for $n=2$ follows from the same arguments, once we combine with the simpler proof of Theorem 3.6 of [Od24a] using the comparison of the orders of the local fundamental groups. ∎
Further, note that the above analogue to resolution of singularities $\tilde{X}$ is often taken canonically if $\mathcal{X}_{0}$ has quasi-regular minimizer of $\widehat{\rm vol}(-)$.
In our companion paper [Od24a], we restrict our focus to cones where the similar method applies to prove the “proper”-ness part of the following theorem.
Theorem 2.10 (K-moduli of Fano cone [Od24a]).
Consider the set (resp., all $\mathbb{Q}$-Gorenstein families) of the $\mathbbm{k}$-isomorphic classes of K-polystable (resp., K-semistable) $\mathbb{Q}$-Fano cone of $n$-dimension, with the fixed normalized volume $V$ of the vertices. Then, it forms a good proper (and separated) moduli algebraic space over $\mathbbm{k}$ and its stacky enhancement, generalizing the K-moduli of $\mathbb{Q}$-Fano varieties.
2.2.2. Statements on the comparison with differential geometry
As a link to differential geometric bubblings i.e., rescaled up pointed Gromov-Hausdorff limits, as studide in [Sun23, dBS24], we first prove the following, which confirms [Sun23, Conjecture 5.4 (i)] and also makes a progress on Conjecture 5.4 (ii) of op.cit (cf., also [dBS24, §4.1]).
Theorem 2.11 (Partial identification with analytic construction of [Sun23]).
We work under the Setup 1 (a), 2, 3, and we further assume that
  • $\pi$-ample $\mathcal{L}$ is put on $\overline{\mathcal{X}}$ so that $g_{t}$ is restriction of the unique Kähler-Einstein metric on $\mathcal{X}_{t}$ whose Kähler class is $2\pi c_{1}(\mathcal{L}_{t}:=\mathcal{L}|_{\mathcal{X}_{t}})\in H^{1,1}(\overline{\mathcal{X}}_{t},\mathbb{R})$.
  • $(\mathcal{X}_{0},\mathcal{L}_{0},g_{0})$ is the polarized limit space of $(\overline{\mathcal{X}}_{t},\mathcal{L}_{t},g_{t})$ for $t\to 0$ in the sense of [DS14] (and [DS17]),
to make [DS17, Sun23] works. Then, if the ambient local coordinates around $\sigma(\Delta)\subset\mathcal{X}$ are compatible with [Sun23, 4.5, 4.6],
  1. ($\mathrm{i}$)
    our algebraic minimal bubble $\mathcal{X}^{\prime}_{\rm min,0}$ can be described by the notations in the proof of [Sun23, Therorem 4.3] after that discussions, and is unique for each $\pi$.
  2. ($\mathrm{ii}$)
    our $\mathcal{X}^{\prime\prime}_{\rm min,0}$ satisfies the conditions (a), (b) of [Sun23, Conjecture 5.4].
The above compatibility of local ambient coordinates mean the coincidence of (constants times) the restriction of global coordinates and $\tilde{f}^{\alpha}_{j,k}$ in [Sun23, §4, after Lemma 4.5]. We expect somewhat weaker asymptotic version is enough but anyhow the above discussions naturally motivate the following question.
Question 2.12 (Coincidence with differential geometric bubbling).
Does our $\mathcal{X}^{\prime\prime}_{\rm min,0}$ in Theorem 2.4, applied with some $\mathcal{T}^{\prime\prime}$ and good ambient local coordinates, coincide with the (differential geometric) minimum bubble in the sense of [Sun23, Definition 1.9] (Definition 2.1), obtained in [Sun23, Theorem 1.10]?
Note that we refer to a priori non-canonicity of the constructions in Theorem 2.4, which we expect to parallel the choice of the family of Kähler metrics $\{g_{t}\}_{t}$ on $\mathcal{X}\to\Delta$. Below, we confirm the above question affirmatively in the following $1$ or $2$-dimensional classical cases considered in [dBS24]. The statements are for deeper bubbles as well, due to the iterative nature.
Theorem 2.13 (Confirmation - coincidence with differential geometric bubbling).
The above question has affirmative answers at least in the following three cases, which form examples of Setup 1. More precise statements are itemized as follows. In all the cases, the semistable bubbles and polystable bubbles coincides i.e., $\mathcal{X}^{\prime}=\mathcal{X}^{\prime\prime}$ for each pair.
  1. ($\mathrm{i}$)
    (“Log curve” case) If we consider $\mathcal{X}:=\mathbb{C}\times\Delta_{t}\to\Delta_{t}$ with flat (singular) Kähler metrics with conical singularities $p_{i}(t)\in\mathbb{C}$ for $i=0,\cdots,k$ of angle $2\pi(1-\beta_{i})$ with $10<\beta_{i}<$ 1, the sequence of bubblings constructed algebro-geometrically by Theorem 2.4 (cf., Remark 2.7) can be taken canonically, coincides with that of [dBS24, §2.2, Lemma2] and hence coincides with differential geometric bubbles i.e., rescaled pointed Gromov-Hausdorff limits (or, polarized limit space in the sense of [DS17], to be more precise).
  2. ($\mathrm{ii}$)
    (ALE case with $n=2$) Consider general smoothing affine family of ADE singularity $\mathcal{X}(\supset\Delta_{t})\to\Delta_{t}$ where $X=\mathcal{X}_{0}$ has only one singularity at $\sigma(0)=0$ which is of $A_{k}$ type. Suppose we have a continuous family of ALE hyperKähler (orbi-)metrics $g_{t}$ on $\mathcal{X}_{t}$ whose Kähler forms $(\omega_{I,t},\omega_{J,t},\omega_{K,t})$ ($I$ is the original complex structure for $X_{t}$s in Setup1) satisfy that
    • $\omega_{I,t}$ is $\partial\overline{\partial}$-exact
    • $\omega_{J,t}+\sqrt{-1}\omega_{K,t}$ is holomorphic on $\mathcal{X}\setminus X$.
    Then, the sequence of bubblings constructed algebro-geometrically by Theorem 2.4 can be taken canonically again, coincides with the construction of [dBS24, §3.3, Theorem 3] and hence coincides with differential geometric bubbles.
  3. ($\mathrm{iii}$)
    ($A_{1}$ case with any $n$) If $\mathcal{X}\to\Delta_{t}$ is open subset of proper holomorphic family over $\Delta$ which is the smoothing of higher dimensional $A_{1}$-singularity, the minimal bubbling constructed algebro-geometrically by Theorem 2.4 is $V(z_{1}^{2}+\cdots+z_{n}^{2}-1)\subset\mathbb{A}^{n+1}$ and coincides with the differential geometric minimal bubble with the Stenzel asymptotic conical metric ([EH79, Ste93]).
We proceed to the proofs of the above Theorems 2.4, 2.6, 2.11, and 2.13.
2.2.3. Details of the algebraic algorithmic construction
In this subsubsection, we prove Theorem 2.4 and 2.6.
proof of Theorem 2.4.
First we discuss the construction of $\mathcal{X}^{\prime}_{\rm min}$, its central fiber $\mathcal{X}^{\prime}_{{\rm min},0}$ and its degeneration to a K-semistable klt cone $W\curvearrowleft T$ (recall at §2.1). We consider the deformations of $W$ and here we a priori face the problem of infinite-dimensionality of the deformation space of affine varieties, especially those with non-isolated singularities (see [SJZ23, §6, 6.9]). Nevertheless, following the idea of negative weight deformation ([CH22], also cf., [Sun23, dBS24]), we can circumvent the problem.
First, we take a temporary embedding of $W$ to $\mathbb{A}^{l}$ via (non-constant) $T$-semiinvariants $z_{1},\cdots,z_{l}$ so that the vertex maps to the origin. Suppose its Reeb vector field $\xi$ which corresponds to the volume minimizer and has weights $w_{1},\cdots,w_{l}\in\mathbb{R}_{>0}$ on each coordinate of $\mathbb{A}_{\mathbbm{k}}^{l}$ i.e., ${\rm deg}_{\xi}(z_{i})=w_{i}$ (cf. [DS17, Od24a, Od24b]). We consider finite generators $f_{1},\cdots,f_{N}$ of the defining ideal which gives $W\subset\mathbb{A}_{\mathbbm{k}}^{l}$ and set $d_{i}:={\rm deg}_{\xi}(f_{i})$. Consider all affine deformations of $W$ of the $\xi$-negative type $V(\{f_{i}+h_{i}\}_{i=1,\cdots,N}\mid 0<{\rm deg}_{\xi}(h_{i})<{\rm deg}_{\xi}(f_{i}))$. Also compare more intrinsic but conditional related construction: Corollary 2.17. Then, by applying the flattening stratification and taking the universal hull [Kol08, 1, 2], we obtain a $T$-equivariant affine flat $\mathbb{Q}$-Gorenstein (locally stable) family $\pi_{\mathcal{Y}}\colon\mathcal{Y}\to{\rm Def}^{-}(W)\subset\mathbb{A}^{m}$ for $m\gg 0$. We consider its $T$-invariant locally closed subset
$Z^{+}:=\{b\in{\rm Def}^{-}(W)\mid\widehat{\rm vol}(0\in\pi_{\mathcal{Y}}^{-1}(b))=\widehat{\rm vol}(0\in W))\},$
by [BL21]. By replacing ${\rm Def}^{-}(W)$ by a $T$-invariant open affine neighborhood by Sumihiro’s theorem ([Sum74]), we can suppose $Z^{+}$ is its closed subset. Further, by [Che24], $\mathcal{O}_{\pi_{\mathcal{Y}}^{-1}(Z^{+})}$ has ideal sequences $\{I_{\lambda_{i}}\}_{\lambda_{i}\in S(W)\subset\mathbb{R}_{>0}}$ such that for any $b\in Z^{+}$$\{I_{\lambda_{i}}|_{\pi_{\mathcal{Y}}^{-1}(b)}\}_{\lambda_{i}\in S(W)\subset\mathbb{R}_{>0}}$ gives the degeneration to the K-semistable Fano cone à la [DS17]. Here, $S(W)$ is the holomorphic spectrum of $W$ also in the sense of [DS17]. Recall these at §2.1.
Now, we take a positive integer $m$ and shrink ${\rm Def}^{-}(W)(\ni[W])$ enough again to a $T$-invariant open affine neighborhood such that all $i=1,\cdots,m$, $I_{\lambda_{i}}/I_{\lambda_{i+1}}$ are all free $\mathcal{O}_{Z^{+}}$-modules of rank $r_{i}$, and have $T$-equivariant splittings of $I_{\lambda_{i}}\twoheadrightarrow I_{\lambda_{i}}/I_{\lambda_{i+1}}\simeq\mathcal{O}_{Z^{+}}^{r_{i}}$ which corresponds to $T$-semiinvariant $\{z_{i,j}\in I_{\lambda_{i}}\mid 1\leq i\leq m,1\leq j\leq r_{i}\}$ ([IS15, 2.5]). By the complete reductivity of $T$, one can also extend $z_{i,j}$ to $T$-semiinvariant regular functions on whole $\mathcal{Y}$, which vanishes at the 00-section. With large enough $m$ and shrunk enough ($T$-invariant) $[W]\in{\rm Def}^{-}(W)$, we can and do assume the above $z_{i,j}$ generates $\Gamma(\mathcal{O}_{\pi_{\mathcal{Y}}^{-1}(b)})$ as $\mathbbm{k}$-algebras for any $b\in{\rm Def}^{-}(W)$. From the construction, it follows that for any $b\in Z^{+}$, the natural $T$-action with the weights $w_{i}$ give the degeneration to K-semistable Fano cone. We here reset ł$\l:=\sum_{i=1}^{m}r_{i}$ and $\xi=(w_{1},\cdots,w_{1},\cdots,w_{m},\cdots,w_{m})$. Then, the above $z_{i,j}$ gives a large reembedding of $\mathcal{Y}\hookrightarrow\mathbb{A}_{\mathbbm{k}}^{l}\times{\rm Def}^{-}(W)$.
Next, we enlarge ${\rm Def}^{-}(W)$ a little just to include $\mathcal{X}_{t}$s. Suppose the reembedding $W\subset\mathbb{A}_{\mathbbm{k}}^{l}$ constructed above is defined by $c$ polynomial equations $g_{1},\cdots,g_{c}$. Then, take an arbitrary extended embedding of $\mathcal{X}$ to $\mathbb{A}_{\mathbbm{k}}^{l}\times_{\mathbbm{k}}\Delta$ and suppose that $\mathcal{X}_{t}$ for generic $t$ is defined by equations of $\xi$-degrees bounded above by some $C>0$. Now we consider the set of all possible polynomials $g_{1},\cdots,g_{c}$ on $\mathbb{A}_{\mathbbm{k}}^{l}$ of the degrees at most by $C$, without constant terms, and consider the corresponding universal family $\mathcal{U}\twoheadrightarrow\mathbb{A}_{\mathbbm{k}}^{p}$ for $p\gg 0$. There is a corresponding morphism $\iota\colon\Delta\to\mathbb{A}_{\mathbbm{k}}^{p}$ to this $\mathcal{X}\to\Delta$.
Then we set $M:={\rm Def}^{-}(W)\cup T\cdot\iota(\Delta)\subset\mathbb{A}_{\mathbbm{k}}^{p}$, possibly after shriking $\Delta$ if necessary. Recall that $\Delta$ in Setup 1 (b) is a pointed smooth affine $\mathbbm{k}$-curve so that $M$ underlies a structure of algebraic $\mathbbm{k}$-scheme. Here, this $T$-action on $\iota(\Delta)\subset\mathbb{A}_{\mathbbm{k}}^{p}$ is given via the re-embedding coordinates and is the same for the whole $T$-action on $M$.
We denote the base change of $\mathcal{U}\to\mathbb{A}_{k}^{p}$ to $M$ as $\mathcal{V}\to M$ and $\mathcal{M}:=[M/T]$. (For those who prefer rank $1$ setup, one could alternatively take $\mathcal{T}$ a (non-canonical) affine test configurations induced by an algebraic subtorus $T^{\prime}$ of $T$ with the rank $1$ (the symbol $\mathcal{T}^{\prime}$ is reserved for different family to appear later as it is more natural notation for it), as a subfamily of the whole family $\mathcal{X}_{\tau}\to U_{\tau}$ of [Od24b, Example 2.9]. For $r>1$ case, this $T^{\prime}$ is clearly non-unique as a rational approximation of the algebraic positive degeneration $X\rightsquigarrow W$ ([LX18, §3], [Od24b, §2]).)
From our construction, the resulting $\varphi\colon[\mathcal{V}/T]\to\mathcal{M}$ is a $\mathbb{Q}$-Gorenstein log terminal family over an Artin stack and all the fibers contain the origin 00.
Now it follows that $[Z^{+}/T]$ has a structure of canonical higher $\Theta$-strata ([Od24b]) on $\mathcal{M}$ near $[W]$ by the above construction, which is also locally liftable. This follows from the above construction of the $T$-acted base, using [Che24], together with the following fact: if you have a log terminal affine morphism with base points $f\colon\tilde{\mathcal{X}}(\supset\sigma(S)=\{0\}\times S)\to S$ and a uniform positive weight degeneration $\tilde{X}_{s}\ni\vec{0}\rightsquigarrow W_{s}\ni\vec{0}$ of the pointed klt affine variety $X_{s}\ni\vec{0}$ to its K-semistable cone $W_{s}$, i.e., if there is a flat family over $S\times_{\mathbbm{k}}U_{\tau}$ with a connected algebraic $\mathbbm{k}$-scheme $S$ and an affine toric variety $U_{\tau}$ which corresponds to a rational polyhedral cone $\tau$, the corresponding minimal normalized volume $\widehat{\rm vol}(f^{-1}(s)\ni 0)$ is constant i.e., not depending on $s\in S$. This follows from the comparison with the leading coefficient of the index character (cf., [MSY08, Li19, CS18] and also [BB73, 4.1], [Od24a, §2]). On the other hand, over a neighborhood of $[W]\in Z^{+}$, our new $T$-action gives a higher $\Theta$-strata by [Che24].
Hence, we can apply semistable reduction type theorem [Od24b, Theorem 3.8 (also cf., 1.1)] (which generalizes [Lan75, AHLH23]) and we obtain a Langton type modification of the original morphism $f\colon\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}\to[M/T]$ (resp., $f\colon\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]\to[M/T]$) which corresponds to $\mathcal{X}$. For Setup 1 (b) case, recalling the arguments in §3 of [Od24b] and its notations, we consider $\overline{f}$ and its restriction $f_{\rm new,\xi}\colon{\rm Spec}(R_{\rm new}(\xi^{\prime}))\to\mathcal{M}$ as in its Corollary 3.10. It corresponds to an affine faithfully flat family over $\mathbbm{k}[[M^{\prime}_{\xi^{\prime}_{\geq 0}}]]$. Since $R_{\rm new}(\xi^{\prime})=\mathbbm{k}[[M^{\prime}_{\xi^{\prime}_{\geq 0}}]]$, we can consider its arbitrary restriction in the sense of Lemma 3.5 of §3.1 of [Od24b] and denote it as $\mathcal{X}^{\prime}_{\rm min}$ (see the statements of Theorem 2.4). For the analytic variant Setup 1 (a) case, it is easy to see that by the above procedure and a restriction in an analogous sense to Lemma 3.5 of op.cit, one obtains $\mathcal{X}^{\prime}_{\rm min}\to\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$  [^10] Here, $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ denotes the discrete valuation ring which consists of the convergent power series of $t$. possibly after replacement of $t$ by its $d$-th root for some $d\in\mathbb{Z}_{>0}$. We are now going to confirm the required conditions of $\mathcal{X}^{\prime}_{\rm min}$ in our Theorem 2.4 and that $\mathcal{W}$ of ($\mathrm{iv}$) also naturally show up during the construction.
As noted in [Od24b], note that the family $\mathcal{X}^{\prime}_{\rm min}$ (as the total space) is non-canonical which corresponds to a choice of $T^{\prime}$ while the central fiber $\mathcal{X}^{\prime}_{\rm min,0}$ does not depend on it. Nevertheless, under certain conditions, they can be constructed canonically. For instance, suppose $W$ has only isolated singularity and have unobstructed deformation i.e., with smooth $T$-equivariant deformation space $[W]\in{\rm Def}(W)$ which is algebraized by [Art69, Pin74]. Then, by [KR82, BH85, LP17], $[W]\in{\rm Def}(W)$ is $T$-equivariantly isomorphic to $0\in\mathbb{A}_{\mathbbm{k}}^{d}$ with linear $T$-action. This gives a $T$-equivariant decomposition ${\rm Def}(W)=\oplus_{i\in I}V_{i}$ to $T$-eigenspaces $V_{i}$. We denote the $\xi$-weights on $V_{i}$ as $W_{i}$. Now we add an assumption that $\iota\colon\Delta\to{\rm Def}(W)$ which corresponds to $\mathcal{X}$ has non-vanishing differential at each each $V_{i}$: we call this condition weakly non-degenerating, comparing with stronger conditions in [BR15]. A subtle point is that because of a priori non-universality of ${\rm Def}(W)$, $\iota$ is not canonical and only its differential $\partial\iota_{t=0}$ is canonical by the versality. On the other hand, the construction of $\mathcal{X}^{\prime}_{\rm min}$ shows that they are determined by $\partial_{t}\iota|_{V_{i}}$ for those $V_{i}$ with $W_{i}=1$. Thus, the weak non-degeneracy assumption implies that the above construction of $\mathcal{X}^{\prime}_{\rm min,0}$ is canonical. This confirms ($\mathrm{vi}$).
The obtained morphism from $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$), or its finite extension $\mathbb{C}[[t^{\frac{1}{d}}]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t^{\frac{1}{d}}]]$) for some $d\in\mathbb{Z}_{>0}$, to $\mathcal{M}$ gives the desired family $\mathcal{X}^{\prime}_{\rm min}\to\Delta$ from the construction.
The K-semistability of $\mathcal{X}^{\prime}_{\rm min,0}$ in the sense of [Sun23, §5] follows from the construction as $W$ is a K-semistable log terminal cone ($\mathbb{Q}$-Fano cone) in the sense of [CS18, CS19]. Furthermore, from the positivity of the weights on the coordinates of $T^{\prime}$-action and the definition of elementary modification [AHLH23, §3], we conclude that the obtained test configuration $\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow W$ is a negative weight deformation in the sense of [CH22]111111just of rank $1$ i.e., with constant $\xi_{i}$ in their definition, [Od24b, §2]. Note that we do not know if any good choice of $T^{\prime}$ gives rise to a $\Theta$-strata in the original sense, and independence of $T^{\prime}$ is also a priori unclear (though it follows from [Od24b, 3.8 (2): canonicity claim]). These are why we need the higher rank setup and arguments in the generalized form [Od24b, Theorem 3.8] rather than the original in [AHLH23, §6].
Now we are ready to discuss algebraic construction of $\mathcal{X}^{\prime\prime}_{\rm min}$ from $\mathcal{X}^{\prime}_{\rm min}$, which follows a similar line of ideas as those for $\mathcal{X}^{\prime}_{\rm min}$, but also with some differences. We take a (non-canonical) algebraic subtorus $T^{\prime\prime}$ of $T$, with the rank $1$, which degenerates $W$ to the K-polystable degeneration as in [LWX21, Theorem 1.2]. Denote the corresponding base i.e., the partial compactification of $T^{\prime\prime}$ as $\overline{T}^{\prime\prime}$ which is isomorphic to $\mathbb{A}^{1}$. Recall that the construction in loc.cit is far from canonical. We still take one of them and denote the corresponding test configuration as $\mathcal{T}^{\prime\prime}\to\overline{T}^{\prime\prime}$. Henceforth we can and do assume $\mathcal{T}^{\prime\prime}$ is also achieved inside $M$, by enlarging $M$ inside the original base $\mathbb{A}_{k}^{p}$ if necessary, and $\mathbb{C}^{l}$ (resp., $\mathbb{A}_{k}^{l}$) by retaking larger $l$, $\mathbb{C}^{l}$ (resp., $\mathbb{A}_{k}^{l}$) and the embedding $\mathcal{X}\hookrightarrow\mathbb{C}^{l}\times\Delta$ (resp., $\mathcal{X}\hookrightarrow\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$) is also realized in $M$ if necessary. Now we consider a couple of morphisms $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}\to M$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]\to M$) corresponding to $\mathcal{T}^{\prime}$ of degeneration of $X^{\prime}\rightsquigarrow W$, and corresponding to $\mathcal{T}^{\prime\prime}$ whose degeneration is of the form $W\rightsquigarrow C$.
We denote the base, with its associated moduli map, of the negative weight degeneration $\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow W$ by $\Delta^{\prime}\to M$. Then, consider the natural rational map from $m\colon\Delta^{\prime}\times\overline{T}^{\prime\prime}\to M$. If it extends to a morphism, one can take the image of $c\times\{0\}(\in\Delta^{\prime}\times\overline{T}^{\prime\prime})$ for $c\neq 0$ as $\mathcal{X}^{\prime\prime}_{\rm min,0}$. Otherwise, we apply [AHLH23, Theorem 6.3] (or [Od24b, Theorem 3.8]) to them, with respect to the simple $\Theta$-strata $[\overline{m(\Delta^{\prime}\times T^{\prime\prime})}/T^{\prime\prime}]$ supported at $m(0\in\Delta)$. Here, $\overline{m(\Delta^{\prime}\times T^{\prime\prime})}$ means the closure of the image $\overline{m(\Delta^{\prime}\times T^{\prime\prime})}$ of $m$. Then, it follows that possibly after passing to a finite ramified covering of the base $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$), we obtain a isotrivial family which degenerates $\mathcal{X}^{\prime}_{\rm min,0}$ to $\mathcal{X}^{\prime\prime}_{\rm min,0}\subset\mathbb{A}_{\mathbbm{k}}^{l}$. In this case, we can avoid the full use of [AHLH23, Od15b] by a more standard direct and canonical arguments using a toric surface as follows.
Note that $T^{\prime}$ and $T^{\prime\prime}$ are commutative so that they are both algebraic subgroups of an algebraic tori $T_{2}$ of rank $2$, which acts equivariantly on the family $\mathcal{U}\to\mathbb{A}_{\mathbbm{k}}^{m}$ and its base change $\mathcal{V}\to M$. Since the obtained elementary modification maps $(0,0)\in[\mathbb{A}^{2}/\mathbb{G}_{m}]={\rm ST}_{k[[t]]}$ (cf., [HL14, AHLH23]) to $[C\in\mathcal{M}]$ with the inertia group $T$, the elementary modification is represented by a $T_{2}$-equivariant morphism $f\colon\mathbb{A}_{\mathbbm{k}}^{2}\to M$. The normalization of its image in $M$ is a toric surface $T_{N\simeq\mathbb{Z}^{2}}\Sigma$ where the fan contains a ray $\mathbb{R}_{>0}(1,0)$ (resp., $\mathbb{R}_{>0}(0,1)$) which corresponds to $\mathcal{T}^{\prime}$ (resp., $\mathcal{T}^{\prime\prime}$) and its support ${\rm Supp}(\Sigma)=\mathbb{R}_{\geq 0}^{2}$ from the construction. Take the ray $\mathbb{R}_{>0}(s_{1},s_{2})(s_{i}\in\mathbb{Z}_{>0})$ which is adjacent to $\mathbb{R}_{>0}(1,0)$ and consider the one parameter subgroup $\lambda\colon\mathbb{G}_{m}\to T_{2}$. Then it gives a test configuration $\mathcal{X}^{\prime}_{\rm min,0}\rightsquigarrow\mathcal{X}^{\prime\prime}_{\rm min,0}$ for some $X^{\prime\prime}\in S$ such that the inverse action of $\mathcal{T}^{\prime\prime}$ degenerates $\mathcal{X}^{\prime\prime}_{\rm min,0}$ to $C$ as desired.
Now we glue a family $\mathcal{X}^{\prime}_{\rm min,0}$ and $\pi_{X^{\prime}}$, by the same technique as [LWX21, §3], [Od22a, §3], to obtain the desired $\mathcal{X}^{\prime\prime}_{\rm min,0}\to\Delta$, as a vertical birational modification of $\mathcal{X}^{\prime}_{\rm min,0}\to\Delta$, which satisfies the conditions listed in the statements of our theorem. The only part which does not follow immediately, is the equality assertion of ($\mathrm{iii}$). It holds because of the condition satisfied by the “minimally” constructed $\mathcal{X}^{\prime}_{\rm min,0}\ni\sigma^{\prime}(0)$ i.e., that it achieves the possible minimum normalized volume, combined with the semicontinuity [BL21, Theorem 1], [Xu20, Theorem 1.3] of $\widehat{\rm vol}$ applied to $\pi_{X^{\prime}}$. Here we complete the construction of $\mathcal{X}^{\prime\prime}_{\rm min,0}$ from $\mathcal{X}^{\prime}_{\rm min,0}$. If $n=2$, since any log terminal singularities are quotient singularities so that K-semistable Fano cones are automatically K-polystable, we have $\mathcal{X}^{\prime\prime}_{\rm min,0}=\mathcal{X}^{\prime}_{\rm min,0}$ from our construction.
Finally, to obtain more bubbles $(\mathcal{X}^{\prime},\mathcal{X}^{\prime\prime})$, we replace the given $\mathcal{X}_{\rm min}$ by $\mathcal{X}^{\prime\prime}_{\rm min}$ and repeat the whole process above. We show this repetition terminates after finite times in Theorem 2.8. We complete the proof. ∎
Remark 2.14.
From the above proof, it easily follows that the analytic/formal germs of $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ around the images of $\sigma^{\prime}$ (resp., $\sigma^{\prime\prime}$) is determined only by the analytic germ of the image of $\sigma\colon\Delta\to\mathcal{X}$. Generalization of the above results to the so-called log pair setup i.e., family of kawamata-log-terminal log pairs $\{(X_{t},D_{t})\}_{t\in\Delta}$ (cf., e.g., [BL21, §2.4] for review of the definition) is also almost verbatim which we omit here to avoid technical complication.
Remark 2.15.
The above replacement $\mathcal{X}$ by $\mathcal{X}^{\prime}$ or $\mathcal{X}^{\prime\prime}$ or the existence of bubbles can first look in a sharp contrast to what happens in analogous global setup i.e., the separatedness of the moduli of log terminal projective Calabi-Yau varieties ([Od12, 4.2 (i), 4.3],  [^12] However, although 4.3 of op.cit is fine, there is an obvious typo in its 4.2(i), mixing up notation with canonically polarized case: in the left hand side $-K_{\mathcal{X}/\mathcal{C}}$ should be read $\mathcal{L}$ and we assume K-stability (log terminality) of the fibers of $\pi$. [Bou14, 2.1]).
proof of Theorem2.6.
Our proof follows that of Theorem 2.4 closely, which essentially proves the case when $R=\mathbbm{k}[[t]]$ or $\mathbb{C}[[t]]^{\rm conv}$. For $\pi^{-1}(0)=\tilde{\mathcal{X}}_{0}$, we apply the same arguments as the beginning of the proof of Theorem 2.8. In particular, we take the same log terminal family $\mathcal{U}\subset\mathbb{A}_{\mathbbm{k}}^{l}\times\mathbb{A}_{\mathbbm{k}}^{m}$, which restricts to $\mathcal{V}\to M$ and consider the map $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{\geq 0}]]\to M\to\mathcal{M}$ which corresponds to $\tilde{\pi}$. As in the previous proof of Theorem 2.4, take the algebraic torus $T$ for $W$, the semistable tangent cone of $X$ and a rational polyhedral cone $\tau$ in the cocharacter lattice $N:={\rm Hom}(\mathbb{G}_{m},T)$ with $\xi\in\tau$ such that the base change of $\mathcal{V}$ over $U_{\tau}$ gives a flat isotrivial degeneration (over possibly higher dimensional base) with the central fiber $W$. If there is a uniform family of such isotrivial degenerations $X\rightsquigarrow W$ over a connected base, then from the same reason as in the proof of Theorem 2.4, i.e., by the use of index character ([MSY08, CS18, CS19]), the normalized volume are constant. Similarly as before, combined with the uniqueness of the $\widehat{\rm vol}$-minimizing valuation, this gives the canonical structure of higher $\Theta_{\tau}$-strata on $M_{c}$ for $c:=\widehat{\rm vol}(X,0)$. Then, we again apply the higher $\Theta$-stable reduction theorem [Od24b, §3.2, Theorem 3.8] to construct $\Gamma_{1,\geq 0}^{\prime}$, $\widetilde{\mathcal{X}}_{1}^{\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{1,\geq 0}^{\prime}]]$ and its section $\sigma_{1}^{\prime}$. The rest of the constructions are almost the same, so we omit the essential repetition. ∎
2.2.4. Comparison with the differential geometric construction
Now, we discuss comparison with the differential geometric construction in [DS14, Sun23] and prove Theorem 2.11 and Theorem 2.13.
proof of Theorem2.11.
We prove the item ($\mathrm{i}$) of Theorem 2.11. In this proof, we closely follow the arguments and notations of [Sun23] (cf., also [DS17]) by fitting it to our notation in Theorem 2.11. In the notation of loc.cit, suppose that his $X_{j}(j=1,2,\cdots)$ are taken as fibers of our $\pi$ i.e., $X_{j}=\overline{X}_{s_{j}}$ for some sequence $s_{j}\in\Delta$ with $s_{j}\to 0$ for $j=1,2,\cdots$ and $X_{\infty}=\pi^{-1}(0)=X$. Our $\xi$ corresponds to the real $N\times N$ diagonal matrix $\Lambda={\rm diag}(2^{\zeta_{1}},\cdots,2^{\zeta_{N}})$ with $\zeta_{i}\in\mathbb{R}_{>0}$s acting on the ambient space $\mathbb{C}^{N}$ in his notation. We crucially use the local holomorphic functions $G^{i},S^{i}_{j},L^{i}_{j}$ (and $H^{i}$, $H^{i}_{j}$ with higher degrees with respect to $\xi$) on the common ambient space $\mathbb{C}^{N}$ of the germs of $p_{j}\in X_{j}$s ($i$ was denoted as $b$ in loc.cit), $p_{\infty}\in X_{\infty}$ and the possible bubbles $Z$, the bounds of rescaling orders $\underline{k}$ and $\overline{k_{j}}$ as prepared in loc.cit. Then, we replace the sequence $\{j\}$ by its subsequence so that there is a sequence of real numbers $b_{j}(j=1,2,\cdots)$ with $b_{j}\to\infty$ and obtain
(8) $W_{1}:$ $=\lim_{j\to\infty}\Lambda^{b_{j}}\cdot X_{j}$
as a closed subscheme of $\mathbb{C}^{N}$. Let us write $\operatorname{deg}_{\xi}$ for the degree with respect to $\xi$ as in [Sun23, §4]. Here, each $L^{i}_{\infty}$ is obtained as a rescaled limit of $L^{i}_{j}$ for $j\to\infty$. Indeed, it can be achieved as follows. For each $i$, consider the set $L(i)$ of monomials $\chi^{\vec{m}}$ on the ambient space $\mathbb{C}^{N}$ whose $\xi$-degree is less than that of $G^{i}$. Its order is finite as $\zeta_{1},\cdots,\zeta_{N}$ are all positive. We denote the coefficient of $\chi^{\vec{m}}$ in $L^{i}_{j}$ as $c^{i}_{j}(\vec{m})$, then of course we have $c^{i}_{j}(\vec{m})\to 0\hskip 11.38109pt(j\to\infty)$ as $L^{i}_{j}\to 0\hskip 11.38109pt(j\to\infty)$. Now, for each fixed $j$, we consider
(9) $b_{j}:=\max_{i,\vec{m}\in L(i)}\dfrac{\log(|c^{i}_{j}(\vec{m})|)}{\log(2)\cdot(\operatorname{deg}_{\xi}(G^{i})-\operatorname{deg}_{\xi}(\chi^{\vec{m}}))}$
which exists by the finiteness of $\{i\}$ (the Noetherian property) and each $L(i)$. Hence, by passing to a subsequence of $j$, we can and do assume that the above maximum is attained by the same $i$ and $m$. Further, by the compactness of subset $\prod_{i}\{t\in\mathbb{C}\mid|t|\leq 1\}^{L(i)}$, as the set of coefficients of $2^{-\operatorname{deg}_{\xi}(G^{i})}(\Lambda^{b_{j}}\cdot L^{i}_{j})$, if we pass to a further subsequence, we can and do assume that for any $i$, $2^{-\operatorname{deg}_{\xi}(G^{i})}(\Lambda^{b_{j}}\cdot(G^{i}+L^{i}_{j}))$ converges when $j\to\infty$. We denote its limit as $G^{i}+c(i)L^{i}_{\infty}$ with some $c(i)\in\mathbb{C}$ for each $i$ and $\xi$-homogeneous polynomials $L^{i}_{\infty}$ with $\operatorname{deg}_{\xi}(L^{i}_{\infty})<\operatorname{deg}_{\xi}(G^{i}).$ Since $\Lambda^{\overline{k_{j}}}\cdot X_{s_{j}}\to W(j\to\infty)$ with the notation of [Sun23], for each i and $\vec{m}$, we have $2^{\overline{k_{j}}}|c^{i}_{j}(\vec{m})|\to 0$ for $j\to\infty$. On the other hand, for some $i$ and $\vec{m}$, we have $2^{b_{j}}|c^{i}_{j}(\vec{m})|\to 1$ for $j\to\infty$ by the construction. Therefore, we have $\lim_{j\to\infty}(b_{j}/\overline{k_{j}})=0$. Further, from the construction of $W_{1}$, it is clear that $(\Lambda^{-1})^{\mathbb{R}_{>0}}\cdot W_{1}$ is enhanced to be a test configuration degenerating to $W$. All these confirm the conditions for the uniqueness theorem [Od24b, Theorem 3.12 (ii)] so that we can apply it to conclude that $W_{1}$ is $\mathcal{X}^{\prime}_{{\rm min},0}$ in Theorem 2.4. Note that $\mathcal{X}^{\prime}_{{\rm min}}$ is the family level enhancement, which can be taken even canonical over a convergent N-type spectrum by [Od24b, Theorem 3.12]. Item (ii) is shown in Theorem 2.4 (but repeated for the comparison with [Sun23, Conjecture 5.4]). ∎
Now we prove Theorem 2.13. The main point is confirmation of the compatibility of our construction in the proof of Theorem 2.4 with the case by case analysis in [dBS24], in each case.
proof of Theorem 2.13.
In all the three cases of Theorem 2.13, $X\ni\sigma(0)$ is locally an isolated singularity with local $\mathbb{C}^{*}$-action (with conical singularities) and hence in particular we can consider its finite dimensional smooth (unobstructed) deformation space ${\rm Def}(X\ni\sigma(0))$ of the singularity.
First, we consider Case ($\mathrm{i}$). In this case, there is a finite covering (the Weyl type covering) $\mathcal{D}:=\mathbb{C}^{k+1}\to{\rm Def}(X\ni\sigma(0))=\mathbb{C}^{k+1}/W\simeq\mathbb{C}^{k+1}$, where $W\subset S_{k+1}$ is the subgroup of the symmetric group of degree $k+1$, as the symmetry of the tuple $(\beta_{0},\cdots,\beta_{k})$, which acts by the permutation. The quotient is a Galois covering with the Galois group $G$. We consider the morphism $\Delta_{t}\to{\rm Def}(X\ni\sigma(0))$ which is associated to $\mathcal{X}(\supset\Delta_{t})\twoheadrightarrow\Delta_{t}$ and its lift $\Delta_{t}\to\mathcal{D}$ possibly after replacement of $\Delta_{t}$ by its finite covering (a component of the base change $\mathcal{D}\times_{{\rm Def}(X\ni\sigma(0))}\Delta_{t}$). We still denote the finite covering as $\Delta_{t}$ for simplicity, and denote as $\varphi\colon\Delta_{t}\to\mathcal{D}=\mathbb{C}^{k}$, $t\mapsto(s_{0}(t),\cdots,s_{k}(t))$ where each $s_{i}(t)$ is holomorphic with $s_{i}(0)=0$. In particular, the vanishing order of $s_{i}(t)$ at $t=0$ is positive which we denote as $d_{i}$. We set $d:=\min_{i}\{d_{i}\}$.
Note that the induced action of $\mathbb{C}^{*}$ as a subgroup of the automorphism group of $X\ni\sigma(0)$ on $\mathcal{D}$ is with the weights $(-1,\cdots,-1)$. Thus, from the proof of Theorem (2.4), $\mathcal{X}^{\prime}_{\rm min}$ naturally corresponds to
(10) $t\mapsto\varphi^{\prime}_{\rm min}(t):=(t^{-d}s_{0}(t),\cdots,t^{-d}s_{k}(t))\in\mathbb{C}^{k+1}=\mathcal{D}.$
Indeed, for $\epsilon>0$, we have that $\lim_{t\to 0}=[X\ni\sigma(0)=0]$, while $\varphi^{\prime}_{\rm min}(0)\neq$. On the other hand, it is easy to see that the flat (conical singular) Kähler metric on $\mathbb{C}$ is uniquely described as
$\prod_{0\leq i\leq k}|z-s_{i}(t)|^{\beta_{i}-1}|dz|$
(compare [dBS24, Lemma 1]), it follows that the differential geometric bubbling is obtained as the flat singular $\mathbb{C}$ given by $\varphi^{\prime}_{\rm min}(0)$. By replacing $\varphi$ by $\varphi^{\prime}_{\rm min}$ and iterative the procedure, we complete the proof of ($\mathrm{i}$) of Theorem 2.13.
The proof for the case ($\mathrm{ii}$) is more or less parallel to the case ($\mathrm{i}$) as above. We consider the well-known miniversal deformation space of $A_{k}$-singularity by Brieskorn-Grothendieck-Kas-Schlessinger-Slodowy. Recall it is also reconstructed as Kronheimer’s construction of a family of hyperKähler quotients [Kro89a], which we use. The deformation space can be taken as $\mathbb{C}^{k}$ and there is a natural Weyl cover $\mathcal{D}$ which kills the monodromy.
In our setup, due to the existence of ALE hyperKähler metrics $g_{t}$ with the $\partial\overline{\partial}$-exact Kähler form $\omega_{I,t}$, $\mathcal{X}\to\Delta_{t}$ can be written as (after base change to the Weyl cover i.e., take a component of $\Delta_{t}\times_{{\rm Def}(X\ni\sigma(0))}\mathcal{D}$)
$X_{t}=V(xy-\prod_{i=0}^{k}(z-s_{i}(t))\subset\mathbb{A}_{\mathbb{C}}^{3},$
where $s_{i}(t)$ is holomorphic. Indeed, each $X_{t}$ can be compared with the Gibbons-Hawking ansatz for appropriate $(0,s_{i}(t))\in\mathbb{R}\times\mathbb{C}\simeq\mathbb{R}^{3}(i=0,\cdots,k)$ with the harmonic potential
$\sum_{0\leq i\leq k}\frac{1}{|z-s_{i}(t)|}$
and the corresponding connection form as an essentially unique solution to the Bogomolony-type monopole equation, as explained in [LB91, p.302-303], [dBS24, §3.2, §3.3] and it follows from Kronheimer’s Torelli-type theorem [Kro89b] that they are isomorphic as hyperKähler varieties. It is obvious that $s_{i}(t)$ should be continuous but is even holomorphic by the second bulleted assumption on the holomorphicity of $\omega_{J,t}+\sqrt{-1}\omega_{K,t}$ on $\mathcal{X}\setminus X$. Now we can apply the discussion very similar to the proof of case ($\mathrm{i}$) i.e., consider the vanishing order of $s_{i}(t)$ at $t=0$ which is positive and denote by $d_{i}$ and set $d:=\min_{i}\{d_{i}\}$. Then we consider the map to the Weyl cover $\varphi^{\prime}_{\rm min}\colon\mathcal{D}(=\mathbb{C}^{k+1})\to{\rm Def}(A_{k})$ exactly as (10). Since the induced action of $\mathbb{C}^{*}$ as a subgroup of the automorphism group of $A_{k}$-singularity $X\ni\sigma(0)$ on $\mathcal{D}$ is again of the weights $(-1,\cdots,-1)$, we see that our algorithm in Theorem (2.4) gives $\mathcal{X}^{\prime}_{\rm min}$ which corresponds to $\varphi^{\prime}_{\rm min}$. Therefore, by comparing with [dBS24, §3.3, Theorem 3] and iterate the construction, we complete the proof.
The case ($\mathrm{iii}$) is easy as the deformation space ${\rm Def}(X=A_{1}\ni 0)$ is $1$-dimensional and smooth, which is realized as the $\mathbb{C}^{*}$-equivariat degeneration to the cone over quadric, which also serves as an example of (universal) negative weight deformation of Corollary 2.17 (a) in the next subsection. Each general fiber comes with the Stenzel metric [Ste93] for the complex quadric and it is easy to see that the construction in Theorem 2.4 gives the each general fiber. Thus, by [dBS24, §4.2 Proposition 1], we complete the proof. ∎
2.2.5. Negative weight deformation as scaling down
We discuss the notion of negative weight deformation in the sense of [CH22] and equivalently rephrase more algebraically in terms of the generalized test configurations of [Od24b]. In op.cit, we also called it scale down deformation. As the name suggests, we aim to provide algebro-geometric formulation for the opposite analogue of Theorem 2.3. This finishes the proof of the nontrivial direction of [Od24b], and answers the Open Question 1.4 (2) of [CH22]. Note that, in the context of the previous section, the degenerations $\mathcal{X}^{\prime}_{0}\rightsquigarrow W$ and $\mathcal{X}^{\prime\prime}_{0}\rightsquigarrow C$ are both negative weight deformations by the construction.
Theorem 2.16 (Negative weight deformation).
  1. ($\mathrm{i}$)
    ([Od24b, Lemma 2.19]) Take an arbitrary algebraic negative weight deformation (cf., Definition 2.6 of [Od24b]) of a Fano cone $T\curvearrowright(Y,\xi)$ as $p\colon\mathcal{Y}\twoheadrightarrow U_{\tau}$, with $\xi\in\tau$. Then, the general fiber of $p$ is a negative $\xi$-weight deformation in the original sense of Definition 1.7 of [CH22].
  2. ($\mathrm{ii}$)
    Suppose that for a Fano cone $(T\curvearrowright Y,\xi)$, $Y$ has only an isolated singularity. Then, conversely to the above ($\mathrm{i}$), any negative $\xi$-weight deformation $X$ of normal affine variety of $Y$ with $\xi$ in the sense of [CH22] arises in this manner.
  3. ($\mathrm{iii}$)
[…]
In general, the direction of ($\mathrm{ii}$) and ($\mathrm{iii}$) has a difficulty to treat a priori infinite dimensional deformation space (cf., [SRZ23, §6]). The latter ($\mathrm{iii}$) is modelled on and include the examples of [Li19, CR17, HN, Sze19, Sze20, Chi22].
Proof.
The above item ($\mathrm{i}$) is proved in [Od24b], so it remains to prove ($\mathrm{ii}$) and ($\mathrm{iii}$).
We start with the case ($\mathrm{ii}$). We first confirm the existence of $T$-equivariant semi-universal deformation of $T\curvearrowright Y$ after Pinkham. The tangent space $\mathbb{T}^{1}$ of the deformation of $Y$ naturally admits $T$-action by [Pin74, Proposition 2.2] and the arguments (2.5) (2.6) of loc.cit. The existence of the equivariant family follows from the arguments in the proof of Schlessinger [Sch68, Theorem 2.11], as in [Pin74, (2.9), (2.10)]. Indeed, the inductive construction of $J_{q}$ in (2.9) of loc.cit (or [Sch68, 2.11]) extends without any change. We similarly take $F^{0}$ as a column vector of $T$-homogeneous equations $f_{i}(i=1,\cdots,N)$ of some $T$-equivariant embedding $Y\hookrightarrow\mathbb{A}^{N}$, and construct the lift $F^{q}$ whose entries are all $T$-equivariant by projection to $T$-eigenspace at the inductive step for each $q=0,1,2,\cdots$. Thus, the upshot of the necessary modification is simply to replace the degree (with respect to the $\mathbb{G}_{m}$-action) to the weight vector with respect to the $T$-action. Summarizing up, we obtained a $T$-equivariant semiuniversal deformation $\mathcal{U}\to{\rm Def}(Y)$ as a formal family. It can be retaken as an algebraic family by Artin’s theorem [Art69], which we denote by the same symbol.
From the construction, note that ${\rm Def}(Y)$ comes with an $T$-equivariant étale morphism $u\colon{\rm Def}(Y)\to\mathbb{T}^{1}$, where the latter means the affine space scheme for $\mathbb{T}^{1}$.
We decompose $\mathbb{T}^{1}$ to $T$-eigensubspaces as $\mathbb{T}^{1}_{\alpha}$ for each character $\alpha\in M:={\rm Hom}(T,\mathbb{G}_{m})$. Then we extract $\xi$-negative part as
$\mathbb{T}^{1}_{\xi<0}$ $:=\oplus_{\alpha\text{ s.t. }\langle\xi,\alpha\rangle<0}\mathbb{T}^{1}_{\alpha},$
${\rm Def}(Y)_{\xi<0}$ $:=\mathbb{T}^{1}_{\xi<0}\cap{\rm Def}(Y),$
$\mathcal{U}_{\xi_{<}0}$ $:=u^{*}({\rm Def}(Y)_{\xi<0}),$
and consider the flat affine deformation $u|_{\xi<0}\colon\mathcal{U}_{\xi_{<}0}\to{\rm Def}(Y)_{\xi<0}$ as the restriction of $u$. Because of the $T$-equivariance of $u$ and the conditions $\langle\xi,\alpha\rangle<0$, ${\rm Def}(Y)_{\xi<0}\to\mathbb{T}^{1}$ is injective, hence an immersion.
Now, if we consider the negative $\xi$-deformation of $Y$ in the sense of [CH22, Definition 1.7] and take the realizing sequence as $(X_{i},\xi_{i})_{i=1,2,\cdots}$. From the conditions (1) and (2) of [CH22, Definition 1.7] (also cf., its (4)), it follows that $[X_{i}]$ are in the same $T(\mathbbm{k})$-orbit of ${\rm Def}_{\xi<0}(\mathbbm{k})$. Hence if we take its normalization of the closure of the $T(\mathbbm{k})$-orbit, we obtain an affine toric variety which we denote as $U_{\tau}$ with $\tau\subset N_{\mathbb{R}}$. Thus we obtain the desired assertion of ($\mathrm{ii}$).
The proof of ($\mathrm{iii}$) is similar, after proving the finite dimensionality of subspace of versal deformation space in concern. Suppose that $X$ is a $T$-equivariant weighted hypersurface cone i.e., there is a polynomial $f(z_{0},\cdots,z_{n})\in\mathbbm{k}[z_{0},\cdots,z_{n}]$ which is in particular $\xi=(w_{0},\cdots,w_{n})$homogeneous i.e., every monomial terms of $f$ has the same degree $d$ with respect to $\xi$.
If $y$ is an isolated singularity, as in ($\mathrm{ii}$), there are only finitely many monomials $g_{1},\cdots,g_{s}$ of $z_{i}$s whose degree with respect to $\xi$ is less than $d$. In this case, the semi-universal negative $\xi$-deformation is written as
$({\rm Def}_{x}(Y))_{\xi<0}$ $=\mathbb{A}^{s}_{t_{1},\cdots,t_{s}},$
$\mathcal{U}_{\xi<0}$ $=V(f+\sum_{i=1}^{s}t_{i}g_{i}).$
If $Y$ have a $1$-dimensional singularity, the singular locus is a line i.e., $1$-dimensional linear subspace of $\mathbb{A}^{n+1}_{z_{0},z_{1},\cdots,z_{n}}$ which we can assume to be of the form $(z_{1}=\cdots=z_{n}=0)$ after a $T$-equivariant linear transformation of the coordinates. Then, we can assume $f$ is a polynomial of $z_{1},\cdots,z_{n}$. Similarly as above, we take a homogeneous basis of the Jacobian ring ${\rm Jac}(f)=\mathbbm{k}[z_{1},\cdots,z_{n}]/\langle\{\frac{\partial f}{\partial z_{i}}\}\rangle$ and denote the lift as $g_{i}$s. Similarly as above, in this case, the semi-universal negative $\xi$-deformation achieved on $\mathbb{A}^{n}$ is written as
$({\rm Def}_{x}(Y))_{\xi<0,\mathbb{A}^{n}}$ $=\mathbb{A}^{s}_{t_{1},\cdots,t_{s}},$
$\mathcal{U}_{\xi<0,\mathbb{A}^{n}}$ $=V(z_{0}+f+\sum_{i=1}^{s}t_{i}g_{i})$
$\simeq({\rm Def}_{x}(Y))_{\xi<0,\mathbb{A}^{n}}\times\mathbb{A}^{n}.$
This is the semi-universal object of those negative deformations which are affine spaces $\mathbb{A}^{n}$ as varieties. The rest of the proof is the same as ($\mathrm{ii}$). ∎
Note that in the proof of above Theorem 2.16 ($\mathrm{ii}$) and ($\mathrm{iii}$), the following is proved, which we extract for later convenience.
Corollary 2.17 (Semi-universal negative weight deformation).
  1. (a)
    For each Fano cone $(T\curvearrowright Y,\xi\in N_{\mathbb{R}})$ with the only isolated singularity $y$, there is a semi-universal negative weight deformation $u_{\xi<0}\colon\mathcal{U}_{\xi<0}\twoheadrightarrow({\rm Def}_{x}(Y))_{\xi<0}\subset{\rm Def}_{x}(Y)$ over a finite type base with a good $T$-action $({\rm Def}_{x}(Y))_{\xi<0}$ (in the sense of [LS13]) such that any (algebraic) negative $\xi$-weight deformation can be obtained by pullbacks of $\pi_{\xi<0}$.
  2. (b)
    For each Fano cone $(T\curvearrowright Y,\xi\in N_{\mathbb{R}})$ with the vertex $y$, we suppose $Y$ is an affine (algebraic) hypersurface with $1$-dimensional singular locus, which is smooth (line).
    Then if we consider all negative deformations $X$ of $(Y,\xi)$ which are isomorphic to $\mathbb{A}^{n}$, there is a semi-universal deformation of them $\mathcal{U}_{\xi<0,\mathbb{A}^{n}}\to({\rm Def}_{x}(Y))_{\xi<0,\mathbb{A}^{n}}$.
After the construction of the semi-universal negative deformation as above, it is natural to consider the corresponding moduli theory (cf., also [SRZ23, §6]). Here are some questions, which we hope to explore more in near future.
Definition 2.18.
For the above Kuranishi space ${\rm Def}_{\xi<0}(Y)$ of Corollary 2.17 (a) (resp., ${\rm Def}_{\xi<0,\mathbb{A}^{n}}(Y)$ of Corollary 2.17 (b)) for the negative $\xi$-weight deformation in Corollary 2.17, note that there is a canonical $G:={\rm Aut}_{T}(Y)$-action (automorphism group of $Y$ as Fano cone) with canonical linearization which is trivial on the origin $0=[Y]$. If $({\rm Def}_{\xi<0}(Y)\setminus 0)\curvearrowleft G$ is prestable or semistable in the sense of [Mum65, Chapter I, §4], we define the corresponding GIT quotient $M^{\xi<0}_{Y}:=({\rm Def}_{\xi<0}(Y)\setminus 0)//G$ (resp., $({\rm Def}_{\xi<0,\mathbb{A}^{n}}(Y)\setminus 0)//G$).
Recall $G$ is reductive as proved in [DS17, Appendix] for $\mathbbm{k}=\mathbb{C}$, and for general $\mathbbm{k}$, it is algebraically proved [Od24a, Corollary 3.13].
Question 2.19.
In the above setup of Definition 2.18, does the corresponding GIT quotients
$M^{\xi<0}_{Y}:=({\rm Def}_{\xi<0}(Y)\setminus 0)//G$
admit moduli interpretation of affine $\partial\bar{\partial}$-exact complete Calabi-Yau metrics with Euclidean volume growth? Does the prestability or semistability condition always hold?
Remark 2.20.
For instance, if $G=T=\mathbb{G}_{m}$, the latter is true and $M_{Y}^{\xi<0}$ always comes with a canonical realization as a closed (projective) subscheme inside a weighted projective space. If $Y$ is toric with only isolated singularity, then it is also true by [Alt94, Alt97] (indeed, for $n\geq 4$, the tangent space is trivial). Note that the recent interesting non-existence result of asymptotically conical Kähler metric on $G_{2}$ type complex symmetric spaces $X$ of rank $2$ with a specified asymptotic (horospherical) cone $(Y,\xi)$ [Ngh24, Corollary D] is due to the discrepancy of the (metric) Reeb vector field and the negative valuation i.e., $X\notin{\rm Def}_{\xi}^{-}(Y)$, hence is of different nature from what is discussed here. We thank T.T.Nghiem for the clarification on this point.
2.3. Analogue for Kähler-Ricci solitons - Properness of moduli and bubbling solitons
This subsection takes a brief and temporary detour from the main discussion, for a few pages, to consider another application of our methods to more global problem. That is, we apply to the (expected) properness problem of moduli of compact shrinking Kähler-Ricci soliton (cf., [PSS23, GPSS18, CSW18, Ino19a, HL20, BLXZ23]), and provides a global analogue of bubbling in that context.
In particular, we prove the expected properness modulo some natural but technical input from birational geometry of K-stability. The reason this material is put here is that technically its proof is very similar to the above proofs of Theorem 2.4, 2.6 (and [Od24a]) and is even somewhat easier, since the objects are proper rather than affine. In this subsection, we assume the base field $\mathbbm{k}$ is algebraically closed field of characteristic 00.
Recall that the work of Chen-Sun-Wang on Fano manifolds [CSW18] provides another $2$-step degeneration
$X\rightsquigarrow(T\curvearrowright X^{\prime},\xi^{\prime})\rightsquigarrow(T\curvearrowright X^{\prime\prime},\xi^{\prime\prime}),$
where $(T\curvearrowright X^{\prime},\xi^{\prime})$ and $(T\curvearrowright X^{\prime\prime},\xi^{\prime\prime})$ are (modified) K-semistable and K-polystable in the sense of [BN14] respectively, hence the latter admitting (weak) Kähler-Ricci solitons as the polarized limit space (Gromov-Hausdorff limit) of the normalized Kähler-Ricci flow. Later it is more algebraically reformulated by the work of [DeSz20, Ino19a, HL20, BLXZ23]. In particular, Dervan-Szekelyhidi [DeSz20] introduces  [^13] Here, we use the opposite sign convention of [DeSz20] following [HL20, BLXZ23]. the $h$-invariant $h(X)\in\mathbb{R}_{\leq 0}$ for any $\mathbb{Q}$-Fano variety $X$ ([HL20] for the kawamata-log-terminal singular pairs case).
So, given our construction [Od24a] of the moduli of possibly irregular Calabi-Yau cones, especially the proof of the properness by using the higher $\Theta$-stable reduction [Od24b], it now becomes natural to expect that an analogous method may work for the moduli of Fano varieties with (weak) shrinking Kähler-Ricci solitons.
As the setup, let $N$ be a free abelian group of finite rank and we consider an algebraic $\mathbbm{k}$-torus $T=N\otimes\mathbb{G}_{m}$ which acts on log terminal $\mathbb{Q}$-Fano varieties $X$ of dimension $n$, together with $\xi\in N\otimes\mathbb{R}$ so that the triple $(T\curvearrowright X,\xi)$ is K-semistable and its anticanonical volume $(-K_{X})^{n}$ is $V$ (or consider its natural logarithmic generalization, for fixed coefficients set). We consider their moduli Artin stack $\mathcal{M}^{s}_{n,V,T,\xi}$, as an easy proof of its existence is included in our Theorem 2.22 for readers convenience, following e.g., [PSS23, Ino19a, BLXZ23]. We also fix $n,V,h$ and consider the finite type moduli Artin stack $\mathcal{M}_{n,V,h}$ (resp., $\mathcal{M}_{n,V,h^{+}}$), after [BLXZ23, 6.1, 6.4], of $n$-dimensional $\mathbb{Q}$-Fano varieties $X$ with anticanonical volume $(-K_{X})^{n}=V$ and $h(X)=h$ (resp., $h(X)\geq h$).
The latter is (essentially) constructed in [BLXZ23, §6, Theorem 6.1 (and 6.4)], as op.cit Theorem 6.1 proves the existence of $\mathcal{M}^{\rm Fano}_{n,V,(1,)h^{+}}$ which is our $\mathcal{M}_{n,V,h^{+}}$, and it follows from their Theorem 6.4 that $\mathcal{M}_{n,V,h}$ is a locally closed substack of $\mathcal{M}_{n,V,h^{+}}$. Indeed, by the same theorem, it even follows that $\mathcal{M}_{n,V,h^{+}}$ has a finite stratification by of $\mathcal{M}_{n,V,h^{\prime}}$ for finite possibilities of $h^{\prime}(\geq h)$. Now we hope that the following analogue of [Che24, 1.1], [Od24a, Theorem 3.5] hold.
Conjecture 2.21 (Higher $\Theta$-stratification by $h$-invariant of Fano varieties).
The $h$-invariants of $\mathbb{Q}$-Fano families and the corresponding first step degeneration of [CSW18, BLXZ23] (as reformulated in [HL20, §2], [Od24b, Example 2.19, Corollary 2.17]) satisfies the following general properties.
  1. ($\mathrm{i}$)
    Consider any finite type $\mathbbm{k}$-scheme $S$ and a $\mathbb{Q}$-Gorenstein 141414locally stable in the sense of [Kol23] family $\pi\colon\mathcal{X}\to S$, whose $h$-invariants and volumes are both constant i.e., $h(X_{s})$, $(-K_{X_{s}})^{n}$ does not depend on closed points $s\in S$, where $X_{s}:=\pi^{-1}(s)$. Then, there is a finite rank free abelian group $N$ and a rational polyhedral cone $\tau\subset N\otimes\mathbb{R}$ together with $\xi\in S$ such that there is a $\mathbb{Q}$-Gorenstein family $\tilde{\pi}\colon\tilde{\mathcal{X}}\to U_{\tau}\times S$ which restricts to $\pi$ (resp., the corresponding first step degeneration of $X_{s}$ [CSW18, BLXZ23]) on $(1,\cdots,1)\times S$ (resp., $\Delta_{\xi}\times s\to U_{\tau}\times s$, on each $s$). Here, $U_{\tau}$ means the affine toric variety corresponding to $N_{\mathbb{R}}\supset\tau$ and $\Delta_{\xi}$ is the N-spectrum in [Od24b, §2] (cf., also [HL20, §2]).
  2. ($\mathrm{ii}$)
    For a positive integer $n$ (the dimension), real numbers $V$ (the anticanonical volume, hence positive) and $h$ (lower bound for the $h$-invariant), the finite type moduli Artin stack of $\mathbb{Q}$-Fano varieties $X$ with $h(X)\geq h$ $\mathcal{M}_{n,V,h^{+}}$ (cf., [BLXZ23, 6.1]) and its natural finite stratification $\{\mathcal{M}_{n,V,h^{\prime}}\}_{h^{\prime}(\geq h)}$ by the values of $h(X)$ ([BLXZ23, Theorem 6.4]) admits the structure of a higher $\Theta$-stratification in the sense of [Od24b, §3.3], by encoding the aforementioned first step degeneration of $\mathbb{Q}$-Fano varieties [CSW18, BLXZ23] (reformulated in [Od24b, Ex 2.19]).
Note that (the special case when $S$ is a smooth curve of) the above ($\mathrm{i}$) easily implies ($\mathrm{ii}$). Indeed, because the $H^{\rm NA}$-invariant for a special $\mathbb{R}$-test configuration only depends on the central fiber with the vector field (cf., [HL20, Example 2.32 (87), §2.5 (73)]), together with the fact that the Duistermaat-Heckman measure of $\mathbb{R}$-filtration is determined by the dimensions of the each filters (cf., [BC11], [Od24b, §2.6]), such $\tilde{\pi}$ can exists only if $h(X_{s})$ is constant. The rest follows from the similar arguments as [Od24a, 3.13 (§3.3, §3.4)] after [HL14, AHR20]. See also a related criterion of the $\Theta$-stratification in [HL14], given by two conditions (S) and (B) therein.
Then if the above ($\mathrm{ii}$) holds, now that [Od24b] prepared, the higher $\Theta$-stable reduction theorem §3 loc.cit applies and we readily obtain its application as below. This is similar to [Od24a, §3] for the Calabi-Yau cone case and our Theorems 2.4, 2.6.
Theorem 2.22 (Compactness).
Under the above setup, we assume the conjecture 2.21 holds. Then, both the moduli algebraic $\mathbbm{k}$-stacks $\mathcal{M}_{n,V,h}$ and $\mathcal{M}^{s}_{n,V,T,\xi}$ are universally closed. 151515i.e., satisties the existence part of the valuative criterion of properness of an algebraic stack. The actual properness requires separatedness in some of the literature.
Also, the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ of (modified) K-semistable pair $(X\curvearrowleft T,\xi)$ exists. If $\mathbbm{k}=\mathbb{C}$ and the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ admits a categorical good  [^16] We refer to [Stacksproject, 106.12] (resp., [Alp13]) for the definition of categorical moduli (resp., good moduli). Before, the author has often used the term “coarse moduli” for the exactly same meaning as the categorical moduli in older papers, i.e., in a broader sense than e.g., [Mum65], not assuming the condition of bijectivity of geometric points as in op.cit. moduli space $M^{s}_{n,V,T,\xi}$, its complex analytification $M^{s}_{n,V,T,\xi}(\mathbb{C})$ is compact.
Remark 2.23.
In particular, by setting $h:=0$, we give a yet another proof of the properness of K-moduli stack of $\mathbb{Q}$-Fano varieties (modulo Conjecture 2.21), different from [BHLLX21, LXZ22] which uses the $\delta$-invariant [FO18, BJ18].
Proof.
Consider any stalk of the structure sheaf on a complete $\mathbbm{k}$-curve at a closed point and denote the DVR as $R$, its fraction field (the function field) as $K$. Take an arbitrary $\mathbbm{k}$-morphism $\varphi^{o}\colon{\rm Spec}K\to\mathcal{M}_{n,V,h_{1}=h}$, corresponding to a $\mathbb{Q}$-Fano family $\pi^{o}\colon\mathcal{X}^{o}\to{\rm Spec}(K)$. By [LX18], after a finite base change (we keep the notation of $R,K$ though), $\pi^{o}$ extends to a $\mathbb{Q}$-Fano family $\pi\colon\mathcal{X}\to{\rm Spec}(R)$. We suppose the $h$-invariant of the central fiber (resp., generic fiber) of $\pi$ is $h_{0}$ (resp., $h_{1}$). By [BLXZ23], we have $h_{0}\leq h_{1}$. Correspondingly, so that $\varphi^{o}$ extends to a $\mathbbm{k}$-morphism from ${\rm Spec}(R)$ to $\mathcal{M}_{n,V,h_{0}^{+}}$ for some $h_{0}\leq h_{1}$.
What we need to prove is that, if $h_{0}<h_{1}$, we have another extension of $\varphi^{o}$ to $\mathcal{M}_{n,V,h_{1}=h}$ as $\varphi\colon{\rm Spec}(R)\to\mathcal{M}_{n,V,h_{1}=h}$ possibly after finite extension of $R$. If conjecture (2.21) ($\mathrm{ii}$) is true, then the assertion is a direct corollary to the the higher $\Theta$-stable reduction along the higher $\Theta$-stratification [Od24b, §3.3, Theorem 3.18]. Note that by boundedness results obtained in [BLXZ23, Jia20], possible values of $h(X)$ for $X\in\mathcal{M}_{n,V,h^{+}}$ are finite.
The existence of $\mathcal{M}^{s}_{n,V,T,\xi}$ as an algebraic $\mathbbm{k}$-stack follows similarly as [Od24a]: it is equivalent to the stack of $T$-equivariant $\mathbb{Q}$-Gorenstein families of $n$-dimensional $\mathbb{Q}$-Fano varieties of volume $V$, whose $H^{\rm NA}$-minimizer is $\xi$, by its uniqueness ([HL20, 1.2], [BLXZ23, 1.2]), and the constructible lower semicontinuity of the $h$-invariant ([BLXZ23, 6.4]). Indeed, the whole stack $\mathcal{M}^{T}$ of $T$-equivariant $\mathbb{Q}$-Gorenstein families of $n$-dimensional $\mathbb{Q}$-Fano varieties $X$ of volume $V$ is an algebraic $\mathbbm{k}$-stack. This comes from the fact that the classical affine cone $C(-NK_{X}):={\rm Spec}\oplus_{m\geq 0}H^{0}(-mNK_{X})$ for a uniform sufficiently divisible positive integer $N$, together with the action of $T\times\mathbb{G}_{m}$ on it, recovers $X$. Indeed, the data $C(-NK_{X})\curvearrowleft(T\times\mathbb{G}_{m})$ forms an algebraic stack due to the existence of the (quasi-projective) multi-Hilbert scheme [HM04] and the hulls and husks construction by Kollár [Kol08]. Then our $\mathcal{M}^{s}_{n,V,T,\xi}$ is straightforwardly seen to be an open substack of $\mathcal{M}^{T}$ by the aforementioned results in [HL20, BLXZ23].
Take a connected component $\mathcal{M}$ of $\mathcal{M}^{s}_{n,V,T,\xi}$, put the corresponding $h$-invariant as $h$ and consider the connected component of $\mathcal{M}_{n,V,h}$ which includes the natural image of $\mathcal{M}$, which we denote as $\mathcal{M}_{h}$. Therefore, the forgetful morphism $\mathcal{M}\to\mathcal{M}_{h}$ between the two algebraic $\mathbbm{k}$-stacks is radiciel and finite. Since the latter is universally closed, so is the former. We complete the proof. ∎
By the above compactness theorem, we can expect very similar methods of constructing the good moduli space $\mathcal{M}^{s}_{n,V,T,\xi}$ of (weak) Kähler-Ricci solitons, to the construction of the good moduli space of the moduli stack of the Calabi-Yau cones in [Od24a]. Also note that the above algebraic compactness result is compatible with the precompactness type result by [PSS23, GPSS18].
2.4. Construction of minimal bubble via coordinates
Now, finishing the digression in the previous subsection, we come back to the original context of affine klt varieties and the local bubbling problem.
In this subsection, we discuss a priori different (but conjecturally equivalent) and more practical reconstruction of the minimum bubble, in terms of coordinates and defining equations. From what follows, for notational simplicity, we work over the Setup 1 (b) $\land$ 2 $\land$ 3. It is almost verbatim, by setting $R:=\mathbb{C}[[t]]^{\rm conv}$, how to discuss a complex analytic version for the Setup 1 (a) $\land$ 2 $\land$ 3.
As in [Od24a, Od24b], we assume (without loss of generality) $X$ and $W$ are embedded into the affine space $\mathbb{A}_{\mathbbm{k}}^{l}$ and the isotrivial degeneration $X\rightsquigarrow W$ is obtained by weights $(w_{1},\cdots,w_{l})\in\mathbb{R}_{>0}^{l}$ corresponding to the coordinates $z_{i}$ of $\mathbb{A}_{\mathbbm{k}}^{l}$. We would like to also put a weight on $t$, the base coordinate, as $w_{0}\in\mathbb{R}_{>0}$.
We further fix the extended embedding of $\mathcal{X}$ into $\Delta\times\mathbb{A}^{l}$ which maps $\sigma$ to the 00-section. By the arguments of proof of Theorem 2.4 and Appendix, this setup is not restrictive. Now, we consider
$\mathcal{W}:=\mathcal{W}(w_{0};w_{1},\cdots,w_{l}):=\mathop{\mathrm{Spec}}\nolimits(gr_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})),$
as a family over $\mathbb{A}^{1}_{t}$ where $t$ is the coordinate of $\Delta$, as in [dBS24]. Here, $gr_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ is the natural graded ring
(11) $\oplus_{m\geq 0}\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m\}/\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)>m\},$
where for $f=\sum a_{i_{0},i_{1},\cdots,i_{N}}t^{i_{0}}\prod_{1\leq j\leq N}z_{j}^{i_{j}}$, the weight $w(f)$ of $f$ means $\min\{\sum_{j}w_{j}i_{j}\mid a_{i_{0},i_{1},\cdots,i_{N}}\neq 0\}$ and $\overline{f}$ means the restriction. Note that this construction (11) is a variant of the graded ring discussed in e.g., [LXZ22, XZ22] but its finite typeness over $\mathbbm{k}$ is easy in this setup. Indeed, if the ideal sheaf for $\mathcal{X}\subset\mathbb{A}^{k}\times\Delta$ is $I$, and $w(\mathcal{O}_{\Delta\times\mathbb{A}^{k},0})$ consists of $m_{1}<m_{2}<\cdots$, if we set $\{f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m_{i}\}=:R_{i}$, we have
(12) $\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m_{i}\}/\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m_{i+1}\}$
(13) $=R_{i}/(R_{i+1}+\{f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m_{i}\}\cap I)\twoheadleftarrow R_{i}/R_{i+1},$
for each $i=1,2,\cdots$ so that ${\rm gr}_{w_{0},\cdots,w_{l}}(\mathcal{O}_{\mathbb{A}^{k},0})=\oplus_{i}R_{i}/R_{i+1}\simeq\mathbbm{k}[z_{1},\cdots,z_{l},t]$ is a finite type algebra over $\mathbbm{k}$.
Proposition 2.24.
In the above context,
  1. ($\mathrm{i}$)
    there is a (natural) critical value $w_{\rm cri}\in\mathbb{R}_{>0}$, such that
    • $w_{0}>w_{\rm cri}$, then $\mathcal{W}$ is canonically isomorphic to $\mathbb{A}^{1}_{t}\times W$ over $\mathbb{A}^{1}_{t}$.
    • $w_{0}=w_{\rm cri}$, then $\mathcal{W}$ is a normal affine faithfully flat and $\mathbb{G}_{m}$-equivariant family with the central fiber $W$, where $\mathbb{G}_{m}$ acts on the base $\mathbb{A}_{t}^{1}$ with the weight $w_{0}$.
  2. ($\mathrm{ii}$)
    In particular, for $w_{0}\geq w_{\rm cri}$, the weights $w_{0},w_{1},\cdots,w_{l}$ gives a unique valuation on $\mathcal{O}_{\mathcal{X},x}$.
This proposition is inspired by [Sun23, dBS24] and the discussions with the authors.
Proof.
As we assumed in the beginning of this subsection, $\mathcal{X}\subset\Delta\times\mathbb{A}^{l}$ is cut out by some defining polynomials $F_{1},\cdots,F_{m}$ of $z_{1},\cdots,z_{N},t$. We (again) take enough $F_{i}$s to form a universal Gröbner basis of the generating ideal ([BCRV22, §5.1]). As $x=0\in\mathcal{X}$, these do not contain constant terms. Then $\mathcal{W}\subset\mathbb{A}_{t}^{1}\times\mathbb{A}^{l}$ is cut out by the initial terms ${\rm in}_{w_{0},w_{1},\cdots,w_{l}}(F_{i})$ of $F_{i}$ with respect to the weights $w_{0},w_{1},\cdots,w_{l}$. Thus in particular, it admits the natural $\mathbb{G}_{m}$-action which equivariantly projects down to $\mathbb{A}_{t}^{1}$ with the weight $w_{0}$.
We set $f_{i}:=F_{i}|_{t=0}$. Since $F_{i}=f_{i}+tg_{i}$ for some polynomials $g_{i}$s, for large enough $w_{0}$, there is a critical value $w_{\rm cri}$ such that for the initial term of $F_{i}$ is $f_{i}$ for any $i$ if and only if $w_{0}>w_{\rm cri}$. From the construction, if $w_{0}=w_{\rm cri}$, ${\rm in}_{w_{0},w_{1},\cdots,w_{l}}(F_{i})|_{t=0}=f_{i}$ for each $i$ so that the central fiber of $\mathcal{W}$ is still $W$. For general weights $w_{0},\cdots,w_{l}$, $gr_{w_{0},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ may not be flat over $\mathbb{A}^{1}$, but in our case, since the central fiber is normal, $n$-dimensional, and $\mathcal{W}$ is $\mathbb{G}_{m}$-acted, $\mathcal{W}(w_{\rm cri},w_{1},\cdots,w_{l})$ is normal and faithfully flat over $\mathbb{A}^{1}$. This completes the proof of the first claim ($\mathrm{i}$) of the proposition. In particular, $gr_{w_{\rm cri},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ is integral so that $w_{\rm cri},w_{1},\cdots,w_{l}$ naturally gives a valuation on $\mathcal{O}_{\mathcal{X},x}$ (cf., [Tei14, proof of Proposition 3.3. a) and comments]). This completes the proof of the item ($\mathrm{ii}$). ∎
We now prepare a key observation to connect the previous deformation theoretic construction of bubbling (Theorem 2.4, 2.6) to approaches via coordinates (or valuations).
In particular, we analyze a general structure of $T$-equivariant morphism $\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})\to M$ (see [Od24b, §3] and the previous section for the notation) which appears in the proof of Theorem 2.4. Here, $\tau^{\prime}$ is a rational polyhedral cone of $N\oplus\mathbb{Z}$ which projects as PL isomorphism (in particular, bijectively) to $\tau\subset N_{\mathbb{R}}$. The rational polyhedral cone $\tau^{\prime}$ includes a vector $(w_{0},\xi)\in(\mathbb{Z}\oplus N)\otimes\mathbb{R}$, which projects to $\xi\in\tau\subset N\otimes\mathbb{R}$.
In the context (proof) of Theorem 2.4, we were concerned with particular $w_{0}$ which is naturally determined by the other data including $\xi$. Nevertheless, here we consider arbitrary $w_{0}$ and corresponding $T$-equivariant morphism $\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})\to M$ and analyze the obtained degenerating families from different perspectives.
Lemma 2.25 (Relation with weighted blow ups).
For $T\curvearrowright M$ and $\Delta\to M$ in the proof of Theorem 2.4, where the latter specifies $\mathcal{X}\to\Delta$ and its (00-)section $\sigma$, we consider an arbitrary $T(=N\otimes\mathbb{G}_{m})$-equivariant morphism $\overline{f}\colon\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})\to M$ for some $\tau^{\prime}$ as the previous two paragraphs right above here. We take a primitive integral $\xi^{\prime}=(\xi,w_{0})\in(\tau^{\prime}\cap N\oplus\mathbb{Z})$ (thus approximating the original direction $\xi^{\prime}$ of the arguments in [Od24b, 3.8] with rational rank $r$, in general) which projects to integral $\xi\in\tau\cap N$.
  1. ($\mathrm{i}$)
    We consider the rank $1$ algebraic subtorus of $T\times\mathbb{G}_{m}$ (resp., the rank $1$ algebraic subtorus of $T$) associated to $\mathbb{R}\xi^{\prime}\cap(N\oplus\mathbb{Z})$ (resp., $\mathbb{R}\xi\cap N$) as $(\mathbb{G}_{m}\simeq)T_{\xi^{\prime}}\subset T\times\mathbb{G}_{m}$ (resp., $(\mathbb{G}_{m}\simeq)T_{\xi}\subset T)$). Consider the corresponding affine family $\tilde{\mathcal{X}}_{\overline{f},\tau^{\prime}}\to\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})$ to $\overline{f}$, restrict it to the closure of $T_{\xi^{\prime}}$-orbit in $\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})$ and denote as $\mathcal{X}^{\prime}(w_{0},\tau^{\prime})\to\mathbb{A}^{1}$.
    Then, if $w_{0}=1$, $\mathcal{X}^{\prime}(w_{0}=1,\tau^{\prime})$ is the weighted blow up of $\mathcal{X}$ for weights $(w_{0},w_{1},\cdots,w_{l})$. For general $w_{0}\geq 1$, $\mathcal{X}^{\prime}(w_{0},\tau^{\prime})$ is the normalization of the base change of the weighted blow up of $\mathcal{X}$ for weights $(w_{0},w_{1},\cdots,w_{l})$ with respect to the $w_{0}$-th power map $\Delta\xrightarrow{s\mapsto s^{w_{0}}=t}\Delta$. In particular, for large enough $w_{0}$, the central fiber (over 00) of $\mathcal{X}^{\prime}(w_{0},\tau^{\prime})\to\mathbb{A}^{1}$ does not depend on $w_{0}$.
  2. ($\mathrm{ii}$)
[…]
By proof of Theorem 2.4, $\mathcal{X}_{\rm min}^{\prime}$ and $\mathcal{W}$ obtained therein can be regarded as particular cases of the above construction ($\mathrm{i}$), ($\mathrm{ii}$).
proof of Lemma 2.25.
Firstly, note that $\tilde{\mathcal{X}}_{\overline{f},\tau^{\prime}}\to\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})$ is affine and faithfully-flat finite type from the definition. We first prove ($\mathrm{i}$). Take the weights of $T_{\xi}$ on the coordinates $z_{1},\cdots,z_{l}$ of $\mathbb{A}^{l}$ as $w_{1},\cdots,w_{l}$. Then, for a fixed $c\in\mathbbm{k}^{*}$, $\mathcal{X}^{\prime}(w_{0},\tau^{\prime})$ is isomorphic to $\overline{\{(t^{-w_{0}},\cdots,t^{-w_{l}})\cdot\mathcal{X}_{t^{w_{0}}c}\}}_{t\in(\Delta\setminus 0)}$ as closed subschemes of $\mathbb{A}^{l}\times_{\mathbbm{k}}\Delta=\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$, where $\cdot$ means the $T_{\xi}$-action, $0\in\Delta$ means the closed point (the origin) and the overline means the (Zariski) closure inside $\mathbb{A}_{z_{1},\cdots,z_{l}}^{l}\times_{\mathbbm{k}}\Delta$. Note that presence of $T_{\xi}$-action on $\mathbb{A}_{\mathbbm{k}}^{l}$ makes the above construction independent of $c$. If $w_{0}=1$, equivalently, it ($\mathcal{X}^{\prime}(w_{0}=1,\tau^{\prime})$) is the restriction of the Zariski closure of $\mathcal{X}\setminus X$ in $\mathbb{A}^{l}_{\frac{z_{1}}{t^{w_{1}}},\cdots,\frac{z_{l}}{t^{w_{l}}}}$ (note the different coordinates), and equivalently, it is nothing but the weighted blow up of $\mathcal{X}$ with weights $(w_{0}=1,w_{1},\cdots,w_{l})$. Hence the assertion of ($\mathrm{i}$) holds in this case. For general $w_{0}$ case, let us denote the $w_{0}$-th power map of the base $\Delta_{s}\xrightarrow{s\mapsto s^{w_{0}}=t}\Delta_{t}$, which we use for base change. Then the normalization of ${\rm Bl}_{(w_{0},\cdots,w_{l})}(\mathbb{A}_{\mathbbm{k}[[t]]}^{l})\times_{\mathbbm{k}[[t]]}\mathbbm{k}[[s]]$ is ${\rm Bl}_{(1,w_{1},\cdots,w_{l})}(\mathbb{A}_{\mathbbm{k}[[s]]}^{l})$ as it is easy to see ring-theoretically. Therefore, as closed their subschemes, there is a finite morphism from ${\rm Bl}_{(1,w_{1},\cdots,w_{l})}(\mathcal{X}\times_{\mathbbm{k}[[t]]}\mathbbm{k}[[s]])$ to ${\rm Bl}_{(w_{0},w_{1},\cdots,w_{l})}(\mathcal{X})\times_{\mathbbm{k}[[t]]}\mathbbm{k}[[s]]$ which is automatically the normalization. Indeed, note that the former is normal because of the previous arguments for $w_{0}=1$ case and the fact that it is normal as the total space of a fiberwise normal family by our assumption on $M$. Hence, the assertion of ($\mathrm{i}$) follows from the case of $w_{0}=1$ and we conclude its proof, except for the last statements, which now follows from Proposition 2.24.
Next, we prove ($\mathrm{ii}$). When we consider small enough perturbation of $(w_{0},w_{1},\cdots,w_{l})$ which takes values in $\mathbb{Q}\oplus N_{\mathbb{Q}}$, note that $\lim_{t\to 0}(t^{-dw_{0}},\cdots,t^{-dw_{l}})\cdot X_{t^{dw_{0}}c}$ for $d\in\mathbb{Z}_{>0}$ such that $dw_{i}\in\mathbb{Z}$ (in the sense of Zariski closure inside $\mathbb{A}^{l}_{\mathbbm{k}}\times_{\mathbbm{k}}\Delta=\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$) remain the same. This fact simply follows from the definition of $\tilde{\mathcal{X}}_{\overline{f},\tau^{\prime}}\to\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})$ as what corresponds to $\overline{f}$. On the other hand, by our definition of graded ring in (11) (compare analogous [LX18, 2.10]) and our arguments in the proof of Proposition 2.24, ${\rm gr}_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ also does not change as a ring (ignoring the grading) for a certain perturbation to $w_{0},w_{1},\cdots,w_{l}\in\mathbb{Q}\oplus N_{\mathbb{Q}}$. Indeed, it also follows from the description of ${\rm gr}_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ in terms of initial ideals (see [Eis13, Chapter 15]). So we can and do assume $w_{i}$ are all integers. Then note that for any $c\in\mathbbm{k}$, admitting $c=0$, $\lim_{t\to 0}(t^{w_{0}},\cdots,t^{w_{l}})\cdot X_{c\cdot t^{w_{0}}}\subset\mathcal{W}_{c}$ for any $c$ by the proof of the above ($\mathrm{i}$). Here, the above limit is as close subscheme of $\mathbb{A}^{l}$. Therefore,
$\lim_{t\to 0}(t^{w_{0}},\cdots,t^{w_{l}})\cdot\mathcal{X}=\mathcal{W}(w_{0},\cdots,w_{l})\times_{\mathbb{A}^{1}_{\mathbbm{k}}}\mathbbm{k}[[t]],$
as closed subschemes of $\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$, since the left hand side is contained in the right hand side while both are irreducible hence they are the same. As they are both restrictions of affine test configurations to over $\mathbbm{k}[[t]]$, one can extend the above isomorphism via the horizontal $\mathbb{G}_{m}$-action to $\mathcal{W}\simeq{\rm Spec}({\rm gr}_{\xi^{\prime}}(\mathcal{O}_{\mathcal{X},x}))$.
Now, we apply the above Lemma 2.25 to the situation of (proof of) Theorem 2.4.
Corollary 2.26 (cf., Conjectures in [dBS24, Sun23]).
In the Setup 1, for a good choice of embedding of $\mathcal{X}$ to $\Delta\times\mathbb{A}_{\mathbbm{k}}^{l}$ and $w_{1},\cdots,w_{l}$, $\mathcal{W}(w_{\rm cri};w_{1},\cdots,w_{l})\to\mathbb{A}^{1}$ (of Proposition 2.24) is a test configuration $\mathcal{T}$, which degenerates the central fiber $\mathcal{X}^{\prime}_{\rm min,0}$ of $\mathcal{X}$ in Theorem 2.4 isotrivially to $W$ (compare the proof of Theorem 2.4).
proof of Corollary 2.26.
Firstly, consider the valuation $v_{X}$ of $\mathcal{O}_{X,x}$ which minimizes $\widehat{\rm vol}(-)$. It exists by [Blu18] and is quasi-monomial [Xu20] and further has a finitely generated graded ring ${\rm gr}_{v_{X}}(\mathcal{O}_{X,x})$ [XZ22], as it matches to the differential geometric perspectives provided in [DS17]. Therefore, as explained in [Od24a, §2], one can take homogeneous generators of ${\rm gr}_{v_{X}}(\mathcal{O}_{X,x})$ and lifts to global sections of $\mathcal{O}_{\mathcal{X}}$ which embeds $\mathcal{X}$ into $\mathbb{A}^{1}\times\mathbb{A}^{l}$. Then, there are weights $w_{1},\cdots,w_{l}$ on the fiber-direction coordinates on $\mathbb{A}^{l}$ and the according construction in the proof of Theorem 2.4 in particular gives $w_{0}$ so that corresponding family $\mathcal{W}(w_{0},\cdots,w_{l})$ is the (nontrivial) negative weight deformation in Theorem 2.4 ($\mathrm{iv}$) with the general fiber $\mathcal{X}^{\prime}_{\rm min,0}$. Further, from the construction of proof of Theorem 2.4, it follows that for any $0<w<w_{0}$, we have $\mathcal{W}(w,w_{1},\cdots,w_{l})$ is $W\times\mathbb{A}^{1}$. On the other hand, by Lemma 2.25 ($\mathrm{ii}$), this $\mathcal{W}(w,\cdots,w_{l})$ can be understood as ${\rm gr}_{w,\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ in the sense of (11) in general. Therefore, combining with Proposition 2.24, we obtain that $w_{0}=w_{\rm cri}$ so that Corollary 2.26 follows. ∎
The above corollary allows explicit descriptions of the K-semistable bubbling $\mathcal{X}^{\prime}_{\rm min,0}$ in terms of weighted blow ups, and is inspired by discussions with deBorbon-Spotti (cf., [dBS24, Remark 11]), S.Sun and S.Donaldson. The author appreciates them. Note also that [dBS24] implicitly confirmed the above conjecture in classical examples in low dimension which inspire above (we also did some similar calculations to loc.cit, confirming the above in higher dimensional examples e.g. generic hypersurfaces). Indeed, in order to obtain the differential minimum bubble in general, we need a good choice of extension of the embedding $X$ into $\mathbb{C}^{l}$ to $\mathcal{X}$ to $\Delta\times\mathbb{C}^{l}$ as the construction of [Sun23, Lemma 4.6] suggests.
2.5. Valuative approach
Taking Theorem 2.4 ($\mathrm{x}$) and Proposition 2.24 ($\mathrm{ii}$) into account, for instance, it is reasonable to pass our context to general valuations and discuss characterization of the critical valuations in the sense of Theorem 2.4 ($\mathrm{x}$). Our statements in this subsection are totally coordinates-free.
We set the space of real valuations $v$ (resp., $v_{X}$) of $\mathcal{O}_{X,x}$ (resp., $\mathcal{O}_{X,x}$) whose center $c_{\mathcal{X}}(v)$ (resp., $c_{X}(v)$) is $x\in\mathcal{X}$ (resp., $x\in X$) as $\mathop{\mathrm{Val}}\nolimits(x\in\mathcal{X})$ (resp., $\mathop{\mathrm{Val}}\nolimits(x\in X)$). Their subsets which consist of quasi-monomial valuations will be denoted by $\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in\mathcal{X})$ (resp., $\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in X)$).
For each $v\in\mathop{\mathrm{Val}}\nolimits(x\in\mathcal{X})$ (resp., $v_{X}\in\mathop{\mathrm{Val}}\nolimits(x\in X)$), we have valuation ideals
$\mathcal{J}_{m}(v)$ :=$:=\{F\in\mathcal{O}_{\mathcal{X},x}\mid v(F)\geq m\}$
[…]
as a sequence of coherent ideals. For later convenience, we also set $\mathcal{J}_{>m}(v):=\{f\in\mathcal{O}_{\mathcal{X},x}\mid v(f)>m\}$. The following may evoke the discussion in [Li19, §6] after [MSY08]. S.Boucksom pointed out it is somewhat analogous to the Gauss extension in the isotrivial case (cf., e.g., [BJ18, §1]) in spirit.
Proposition 2.27.
  1. ($\mathrm{i}$)
    For each $v_{X}\in\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in X)$, $a\in\mathbb{R}_{>0}$, with $a\ll 1$, say $a\in(0,a_{0})$ for some $a_{0}\in\mathbb{R}_{>0}$, we have a unique valuation $v\in\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in\mathcal{X})$ with $\mathcal{J}_{m}(v_{a})|_{X}=J_{m}(v_{X})$ for all $m$ (or equivalently, for all $m\gg 0$) and $v_{a}(t)=\frac{1}{a}$.
  2. ($\mathrm{ii}$)
    Once one fixes a holomorphic section $\sigma\colon\Delta\to\mathcal{X}$, and a SNC model of $X$ (resp., [BFJ08, BFJ16]) which realizes $v$ as a quasi-monomial valuation, then in a longer interval (than ($\mathrm{i}$)) of $a$ as $a\in(0,a_{1})$ with $a_{1}>a_{0}$, $v_{a}$ of ($\mathrm{i}$) is still canonically determined. Henceforth, we sometimes denote this $a_{1}$ as $a_{1}(\varphi)$.
The reason of our notation of the weight as reciprocal $\frac{1}{a}$ comes from our interest in $a=0$ case as initial point (cf., also Remark 2.28).
Proof.
There is a log resolution $\varphi\colon\tilde{X}\to X$ where the center of $v_{X}$ is a strata of simple normal crossing divisor $S=\cap_{i\in I}E_{i}$. We write $E_{i}=(z_{i}=0)$ with $z_{i}\in\mathcal{O}_{\tilde{X},\eta_{S}}$ for each $i\in I$, where $\eta_{S}$ means the generic point of $S$. By replacing $\varphi$ by a further blow up if necessary, we can and do assume $\varphi$ extends to a log resolution $\tilde{\varphi}\colon\tilde{\mathcal{X}}\to\mathcal{X}$. Then from the quasi-monomiality assumption of $v_{X}$, there is a set of real positive numbers $c_{i}$s that the $v_{X}$ is determined by $E_{i}$s with the condition $v_{X}(z_{i})=c_{i}$. For convenience, we also take other coordinates $z_{i}(i\notin I)$ which vanishes at the center of $v_{X}$ in $S\subset\tilde{X}$ to form an analytic local coordinate system $\{z_{i}\}$. Take an extension $z_{i}$ to $\mathcal{O}_{\tilde{\mathcal{X}},\eta_{S}}$ as $\tilde{z_{i}}$ and consider the quasi-monomial valuation $v_{a}$ determined by $\tilde{\varphi}$, $c_{i}$s and $v_{a}(t)=\frac{1}{a}$. Hence the existence claim is proven.
Next we consider the uniqueness and canonicity of $v_{a}$. For a different choice of $\tilde{z_{i}}$ as $\tilde{z_{i}}^{\prime}$, there is a holomorphic function $f_{i}$ of $z_{j}$s such that
(14) $\tilde{z_{i}}^{\prime}=z_{i}+tf_{i}(z_{1},\cdots,z_{n}).$
If $v_{a}(t)=\frac{1}{a}>v_{a}(z_{i})$, then by (14), the independencity of $v_{a}$ with respect to the lifted coordinates $\tilde{z_{i}}$ easily follows. This confirms ($\mathrm{i}$). On the other hand, if we fix a section $\sigma$, then we can impose the condition $\tilde{z_{i}}(\sigma(t))=0$ for each $i$. Under such constraints, $f_{i}$ of (14) does not have constant terms i.e., positive multiplicities with respect to $\tilde{z_{i}}$s. In that case, if $v_{a}(t)=\frac{1}{a}>\max\{c_{i}\}-\min\{c_{i}\}$, then the obtained quasi-monomial $v_{a}$ does not depend on the lift $\tilde{z_{i}}$s. Hence the proof of ($\mathrm{ii}$). ∎
Remark 2.28.
S.Boucksom kindly suggested to compare the construction with the composite (cf., e.g., [FK18, §6.4]) of the discrete valuation ${\rm ord}_{\mathcal{X}_{0}}$ on $\mathcal{O}_{\mathcal{X},x}$ and $v_{X}$ on $\mathcal{O}_{\mathcal{X}_{0},x}$. Although it gives a height $2$ (not real-valued) valuation, the author sees it as a kind of infinitesimal analogue of the above construction $\{v_{a}\}_{a}$, which may be its integral curve in some unknown sense.
Example 2.29.
Suppose $\mathcal{X}$ has a simultaneous log resolution, such as $2$-dimensional ADE singularity case. Then, we can apply the above proof to that simultaneous resolution so that the central fiber of $\tilde{X}$ is smooth (irreducible) and $\tilde{z}_{i}$ are all different from $t$, so that we see $a_{0}=\infty$. On the other hand [dBS24, Example 7] infers that $a_{0}$ could be finite in general, for $n(>2)$-dimensional $A$-type singularities.
In this paper, we introduce yet another couple of “Yau-Tian-Donaldson type” conjectures for mild degenerations, after [Sun23, dBS24]. It hopefully connects the complex differential geometry of bubble (tree) to our algebro-geometric or non-archimedean framework.
Now, again take a section $\sigma\colon\Delta\to\mathcal{X}$ as Setup 2 and fix it. Then we have a metric analogue of (2.27):
Definition 2.30.
Under the assumption (3), for $a\in\mathbb{R}_{>0}$, we define $v^{\rm DG}_{a}\colon\mathcal{O}_{\mathcal{X},x}\setminus\{0\}\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$  [^17] We do not expect this value can take $\infty$ in reality. as follows. For $G\in\mathcal{O}_{\mathcal{X},x}$, we set
$v^{\rm DG}_{a}(G):=\limsup_{t\to 0}\dfrac{\log\sup_{y\in B_{|t|^{a}}(\sigma(t))}(|G(y)|)}{-\log|t|}.$
Here, $B_{|t|^{a}}(\sigma(t))$ denotes the geodesic ball in $\pi^{-1}(t)=X_{t}$ centered at $\sigma(t)$ with radius $|t|^{a}$.
The following conjecture for the minimum bubble refines the inspiring predictions and discussions in [Sun23, dBS24] via valuations, especially by variation of log discrepancies and the above type valuation (Definition (2.30)).
Conjecture 2.31.
Under the Setup 1 with a fixed closed base point $x\in X$, we now again take the quasi-monomial valuation $v_{X}$ of $\mathcal{O}_{X,x}$ which minimizes $\widehat{\operatorname{vol}}_{X}(-)$ ([DS17, §3.2], [Li19, Blu18]) and apply Proposition 2.27 and Definition 2.30 to consider $v_{a}$ and $v_{a}^{\rm DG}$. Then the following holds:
  1. ($\mathrm{i}$)
[…]
    • The general fibers of a test configuration ${\rm Spec}({\rm gr}_{v_{\rm min}}\mathcal{O}_{\mathcal{X},x})\to\mathbb{A}^{1}$ is the semistable minimal bubble $\mathcal{X}^{\prime}_{{\rm min},0}$. (compare Corollary (2.26)). In relation with it, if $r=1$, then $v_{\rm min}$ should be the minimal critical valuation (Theorem 2.4 ($\mathrm{x}$)).
    • $a\in(0,a_{\rm min}]$, $A_{(\mathcal{X},X)}(v_{a})$ stays constant (cf., Proposition 2.32 below) while
    • $A_{(\mathcal{X},X)}(v_{a})>A_{(\mathcal{X},X)}(v_{a_{\rm min}})$ for $a>a_{\rm min}$.
  2. ($\mathrm{ii}$)
    After suitably multiplying $v_{X}$ by a positive constant, $(\sigma(t)\in X_{t},|t|^{-a_{\rm min}}g_{t})$ converges to a minimum bubble in the sense of [Sun23].
  3. ($\mathrm{iii}$)
[…]
  4. ($\mathrm{iv}$)
    (family version of local normalized volume) For an arbitray valuation $v$ of $\mathcal{O}_{x,\mathcal{X}}$ with center $x\in\mathcal{X}$, we consider $\widetilde{\operatorname{vol}}_{\mathcal{X}}(v):=(A_{(\mathcal{X},X)}(v))^{n}\cdot\operatorname{vol}_{\mathcal{X}}(v).$ Then, we have  [^19] In the version 1 on arXiv, we include the proof of $\geq$ direction and other related statements. We temporarily cut the details just to make the exposition more concise with clearer focus. $\inf_{c_{\mathcal{X}}(v)=x}\widetilde{\operatorname{vol}}_{\mathcal{X}}(v)=\widehat{\operatorname{vol}}(x\in X).$
The above conjectures are obtained after experimental calculations of various examples, including non-degenerate hypersurfaces. We partially confirm it as follows.
Theorem 2.32.
Under the same setup, Conjecture 2.31 is partially true in the following sense: there is a positive real number $a_{\rm min}$ and a family of quasi-monomial valuations $v_{a}$ of $\mathcal{O}_{\mathcal{X},x}$ ($0<a\leq a_{\rm min}$) such that the first and second bulleted items in ($\mathrm{i}$) hold.
  • (compare Proposition 2.27) The valuative ideals satisfy $\mathcal{J}_{m}(v_{a})|_{X}=J_{m}(v_{X})$ for all $m$ and $v_{a}(t)=\frac{1}{a}$.
  • The general fibers of a test configuration ${\rm Spec}({\rm gr}_{v_{a_{\rm min}}}\mathcal{O}_{\mathcal{X},x})\to\mathbb{A}^{1}$ is the semistable minimal bubble $\mathcal{X}^{\prime}_{{\rm min},0}$. (compare Corollary (2.26)).
  • $A_{(\mathcal{X},X)}(v_{a})$ stays constant for $0<a\leq a_{\rm min}$ (cf., Proposition 2.32 below)
Note that the last item could be roughly regarded as a difficult version of adjunction, once we have $\mathcal{J}_{m}(v_{a})|_{X}=J_{m}(v_{X})$ for all $m$, but is much harder and require $a<a_{\rm min}.$
Proof.
The following depends on the arguments used in previous (sub)sections and [Od24b, §3]. As before, we take the volume-minimizing valuation $v_{X}$, realize it by weights on the coordinates of local embedding of $X$ to affine spaces, consider the corresponding Reeb vector $\xi\in N\otimes\mathbb{R}$ to $v_{X}$ and replace it by an (integral) element of $N$ whose direction approximates the original $\mathbb{R}_{>0}\xi$ i.e., in $\tau\cap N$ in the notation of [Od24b, §3]. It is easy to see that assertions of the above proposition for such replacement $\xi$ implies the same for the original $\xi$.
Recall that the proof of Theorem 2.4 shows that the condition of Lemma 2.25 is satisfied for certain $w_{0}\in\mathbb{Z}_{>0}$, after replacing $\xi$ by its positive integer multiple, so that one can apply it. If we set $a_{\rm min}:=\frac{1}{w_{0}}$, Lemma 2.25 (i) shows that the filtration on $\mathcal{O}_{\mathcal{X},x}$ with respect to the weights $(a\xi,1)$ with $a<a_{\rm min}$ gives a valuation on $\mathcal{O}_{\mathcal{X},x}$, which we denote as $av_{a}$. This is because the spectrum of the graded ring is still isomorphic to the K-semistable Fano cone $W$ after the perturbation of $\xi$ (see e.g., [LX18, 2.10]) and in particular integral. This gives the valuation corresponding to the exceptional divisor of the weighted blow up with the weights $(a\xi,1)$. Further, it follows from the construction that $\mathcal{J}_{m}(v_{a})|_{X}=J_{m}(v_{X})$ for all $m$ and $v_{a}(t)=\frac{1}{a}$.
Below, we show that $A_{(\mathcal{X},X)}(v_{a})$ does not depend on $a$ which are integers. By rescaling $v_{X}$ and base change of $\mathcal{X}\to\Delta$ it implies the constancy of $A_{(\mathcal{X},X)}(v_{a})$ for all rational $a<a_{\rm min}$ hence the same for all real $a<a_{\rm min}$ by its continuity with respect to $a$. Therefore, we can and do assume $a(<a_{\rm min})$ is positive integer and only need to prove the independency of $A_{(\mathcal{X},X)}(v_{a})$. Following Lemma 2.25 (i) we consider the corresponding normalized weighted blow up of $\mathcal{X}$ for the same weights $(a\xi,1)$ which we denote as $\tilde{\rm Bl}_{(a\xi,1)}(\mathcal{X})(\to\mathcal{X})\to\Delta$. Its central fiber consists of two irreducible components i.e., the strict transform $X^{\prime}$ of $X$ and a compactification of $W$. Note that since the latter is the exceptional divisor and the total central fiber is principal divisor of $\tilde{\rm Bl}_{(a\xi,1)}(\mathcal{X})$, both components are $\mathbb{Q}$-Cartier. Further, from [LX20, 1.3], [LX18, 2.10, 4.7], one can and do assume that our approximant $\xi$ corresponds to the exceptional divisor $e$ of a plt blow up of $X$ i.e., ${\rm val}_{\xi}=c{\rm ord}_{e}$ as valuations on $\mathcal{O}_{X,x}$ with some positive real constant $c$. Then, for each $\mathbb{Z}_{>0}\ni a(<a_{\rm min})$, we have
$A_{(\mathcal{X},X)}(v_{a})$ $=\frac{1}{a}A_{(\mathcal{X},X)}({\rm ord}_{W})$
$=\frac{1}{a}A_{X}({\rm ord}_{W}|_{X})$
$=A_{X}(v_{X})=A_{X}(c{\rm ord}_{e}),$
which does not depend on $a$. Thus we confirm the third bulleted item. Also, Lemma 2.25 (ii) implies that the general fibers of ${\rm Spec}({\rm gr}_{v_{a_{\rm min}}}\mathcal{O}_{\mathcal{X},x})\to\mathbb{A}^{1}$ is the semistable minimal bubble $\mathcal{X}^{\prime}_{{\rm min},0}$. We complete the proof of the second bulleted item as well. ∎
3. Discussions on more general bubbles
This last short section takes a few pages to discuss conjectural pictures for future direction. In the first subsection §3.1, we still discuss klt degenerations as in Setup 1 (“type I”, but essentially polarized) and speculate on the whole structure formed by deeper bubbles. Then §3.2 reviews our speculations on the structure of bubblings for so-called maximal degeneration case or intermediate case (“type II” or higher), following [Od22a, Od22b].
3.1. Deeper bubbles for klt degenerations
For the structure formed by all the bubbles for the same klt type degenerations as Setup 1, we only have rougher brave predictions as follows. For convenience of discussions, we implicitly take a rescaling normalization of valuations as $v(t)=1$ in this section i.e., multiply $a$ to those considered in section 2.5.
Conjecture 3.1.
We fix an affine finite type log-terminal faithfully flat morphism $\pi\colon\mathcal{X}\to\Delta^{\rm conv}(\mathbb{C})$, where $\Delta^{\rm conv}(\mathbb{C})$ denotes the convergent spectrum $\mathop{\mathrm{Spec}}\nolimits(\mathbb{C}[[t]]^{\rm conv}:=\mathcal{O}_{\mathbb{A}^{1},0}^{\rm an})$, with $v(t)=1$. (Note that in particular, we have its restrictions of the form $\mathcal{Y}\to\Delta_{\epsilon}:=\{t\in\mathbb{C}\mid|t|<\epsilon\}$ for some $\epsilon>0$ with the same central fibers.) We denote the generic fiber as $\mathcal{X}_{\eta}$ and its Berkovich analytification as $\mathcal{X}_{\eta}^{\rm an}$ Then the following holds.
  1. ($\mathrm{i}$)
    Consider the rescaling equivalence classes of non-flat bubbles at $x\in\mathcal{X}$, which we write $\mathfrak{B}(x\in\mathcal{X})=\{(p\in B,g)\}/\mathbb{R}_{>0}$, and the class $(p\in B,g)$ represents as $[p\in B,g]$. There is a bijective correspondence between $\mathfrak{B}(x\in\mathcal{X})=\{(p\in B,g)\}/\mathbb{R}_{>0}$ and a set of valuations $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})\subset\mathcal{X}_{\eta}^{\rm an}$. We write the latter as $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})$ and the bijective maps as
    $\nu\colon$ $\mathfrak{B}(x\in\mathcal{X})\xrightarrow{\,\smash{\raisebox{-1.50696pt}{$\scriptstyle\simeq$}}\,}\mathop{\mathrm{Crit}}\nolimits(x\in X)\subset\mathcal{X}_{\eta}^{\rm an},$
    $B\colon$ $\mathop{\mathrm{Crit}}\nolimits(x\in X)\xrightarrow{\,\smash{\raisebox{-1.50696pt}{$\scriptstyle\simeq$}}\,}\mathfrak{B}(x\in\mathcal{X}),$
    which are inverse to each other. In particular, the set of bubbles (with both the algebraic structures and metrics) has an natural embedding into a non-archimedean Berkovich analytic space.
  2. ($\mathrm{ii}$)
    $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})\subset\mathcal{X}_{\eta}^{\rm an}$ and compatibly the moduli of bubble classes $\mathfrak{B}(x\in\mathcal{X})$ has a locally PL structure.
  3. ($\mathrm{iii}$)
    $\mathfrak{B}(x\in\mathcal{X})$ and $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})$ only depend on the germ of $\pi$ around $x$.
To make sense of the item ($\mathrm{ii}$) in general, we prepare the following after the general ideas in [Berk99, KS06, BFJ08, BFJ16] for the case $\mathcal{X}_{\eta}$ is smooth.
Lemma 3.2.
In the above setup, we further assume $\mathcal{X}$ descends to the family over $\Delta$ for simplicity. Consider the projective system $(S,>)$ of isomorphism classes of $\varphi\colon\tilde{\mathcal{X}}\to\mathcal{X}$ which run over birational projective morphism such that $E_{x}$ is a divisorial part of $\varphi^{-1}(x)$ and $(\tilde{\mathcal{X}},\tilde{\mathcal{X}}_{0}+E_{x})$ is dlt and $\mathbb{Q}$-factorial, so that the dual intersection complex $\Delta(E_{x})$ is well-defined. (In case $x\in X$ is an isolated singularity, we can instead put stronger assumptions $\tilde{\mathcal{X}}$ is smooth and $\tilde{\mathcal{X}}_{0}+E_{x}$ is simple normal crossing for simplicity.) $(\varphi_{1}\colon\tilde{X}_{1}\to\mathcal{X})>(\varphi_{2}\colon\tilde{X}_{2}\to\mathcal{X})$ is simply defined for domination by a commuting birational morphism $\mathcal{X}_{1}\to\mathcal{X}_{2}$.
  1. ($\mathrm{i}$)
    For each $\tilde{\mathcal{X}}\to\mathcal{X}$ in $S$, consider $\Delta(E_{x})$, the dual intersection complex (of the largest normal crossing open subset of $E_{x}$). Then there is a natural injective continous map $\Delta(E_{x})\hookrightarrow{\rm Val}^{\rm qm,t\mapsto 1}(x\in\mathcal{X})$. Through this map, we identify $\Delta(E_{x})$ as a subset of ${\rm Val}^{\rm qm,t\mapsto 1}(x\in\mathcal{X})$.
  2. ($\mathrm{ii}$)
    Further, we have the following structure:
[…]
    (16) $\subset\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$
[…]
    (18) $\subset\mathcal{X}_{\eta}^{\rm an},$
    where $\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$ refers to the set of valuations of $\mathcal{O}_{\mathcal{X},x}$ with the center $x$ which maps $t$ to $1$, $\varphi\in S$ and the domination “$>$” induces PL maps as in loc.cit. In particular, (17) is a dense inclusion while (18) is not.
Proof.
[…]
3.2. Towards bubbling for semi-log canonical degeneration
We make some further discussions towards understanding of bubbling along worse degenerations than log terminal singularities, following our previous works ([Od22a, Od22b]). This subsection is independent of other (sub)sections. As a common setup of diverging or collapsing Ricci-flat Kähler metrics in the literature, one often consider the cases when the section hits a non-log-terminal, but strictly (semi-)log canonical point, like a non-smooth point at simple normal crossing degeneration. In such a situation, natural metric limits are expected to collapse. Therefore, to extend our approach, one can instead naturally ask the existence of minimal non-collapsing bubble (with infinite diameter) as defined in [Od22a, Definition 4.2] if it exists. Nevertheless the situation is more intricate (cf., e.g., [Od22a, 4.4], [HSVZ22, SRZ23]) and far beyond the main setup of this paper.
Here, we fix such degeneration of polarized Calabi-Yau varieties over a punctured disk $\Delta^{*}$ to which we cannot fill in any log terminal polarized Calabi-Yau variety as the central fiber. For that, we define some subclass of possible irreducible components of degenerations, which we conjecture to form a large subset of minimal non-collapsing limits of the Calabi-Yau metrics. In this case, it sounds very natural to consider dlt minimal models to study these general bubbles, which led to the notion of “galaxy models” ([Od22b, §2]): connected locally ringed space which contains infinitely many bubbling log Calabi-Yau varieties densely. Then, we have been conjecturing that
Conjecture 3.3 ([Od22a, §4], [Od22b, §2]).
For any non-log-terminal semi-log-canonical (locally stable in the sense of Kollár) degeneration of polarized log terminal projective Calabi-Yau varieties $(\mathcal{X},\mathcal{L})\to\Delta$, we enhance it to (any) polarized galaxy model $\tilde{\mathcal{X}}\to\mathbbm{k}[[t^{\mathbb{Q}}]]$ and consider the klt locus of $\tilde{\mathcal{X}}^{\rm klt}_{0}$ its central fiber $\tilde{\mathcal{X}}_{0}$.
Each of its connected component appear as the minimal non-collapsing (in the sense of [Od22a]) pointed Gromov-Hausdorff limits (bubbling) of the Calabi-Yau metrics, for some sequence of base points $p_{t}$ in the general fibers $\mathcal{X}_{t}$.
In the case of families of polarized abelian varieties this is confirmed by explicit matrix calculations after Siegel [Od19] and further we know these galaxy models represent the non-archimedean Calabi-Yau metric as confirmed essentially in [Gub07, Gub10] (see [Od22b, §2]). Although the converse to the above conjecture is not true for non-flat cases, the author believes that the above bubbles forms the basic large subclass of minimal non-collapsing limits in general.
Acknowledgements   The author heartily thanks S.Sun and C.Spotti for the helpful conversations over many years on this topic. He also thanks H.Blum, S.Boucksom, M.deBorbon, S.K.Donaldson, D.Halpern-Leistner, M.Jonsson, Y.Kusakabe, M.Mauri and T.T.Nghiem for helpful comments and discussions. During the research, the author was partially supported by Grant-in-Aid for Scientific Research (B) 23H01069, 21H00973, Grant-in-Aid for Scientific Research (A) 21H04429, 20H00112, and Fund for the Promotion of Joint International Research (Fostering Joint International Research) 23KK0249. Part of this work is done during the author’s stay at Paris, notably Institut de mathématiques de Jussieu - Paris Rive Gauche from March 2024 and Imperial College London in June 2024. The author received there above-mentioned benefitial comments, and nice pleasant working environments. We appreciate their kind hospitality.
Appendix A Higher $\Theta$-stable reduction revisited
Recall that [Od24b, §3] introduces a higher rank generalization of the stable reduction avoiding $\Theta$-strata [HL14, AHLH23] and the construction in our Theorem 2.4, 2.6 and the properness of the moduli of Calabi-Yau cone in [Od24a] are partly its applications (in the irregular case). Since our original proof in [Od24b, §3] depends on sophisticated stack theoretic framework and techniques ([AHHLR22, HL14]), here we give more scheme theoretic reexplanation a priori conditionally (the conditions are often satisfied in practice e.g., for the above-mentioned applications).
First, we give scheme theoretic version of the higher $\Theta$-strata.
We work over any field $\mathbbm{k}$. Let us set $\mathcal{M}:=[M/G]$, where $M$ is a finite type $\mathbbm{k}$-scheme acted by an algebraic $\mathbbm{k}$-group $G$. We follow and partially use the basic notations of [Od24b]. For instance, $N\simeq\mathbb{Z}^{r}$ is a lattice, $T=\mathbb{G}_{m}\otimes N$ is an algebraic $\mathbbm{k}$-split torus, $\tau\subset N\otimes\mathbb{R}$ is a rational polyhedral cone and corresponding affine toric variety is denoted as $U_{\tau}$. In addition, here we denote $m\colon T\times T\to T$ is the multiplication and $a\colon G\times M\to M$ is the action.
Definition A.1 (after [Od24b, §3]).
We call a higher $\Theta$-strata on $[Z^{+}/G]\subset\mathcal{M}$ for $G$-invariant locally closed subscheme $Z^{+}(\subset M)$ of type $\tau\subset N\otimes\mathbb{R}$ is locally liftable if the corresponding $\mathcal{Z}^{+}\subset\underline{\rm Map}(\Theta_{\tau}=[U_{\tau}/T],[M/G])$ with ${\rm ev}={\rm ev}_{(1,1,\cdots,1)}\colon\mathcal{Z}^{+}\simeq[Z^{+}/G]$ is given by the following data:
  • a Zariski open covering $\{M^{\prime}\}$ of an open neighborhood of $Z^{+}$ in $M$.
  • For each $M^{\prime}$, we have an algebraic $\mathbbm{k}$-morphism $M^{\prime}\to{\rm Hom}_{\rm alg.\mathbbm{k}-grp}(T,G)$ (homomorphism family), which we encode by $\psi\colon T\times M^{\prime}\to G$, that satisfies the following:
    1. ($\mathrm{i}$)
      $a\circ(\psi\times{\rm id}_{M^{\prime}})\colon T\times M^{\prime}\to M$ gives $a^{\prime}\colon T\times M^{\prime}\to G\times M\to M$ whose image lies in $M^{\prime}$ and is a $T$-action (action-ness),
    2. ($\mathrm{ii}$)
      $a^{\prime}|_{(M^{\prime}\cap Z^{+})}$ extends to $U_{\tau}\times Z^{+}\to Z^{+}$.
    Here, ${\rm Hom}_{\rm alg.\mathbbm{k}-grp}(T,G)$ naturally means the algebraic $\mathbbm{k}$-scheme consists of homomorphisms of the algebraic $\mathbbm{k}$-groups, which naturally contains $G/Z(T)$ with the centralizer $Z(T)$ of $T$.
In practice, many examples of (higher) $\Theta$-strata satisfies the above local liftability condition and indeed we do not know any example which do not satisfies it. Under this a priori assumption, we can give a purely scheme-theoretic proof of the higher $\Theta$-stable reduction theorem of [Od24b]) as follows. We recall and formulate the statement in a simplified manner, as it is practically enough for above mentioned applications.
Theorem A.2 (cf., [Od24b, 3.8]).
Suppose a $\mathbbm{k}$-morphism $f\colon{\rm Spec}\mathbbm{k}[[t]]\to M$ satisfies that the closed point maps to a locally liftable (Definition A.1) higher $\Theta$-strata $Z^{+}$ and the generic point maps to outside $Z^{+}$. Replace $f$ by its base change $f\colon\mathbbm{k}[[t^{\frac{1}{d}}]]\to M$ for some $d\in\mathbb{Z}_{>0}$. Then, (as an elementary modification [AHLH23, §3.5]) there is a $\mathbbm{k}$-morphism $f^{\prime}\colon{\rm Spec}\mathbbm{k}[[t]]\to M\setminus Z^{+}$ with $f^{\prime}|_{\mathbbm{k}((t))}\in G(\mathbbm{k}((t)))\cdot f(\mathbbm{k}((t)))$.
The following arguments reproves Theorem 3.8 of [Od24b], but it does not depend on sophisticated results on algebraic stack theory, such as [AHHLR22], [HL14, 4.4.6] nor that of Białynicki-Birula type result ([HL14, 1.4.7, 1.4.8]). Nevertheless, the basic scheme-theoretic part of the proof of Theorem 3.8 of [Od24b] is still the same.
Proof.
We follow the proof of [Od24b, Theorem 3.8] but modify the arguments as follows. First, we take $(Z^{+})^{\prime}\subset M$ as Definition A.1 which contains $f(p)$ where $p$ denotes the closed point of ${\rm Spec}\mathbbm{k}[[t]]$. Set $a^{\prime}(p_{\tau},p)=:p^{\prime}$. Then we consider the $T$-action $a^{\prime}$ on a neighborhood of $p^{\prime}$ as $M^{\prime}\subset M$ and resolve the indeterminancy of the rational map $U_{\tau}(\supset T)\times\Delta\dashrightarrow M^{\prime}$ by applying Lemma 3.11 of [Od24b]. We recall what the claim of loc.cit implies briefly: after replacing $\tau$ to smaller enough subcone, it gives an appropriate rational polyhedral subcone of $\tau^{\prime}\subset\tau\times\mathbb{R}_{\geq 0}$ which projects isomorphically to $\tau^{\prime}$, set $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}$ and consider the corresponding affine toric variety $U_{\hat{\tau}^{\prime}}$, as well as its $1$-dimensional orbit $V(\tau^{\prime})(\simeq\mathbb{G}_{m})$ which corresponds to $\tau^{\prime}$. Then, $f$ extends to $U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]\to M$ so that $V(\tau^{\prime})\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$ maps outside $Z^{+}$, due to loc.cit.
Further, taking a rational ray $\mathbb{R}_{\geq 0}(v,d)\subset\tau^{\prime}$ for a primitive integral vector $v\in\tau$ and $d\in\mathbb{Z}_{>0}$, we obtain ${\rm Spec}\mathbbm{k}[[t^{\frac{1}{d}}]]\to U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$. Composing with the above $U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]\to M$, we obtain the required replacement $f^{\prime}$ as a desired modification. ∎
We note that in the proof of our Theorem 2.4 where we applied the original higher $\Theta$-stable reduction (Theorem 3.8 of [Od24b]), we proved the local liftability of the higher $\Theta$-strata in the above sense (Definition A.1), so that our above easier arguments A.2 suffices.
Now we assume $\mathbbm{k}$ is algebraically closed field of characteristic 00.
Then, we similarly show that the higher $\Theta$-strata which appears in the proof of properness of the moduli of Calabi-Yau cones is locally liftable in the sense of Definition A.1, so that the above easier version (Theorem A.2) suffices.
Theorem A.3 ([Od24a, 3.12]).
Fixing any positive integer $n$, $V$, Reeb vector field $\xi$, consider the moduli stack $\mathcal{M}_{V}^{\rm ss}(\xi)$ of $n$-dimensional K-semistable Fano cones of the minimized normalized volume $V$ and Reeb vector field $\xi\in N\otimes\mathbb{R}$. Then, $\mathcal{M}_{V}^{\rm ss}(\xi)$ is universally closed i.e., satisfies the existence part of the valuative criterion of properness.
We follow the arguments of §3.2 and §3.5 of op.cit. We write $T:=N\otimes\mathbb{G}_{m}$.
Proof.
We prove by the valuative criterion. Recall that $\mathcal{M}_{V}^{\rm ss}(\xi)$ is defined as a quotient stack $[B_{V}^{\rm ss}/G]$ in loc.cit, and $B_{V}^{\rm ss}$ can be enlarged to a larger $G$-acted algebraic $\mathbbm{k}$-scheme $B_{V}$ which is stratified as $B_{V}^{\rm ss}=B_{V}(0)\subset B_{V}(1)\subset\cdots\subset B_{V}(m^{\prime})=B_{V}$ by the value of the minimized normalized volumes again as in [Od24a, §3.2]. Thanks to the boundedness result [XZ24], now one can and do assume that each $B_{V}(i)$ contains all $n$-dimensional K-semistable Fano cone of the same volume density $V_{i}$, with $V=V_{0}>V_{1}>\cdots V_{m^{\prime}}$.
We take an arbitrary $\mathbbm{k}$-morphism ${\rm Spec}\mathbbm{k}[[t]]\to B_{V}$ such that the image $f(p)$ of the closed point $p$ lies in a connected component $Z^{+}$ of $B_{V}(i)\setminus B_{V}(i-1)$. (Recall that by the arguments in [Od24a], following [LX14], it follows that any $\mathbbm{k}$-morphism $\mathbbm{k}$-morphism ${\rm Spec}\mathbbm{k}((t))\to B_{V}^{\rm ss}$ can be extended to such for sufficiently large $B_{V}$ and $i$).
[…]
Then, if we restrict to $Z^{+}=B$, $\pi^{-1}(B)=:\mathcal{U}_{B}\to B$ has $(T\times)T_{B}$-action and canonical ideal sequences $I_{\lambda_{i}}\subset\mathcal{O}_{\mathcal{U}_{B}}$ for a discrete subset $\lambda_{i}\subset S(B)\subset\mathbb{R}_{\geq 0}$ with $0=\lambda_{0}<\lambda_{1}<\cdots$, which satisfies that $I_{\lambda_{i}}/I_{\lambda_{i+1}}$ is locally free $\mathcal{O}_{B}$-module of rank $r_{i}$ ([Che24, 1.1 (also §5)]). Here, $S(B)$ is the holomorphic spectrum (in the sense of [DS17]) corresponding to the $\pi$-fibers over $B$. Thanks to the fiberwise $T=N\otimes\mathbb{G}_{m}$-action, one can decompose $I_{\lambda_{i}}$ as $\oplus_{m\in M=N^{\vee}}I_{\lambda_{i},m}$ to the finite dimensional $T$-eigenspaces. Then, each $I_{\lambda_{i},m}/I_{\lambda_{i+1},m}$ is locally free $\mathcal{O}_{B}$-module of the finite rank $r_{i,m}$.
Now, consider the K-semistable degeneration of $[X]=f(p)\in B_{V}$ as $X\rightsquigarrow W$ and realize it as the $T_{B}$-action for some inclusion $T_{B}\subset G$ induced by the splitting of $\mathcal{O}_{X}\supset I_{\lambda_{1}}|_{X}\supset I_{\lambda_{2}}|_{X}\supset\cdots$. The limit $[W]$ still lies inside $B_{V}$ by the construction of $B_{V}$ mentioned above, as the minimized normalized volume remains the same as that of $X$ (see §2.1 and [XZ24]). We extend this inclusion $T_{B}\subset G$ as a homomorphism family locally around $[W]$ as follows. Since the $G$-action on $\mathcal{U}_{B}$ and the fiberwise $T$-action commutes, for each $\lambda_{i}$ and $m\in M^{\prime}$, we have a canonical $T_{B}(\times T)$-equivariant splitting of $\mathbbm{k}$-vector spaces $I_{\lambda_{i},m}|_{W}\twoheadrightarrow(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})|_{W}$ due to the complete reductivity of $T_{B}$. We extend it as follows. If we consider $I_{\lambda_{i},m}$ and $(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})$ as $\mathcal{O}_{B}$-modules, they are both locally free and inherit the $T_{B}$-action by the canonicity of $I_{\lambda_{i}}$s. Then, if we shrink $B$ to a small enough $T_{B}$-invariant open affine neighborhood of $[W]$ in $B$ by Sumihiro’s theorem, the $T_{B}$-equivariant vector bundles restricted to it are the direct sum of $T_{B}$-equivariant line bundles (cf., [IS15, 2.5], [Pay08, 2.1], [Kly89]).
By the same arguments as loc.cit, the $T_{B}$-equivariant surjection between $I_{\lambda_{i},m}\twoheadrightarrow(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})$ is $T_{B}$-equivariantly splittable by the complete reductivity of $T_{B}$ as $(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})\to I_{\lambda_{i},m}$ which corresponds to $T_{B}$-semiinvariant fiberwise coordinates on $\mathcal{U}_{B}=\pi^{-1}(B)$. We extend it to $T_{B}$-semiinvariant functions on $\pi^{-1}(B_{V}^{o})$ for a $T_{B}$-invariant open affine neighborhood $B_{V}^{o}$ of $[W]$ in $B_{V}$ which vanish at the 00-section($\simeq B_{V}^{o}$). If we collect them for all $m\in M^{\prime}$ and many enough $i$, and compare with the embedding of $\mathcal{U}\to\mathbb{A}^{l}\times B_{V}$ (for the same $M^{\prime}$), then this gives family of linear transformation $g(b)\in G\subset{\rm GL}(l)$ for each $b\in B_{V}^{o}$, which a priori non-trivially varies over $B_{V}^{o}$.
Then the family of homomorphism $((Z^{+})^{\prime}:=)B_{V}^{o}\to{\rm Hom}(T_{B},G)$ given by $b\mapsto g(b)^{-1}T_{B}g(b)$ ensures the local liftability of the higher $\Theta$-strata structure on $Z^{+}=B$, so that Theorem A.2 applies. By iterating the same procedure $i$-times, we obtain the desired modification ${\rm Spec}\mathbbm{k}[[t]]\to B_{V}^{\rm ss}$ of $f$, so that we conclude the assertion by the valuative criterion. ∎